### v1

# Nested and outlier embeddings into trees

Abstract
In this paper, we consider outlier embeddings into HSTs and ultrametrics. In particular, for $(X,d)$, let $k$ be the size of the smallest subset of $X$ such that all but that subset (i.e. the “outlier set”) can be probabilistically embedded into the space of HSTs with expected distortion at most $c$. Our primary result is showing that there exists an efficient algorithm that takes in $(X,d)$ and a target distortion $c$ and samples from a probabilistic embedding with at most $O(\frac{k}{\epsilon}\log^{2}k)$ outliers and distortion at most $(32+\epsilon)c$, for any $\epsilon>0$. This leads to better instance-specific approximations for certain instances of the buy-at-bulk and dial-a-ride problems, whose current best approximation algorithms go through HST embeddings.
In order to facilitate our results, we largely focus on the concept of compositions of nested embeddings introduced by [20]. A nested embedding is a composition of two embeddings of a metric space $(X,d)$ – a low distortion embedding of a subset $S$ of nodes, and a higher distortion embedding of the entire metric. The composition is a single embedding that preserves the low distortion over $S$ and does not increase distortion over the remaining points by much. In this paper, we expand this concept from the setting of deterministic embeddings to the setting of probabilistic embeddings. We show how to find good nested compositions of embeddings into HSTs, and combine this with an approximation algorithm of Munagala et al. [39] to obtain our results.
1 Introduction
Low distortion metric embeddings are an indispensable algorithmic tool in many contexts. Many algorithmic problems are easier to solve on trees than on general metrics, so embedding a given metric into a tree and solving the problem on the tree provides a way to obtain an approximately optimal solution on the original metric. However, the final approximation factor depends on the distortion (the multiplicatively worst change in any pairwise distance) of the embedding. This motivates the problem of finding an embedding from general metrics into trees of the least possible distortion. A series of works [3, 11, 12, 28] showed that although some metrics require linear distortion to embed into trees, probabilistic embeddings can do much better: it is possible to embed any $n$ point metric into a distribution over trees such that the expected distance between any pair of points increases by at most $O(\log n)$ [28], while the cycle requires a factor of $\Omega(n)$ for any deterministic tree embedding [3]. In fact, the probabilistic upper bound applies to embeddings into a special family of trees called hierarchically separated trees, or HSTs. The particular structure of these rooted trees – with edge lengths decreasing geometrically over levels going from the root to a leaf – is algorithmically useful in a variety of problems.
While uniform distortion bounds are helpful in understanding the general effectiveness of using metric embeddings to solve a problem, from a practical perspective, a more pertinent question is how well a particular given metric $(X,d)$ can embed into trees. Indeed, for some metrics the optimal distortion is $\Omega(\log n)$ but for others it may be much smaller, e.g. $O(1)$, providing much better algorithmic results. Is it possible to efficiently obtain the best embedding for the given metric into trees? Munagala et al. [39] considered this question from the viewpoint of approximation and provided a constant factor approximation to the instance-optimal distortion.
In this paper we further investigate instance-optimal distortion for probabilistic embeddings into trees by exploring the role of outliers. For a finite metric $(X,d)$, a $(k,c)$-outlier embedding is a map from “most of” $X$, namely all but $k$ vertices that we call outliers, into another metric space $Y$ such that the distortion over the non-outliers is bounded by $c$. Intuitively, we can view these outliers as noisy or adversarial data that may hurt the quality of the instance-optimal embedding from $X$ into $Y$. The goal of the outlier embedding is to identify and remove these outliers upfront so as to obtain a much better embedding over the remaining points. Algorithmically, given a target distortion $c$, we want to find the smallest $k$ such that the removal of $k$ outliers enables embedding the rest of the metric with distortion at most cc. [^1] In Section 3.1, we will discuss some specific examples where removing a small number of points can greatly improve overall distortion. For deterministic embeddings, this problem turns out to be NP-hard when the host metric is a tree or ultrametric [44]. We extend the notion of outlier embeddings to probabilistic output, and give bicriteria approximations focusing on minimizing the outlier set size when given a target distortion for the final probabilistic embedding. Our main result is as follows.
Theorem 1.1:
Let $(X,d)$ be any metric that admits a $(k,c)$ probabilistic outlier embedding into HSTs. Then for any constant $\epsilon>0$, there exists an algorithm that efficiently samples from a probabilistic embedding into HSTs of some $S\subseteq X$ with $|X\setminus S|\leq O(\frac{1}{\epsilon}k\log^{2}k)$ and expected distortion at most $(32+\epsilon)c$.
This result allows us to obtain improved instance-specific approximation ratios for several graph optimization problems (e.g. the buy-at-bulk and dial-a-ride problems) on input metrics that meet certain structural requirements. We describe these applications in Section 3.
Technical tools and challenges
A key technical tool in our approach is the concept of compositions of nested embeddings that was first introduced by Chawla and Sheridan [20] in the context of outlier embeddings into the Euclidean metric. A nested embedding seeks to take some “good” embedding on a subset $S\subseteq X$, as well as a set of “okay” embeddings on all subsets of $X$ with size at most $|X\setminus S|+1$; It then finds a new embedding on the entire set $X$ such that distortion is still good for pairs of points in $S$ and is not too bad for any pair of points. Chawla and Sheridan use nested embeddings to ensure a bound on the optimal solution to an SDP, which they then round to determine their outlier set and embedding.
We combine the ideas of Chawla and Sheridan with the approximation algorithm of Munagala et al. [39] to obtain outlier embeddings into HSTs. Munagala et al.’s algorithm rounds a natural linear programming (LP) relaxation of probabilistic embeddings into HSTs. We extend this LP to incorporate outliers; develop a nested composition into HSTs to guide the choice of distortion constraints in the LP; and provide a rounding similar to that in [39]. The way we use nested compositions to enable finding bicriteria approximations through an LP rounding scheme uses some of the techniques of Chawla and Sheridan [20]. However, we still have two key technical challenges.
Nested compositions into HSTs with bounded contraction.
To bound the modified LP of Munagala et al. [39] that we use in this setting, we must show existence of nested compositions into HSTs that are non-contracting. In particular, Munagala et al.’s LP essentially sets up a metric labeling problem with the goal of assigning each vertex to a representative at geometrically increasing levels. Each vertex $v$ is assigned a representative at each level $i$ that is at a distance no more than $2^{i}$ from $v$. Importantly, the level-wise labeling constraints cannot be weakened for fractional outliers – each vertex, whether or not it’s an outlier, must still be assigned at level $i$ to a representative within distance $2^{i}$ from it. Even slightly weakening this constraint can potentially introduce errors at all levels of the hierarchy, leading to very large distortion (depending on $|X|$ rather than $k$ and $c$).
Nested compositions successively “stitch together” pairs of embeddings on the outliers and the non-outliers. For normed metrics, such as those considered by [20], when contraction is allowed, it suffices to directly concatenate the respective embeddings. Our setting, in contrast, calls for a careful alignment and stitching of the two embeddings that does not contract any distances. We call this procedure a “merge” and characterize the conditions under which it returns a correct nested embedding. We show that such a merge function exists for nested embeddings into HSTs.
Probabilistic nested embedding.
The second technical challenge is in extending the notion of nested embeddings to probabilistic embeddings. [^2] Note that while the embeddings are probabilistic, we require the outlier set to be deterministic. In particular, our nested embedding guarantees are designed specifically with respect to a deterministic outlier set as, by definition, we obtain tighter bounds on the distortion on non-outliers relative to the outliers. Allowing the outlier set to be random is an interesting direction for future work, but would require a significant modification to the manner in which we use nested embeddings to obtain outlier embeddings. Although we are not aware of explicit examples of metrics where removing a random set of outliers provides even better distortion than removing a deterministic outlier set, it is possible that this extension would provide further improvement in the instance-specific approximation factors for some applications. This is important in the context of tree embeddings because, as mentioned previously, probabilistic tree embeddings can obtain exponentially smaller distortion relative to deterministic tree embeddings.
Consider extending a probabilistic embedding of some subset $S$ of a given metric $(X,d)$ into a host space $Y$ to a probabilistic embedding of the entire metric $X$ into $Y$. One potential way of leveraging a deterministic nested embedding is to draw a random embedding $\alpha$ mapping $S$ to $Y$, extend this embedding, and return the extension. What can we say about the distortion of this random extension? Let $c_{\alpha}:=\max_{u,v\in X}(d_{\alpha}(u,v)/d(u,v))$ denote the (probabilistic) distortion achieved by the (probabilistic) embedding $\alpha$, where $d_{\alpha}$ is the distance function in the host metric composed with the mapping $\alpha$. Then, the deterministic nested embedding we construct guarantees us a distortion linear in $c_{\alpha}$. This implies that the distortion of the final random embedding returned by this approach is at worst linear in $\mathbb{E}_{\alpha}[c_{\alpha}]$. However, $\mathbb{E}_{\alpha}[c_{\alpha}]$, can be much larger than the expected distortion of the original embedding, $\max_{u,v\in X}(\mathbb{E}_{\alpha}[d_{\alpha}(u,v)]/d(u,v))$. To avoid this issue, we show that our nested embeddings into HSTs satisfy a certain obliviousness property, and we show that this property lets us obtain a final distortion guarantee that depends on $\max_{u,v\in X}(\mathbb{E}_{\alpha}[d_{\alpha}(u,v)]/d(u,v))$.
To summarize, our primary contributions are as follows.
  • We extend the concept of nested embeddings to the setting of probabilistic embeddings. We give a general algorithm for composing probabilistic nested embeddings into a given host space, so long as we can design a “merge” algorithm with certain properties. We show that the set of embeddings into HSTs has such a merge algorithm and as a result has such an algorithm for composing embeddings.
  • We discuss how to extend the notion of outlier sets to probabilistic embeddings and show that if a given metric $(X,d)$ has an embedding into HSTs with at most $c$ distortion and at most $k$ outliers, we can sample from a probabilistic embedding with at most $(32+\epsilon)c$ distortion and at most $O(\frac{k}{\epsilon}\log^{2}k)$ outliers, for any $\epsilon>0$.
  • We also achieve similar bicriteria approximation results for the weighted outlier problem, where the cost of an outlier set is the sum of the weights assigned to its nodes, and the goal is to find a minimum cost outlier set such that the remaining metric embeds into HSTs with distortion $c$.
  • We present several algorithmic applications of outlier embeddings into HSTs.
Related work
Embeddings are a widely studied concept, and there are many algorithms that obtain “uniform” distortion bounds for embedding arbitrary $n$-point metrics into a particular target space. A series of results [3, 12, 18, 28] beginning with Bartal [11] studied how to probabilistically embed a general metric into HSTs with only poly-logarithmic distortion. The seminal paper of Fakcharoenphol, Rao, and Talwar [28] showed that all $n$-point metrics have a probabilistic embedding into the space of hierarchically separated trees with distortion at most $O(\log n)$. We will use this general algorithm to construct embeddings on small subsets of a metric that we hope to compose with an “optimal” embedding on some larger subset of that metric.
There is also a variety of papers that study instance-optimal distortion for embedding finite metrics into a given target space. That is, the goal in such embeddings is to minimize the distortion needed for the given specific $(X,d)$, which in some cases may be smaller than the best bound achievable for all metrics of that size. Some examples of such papers cover embeddings into constant dimensional Euclidean space [8, 26, 27, 37, 45], the line [37, 29, 7, 41], trees [21, 9], and ultrametrics [4, 39].
Because of their simple structure, HST embeddings are useful for a wide variety of algorithmic applications, including ones related to the $k$-server problem [24, 10], hierarchical clusterings [15, 43, 25], Steiner trees [30], buy-at-bulk network design [6], metric labeling [34], vehicle routing problems [32], and online clustering [22].
The notion of outlier embeddings was initially defined by Sidiropoulos, Wang, and Wang [44] where they studied deterministic embeddings into ultrametrics, trees, and constant-dimensional $\ell_{2}$ space. But their techniques are largely unrelated to ours. One part of their work considers deterministic, isometric outlier embeddings (where the distortion $c$ is 1), and uses structural properties to find approximations. These structural properties don’t apply to non-isometric embeddings. Another part considers bicriteria approximations, but bounds additive distortion (that is, the maximum absolute difference between old and new distances, rather than their ratio). Techniques for bounding additive distortion tend to be quite different from those for multiplicative distortion, the focus of this work. Moreover, additive distortion does not provide approximation guarantees for algorithmic applications.
Chubarian and Sidiropoulos [23] continued the work on outlier embeddings and studied non-isometric outlier embeddings of unweighted graph metrics into the line. Later, Chawla and Sheridan [20] studied deterministic outlier embeddings into $\ell_{2}$ space with unbounded dimension. There is a notion related to outlier embeddings, called embeddings with slack, in which the goal is to minimize the number of edges that are distorted by more than some target factor, and there has been significant work in that area as well [1, 16, 35]. Outlier embeddings are also closely related to metric Ramsey theory [13, 38]. Given inputs $n$ and $c$, metric Ramsey theory asks: What is the largest $m$ such that all metrics of size $n$ have a submetric of size $m$ that embeds into some target space with distortion at most $c$? In this paper we are primarily focused on instance-optimal solutions rather than universal bounds that apply to all metrics, and we seek to approximate the size of the outlier set rather than the size of the non-outlier set. Existing results for metric Ramsey theory give universal bounds for deterministic embeddings into $\ell_{p}$ spaces and ultrametrics.
Our notion of nested embeddings was first defined in [20], but many other types of embedding compositions or extensions have been considered before. One particular example is the notion of a Lipschitz extension [33, 40, 14], which seeks to leave non-outliers unaltered from the original embedding and ensure the expansion (but not necessarily the contraction) of every pair of nodes is bounded. In fact, our nested embedding algorithm is very similar to a part of the Lipschitz extension algorithm presented by Calinescu et al. [14] as a part of their solution to the 0-extension problem. Other works study a variety of different embedding compositions [2, 5, 19, 42, 36]. These notions largely differ from ours in that they emphasize bounding the worst distortion for any pair in the entire metric, while our notion focuses on bounding the worst distortion between any pair subject to a “special set” having even stricter distortion constraints.
2 Definitions and results
In this section, we will define terms used in this paper and summarize our results.
Expanding embeddings and distortion
A metric space is a set of points $X$ with distances $d_{X}(\cdot,\cdot)$ between them that obey the triangle inequality. That is, for all $x,y,z\in X$, $d_{X}(x,z)\leq d_{X}(x,y)+d_{X}(y,z)$. We assume without loss of generality (via scaling) that in the metrics we consider, the minimum distance between distinct pairs of points is at least $1$: $\min_{x\neq y\in X}d_{X}(x,y)\geq 1$.
A deterministic embedding of a space $(X,d_{X})$ into a space $(Y,d_{Y})$ is a map $\alpha:X\rightarrow Y$. We are interested in the distortion (i.e. multiplicative change) of pairwise distances under this map. We use, as a convenient shorthand, $d_{\alpha}(x,y):=d_{Y}(\alpha(x),\alpha(y))$ to denote the distance between the images of two points $x,y\in X$ under this map, and $d(x,y)$ to denote the original distance $d_{X}(x,y)$ when the space $(X,d_{X})$ is clear from the context. We will without loss of generality consider only expanding embeddings, or embeddings for which $d_{\alpha}(x,y)\geq d(x,y)$ for all $x,y\in X$.
We can further expand these definitions to the setting of probabilistic embeddings. In particular, $\mathcal{D}$ is a probabilistic embedding of $(X,d_{X})$ into $\mathcal{Y}$ if for each function $\alpha$ in the support of $\mathcal{D}$, $\alpha:X\rightarrow Y(\alpha)$ for some $(Y(\alpha),d_{Y(\alpha)})\in\mathcal{Y}$. This embedding has distortion $c$ if for all $u,v\in X$ we have:
$d_{\alpha}(u,v)$ $\geq d(u,v)\quad\quad\forall\alpha\in\operatorname{support}(\mathcal{D})$
$\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq c\cdot d(u,v).$
Observe that every embedding in the support of $\mathcal{D}$ is expanding, but pairwise distortion is bounded in expectation over the randomness in $\mathcal{D}$.
We use the notation $\alpha:X\xhookrightarrow{c}Y$ to denote a deterministic embedding with distortion $c$ from $(X,d_{X})$ into $(Y,d_{Y})$ and $\mathcal{D}:X\xhookrightarrow{c}_{p}\mathcal{Y}$ to denote a probabilistic expanding embedding with distortion $c$ from $(X,d_{X})$ into $\mathcal{Y}$.
Outlier embeddings
In this paper, we extend the concept of outlier embeddings to its probabilistic counterpart. Note that probabilistic embeddings can, in general, obtain much smaller distortion into a class of metric spaces than deterministic embeddings, which motivates extending the study of such embeddings to this setting. If $\mathcal{D}$ is a probabilistic embedding from $(X,d_{X})$ into $\mathcal{Y}$ and $K\subseteq X$, define $\mathcal{D}_{X\setminus K}$ to be the distribution identical to $\mathcal{D}$ but with all functions $\alpha$ in the support restricted to the set $X\setminus K$ (i.e. we can sample from $\mathcal{D}_{X\setminus K}$ by sampling from $\mathcal{D}$ and restricting the domain of the function drawn).
Definition 2.1:
A probabilistic embedding $\mathcal{D}$ of $(X,d_{X})$ into $\mathcal{Y}$ is a probabilistic $(k,c)$-outlier embedding if there exists $K\subseteq X$ such that $|K|\leq k$ and $\mathcal{D}_{X\setminus K}$ is a probabilistic embedding of $(X\setminus K,d_{X})$ into $\mathcal{Y}$ with distortion $c$.
Note that in Definition 2.1, there is a deterministic subset $K$ whose distances to other nodes (and each other) we don’t care about. The randomness is only on the embedding for the remaining nodes in $X\setminus K$. One motivation for outlier embeddings is as a way to prevent noise or adversarial data from preventing good embeddings, and we believe that making the outlier set itself deterministic is consistent with this motivation. In Section 3.4 we will also discuss a specific problem where the application naturally lends itself to finding a deterministic outlier set.
Composition of nested embeddings
Chawla and Sheridan [20] introduce the concept of a composition of nested embeddings. This concept considers two (partial) embeddings from a metric space $(X,d_{X})$ into space $(Y,d_{Y})$. One embedding $\alpha_{S}$ maps a subset $S\subset X$ into $Y$ with distortion $c_{S}$, and the other maps the entire space $X$ into $Y$ with distortion $c_{X}>c_{S}$. The goal is to combine these embeddings into a single embedding from $X$ into $Y$ that preserves the distortion $c_{S}$ over pairs vertices in $S$ and obtains distortion comparable to $c_{X}$ over the remaining pairs of vertices.
We will now extend this concept to a probabilistic setting. We assume that both the “good” embedding from $S$ into a space $\mathcal{Y}$ as well as the “poor” embedding from $X$ into $\mathcal{Y}$ are probabilistic. Furthermore, as in [20] the composition only needs to apply the “poor” embedding over subsets of $X$ of size at most $k+1$ where $k:=|X\setminus S|$. Therefore, for the tightest distortion bounds, we will assume that we are given an assortment of probabilistic embeddings (one for each subset of $X$ of size $k+1$) into $\mathcal{Y}$ and we will use $c_{X}$ to denote the worst distortion of any embedding in this assortment.
Definition 2.2 (Probabilistic composition of nested embeddings, modified from [20]):
Let $(X,d_{X})$ be a metric space with $S\subset X$, $k:=|X\setminus S|$, and $\mathcal{Y}$ a class of metrics. Consider an assortment of embeddings: $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$, and for all subsets $K^{\prime}\subseteq X$ with $|K^{\prime}|\leq k+1$, $\mathcal{D}_{K^{\prime}}:K^{\prime}\xhookrightarrow{c_{X}}_{p}\mathcal{Y}$.
A (strong) $(\ell,g)$-nested composition of the assortment $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$, where $\ell$ is a constant and $g:[1,\infty)^{2}\times\mathbb{N}\rightarrow[1,\infty)$, is an algorithm that given the assortment returns a probabilistic embedding $\mathcal{D}:X\xhookrightarrow{}_{p}\mathcal{Y}$ such that:
$\text{for all }u,v\in S,\text{ and for all }\alpha\in\operatorname{support}(\mathcal{D}),$ $d_{\alpha}(u,v)$ $\geq d(u,v)$ (1)
$\text{for all }u,v\in S,$ $\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq\ell\cdot c_{S}\cdot d(u,v),$ (2)
$\text{and, for all }u,v\in X,$ $\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq g(c_{S},c_{X},k)\cdot d(u,v).$ (3)
Chawla and Sheridan primarily considered weak nested compositions that do not require non-contraction of distances for pairs of points not both in $S$. For our application to outlier embeddings into HSTs, restricting contraction of all points is necessary.
We do consider weak nested compositions as well, and we will briefly discuss some minor results for probabilistic Lipschitz extensions in Appendix D.
Hierarchically Separated Trees (HSTs)
In this paper, we are particularly interested in embedding into distributions of hierarchically separated trees, which are metrics defined as follows:
Definition 2.3:
A metric $(X,d)$ is a $\beta$-hierarchically separated tree ($\beta$-HST) if there is a map from $X$ to the leaves of a tree such that the following hold:
  • Each node $u$ in the tree is associated with some value $\eta_{u}$, and $\eta_{v}=\beta\cdot\eta_{u}$ whenever $u$ is a child of $v$.
  • For all leaves $v$, $\eta_{v}=1/\beta$.  [^3] In this paper, we will use the term $\beta$-HSTs to refer to what other sources sometimes call exact $\beta$-HSTs. These sources typically define non-exact $\beta$-HSTs as only requiring $\eta_{v}\geq\beta\cdot\eta_{u}$.
  • For two leaves of the tree $x,y\in X$ with least common ancestor $z$, $d(x,y)=\eta_{z}$.
For simplicity, we will primarily focus on $2$-HSTs, and in referring to HSTs, we mean $2$-HSTs unless otherwise noted. Note that focusing on $\beta$-HSTs for a different constant $\beta$ at worse changes results by a constant factor. We may also write $\mathcal{H}$ to mean the set of $2$-HSTs.
Main results
In order to achieve our main outlier embedding result, we will show in Section 4 that under the right conditions there exist good nested compositions of probabilistic embeddings. The right conditions will depend on the existence of a separate “merging” algorithm for combining embeddings, and we will show in Section 5 that such an algorithm exists for HSTs. This gives us following theorem that we will use to prove Theorem 2.5.
Theorem 2.4 (HSTs have strong nested embeddings):
Given a probabilistic embedding $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}$ HSTs, there exists an embedding $\mathcal{D}:X\xhookrightarrow{}_{p}\mathcal{H}$ with distortion at most $4c_{S}$ on pairs of nodes in $S$ and distortion at most $\zeta\cdot(c_{S}\log k+\log^{2}k)$ on other pairs, where $k=|X\setminus S|$ and $\zeta$ is a constant.
Further, we can efficiently sample the from the distribution $\mathcal{D}$ if we can efficiently sample from $\mathcal{D}_{S}$. We will also briefly discuss the implications of this theorem for ultrametric embeddings in Appendix E
Using Theorem 2.4 and a modification of the LP from Munagala et al. [39] we get the following result.
Theorem 2.5:
Given a metric $(X,d)$ and target distortion $c$, let $k$ be the optimal outlier set size for probabilistically embedding $(X,d)$ into HSTs. Then for any constant $\epsilon>0$, there exists an algorithm that efficiently samples from a probabilistic embedding into HSTs of some $S\subseteq X$ with $|X\setminus S|\leq O(\frac{1}{\epsilon}k\log^{2}k)$ and expected distortion at most $(32+\epsilon)c$.
The LP-based techniques we use also allow us to obtain the following result for the weighted variant of the problem, in which we are given weights $w_{i}$ on all nodes $i$ in the metric space. The cost of a subset $K$ of the metric space is $w(K):=\sum_{i\in K}w_{i}$, and the goal is to find an outlier embedding with some target distortion such that the cost of the outlier set is minimized.
Corollary 2.6:
Given a metric $(X,d)$, node weights $w$, and target distortion $c$, let $w^{*}$ be the optimal outlier set cost for probabilistically embedding $(X,d)$ into HSTs with $c$ distortion and let $k$ be the size of such an optimal outlier set. Then for any constant $\epsilon>0$, there exists an algorithm that efficiently samples from a probabilistic embedding into HSTs of some $S\subseteq X$ with $w(X\setminus S)\leq O(\frac{1}{\epsilon}w^{*}\log^{2}k)$ and expected distortion at most $(32+\epsilon)c$.
The rest of the paper will proceed as follows:
  • In Section 3, we discuss the buy-at-bulk, dial-a-ride, and minimum communication cost tree problems as algorithmic settings where Theorem 2.5 may be useful. We also discuss the existence of metrics for which removing a small number of points greatly improves their minimum distortion for probabilistic HST embeddings.
  • In Section 4, we show how to generalize the $\ell_{1}$ nested composition algorithm of Chawla and Sheridan [20] and characterize the type of input merge functions that allow this algorithm to work. We also show how the “obliviousness” of this algorithm allows it to work in both deterministic and probabilistic settings.
  • In Section 5, we show how to find a good merge function for HST embeddings, and we show how to modify the algorithm of Munagala et al. [39] such that the existence of good nested compositions for HST embeddings results in a good approximation factor for this modified algorithm for the outlier setting.
Finally, in the appendices we will discuss some of the proofs omitted from the main paper and we will discuss a set of conditions on deterministic Lipschitz extensions that allow them to work in the probabilistic setting.
3 Applications
In this section, we will discuss several algorithmic applications of probabilistic outlier embeddings into HSTs in order to motivate our overall results.
3.1 Bad metrics
To motivate this work, we may begin by asking: do there exist metrics for which removing a small number of points results in a significantly better probabilistic embedding into HSTs? In this subsection, we will discuss types of metrics for which this is the case. In particular, in Figure 1, we give an example of an unweighted graph metric that is formed by “linking” two unweighted graphs through an edge. If we select one of these graphs to be $K_{n-\log n}$ (the complete graph on $n-\log n$ nodes) and the other to be a constant degree expander graph on $\log n$ nodes (denoted $G_{\log n}$), we get a graph on $n$ total nodes. However, Bartal [11] shows that, for any $x$, certain expander graphs $G_{x}$ incur a distortion $\Omega(\log x)$ when probabilistically embedding into HSTs. Thus, the graph in Figure 1 requires $\Omega(\log\log n)$ distortion for HST embeddings (inheriting the distortion bound of the submetric $G_{\log n}$), whereas the same graph with $G_{\log n}$ removed can be embedded with $O(1)$ distortion.
$K_{n-\log n}$${\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}G_{\log n}}$
Figure 1: We can consider an unweighted graph on $n$ nodes such that one part of the graph is $K_{n-\log n}$ (the complete graph on $n-\log n$ nodes), and the remaining $\log n$ nodes form some constant-degree expander graph $G_{\log n}$. The two parts of the graph can then be connected by a single edge. Note that the subgraph induced by each of these two sub-parts of this graph is an isometric subgraph of the original graph.
This graph is highly specific and the “good” and “bad” parts of the graph are easy to spot. We will use this graph as an example in the coming subsections because its simplicity makes it easy to analyze, but we also note that the class of metrics whose optimal distortion can be improved by removing a small number of vertices is much more diverse than this simple example.
To demonstrate a large family of metrics with good outlier embeddings, we will turn to the notion of metric composition as defined by Bartal et al. [13].
Definition 3.1 (Metric composition [13]):
Let $(M,d_{M})$ be a metric space with minimum distance exactly $1$, and let $\mathcal{N}=\set{(N_{x},d_{N_{x}})}_{x\in M}$ be a set of metric spaces. Let $\Delta:=\max_{x\in M}\set{diameter(N_{x})}$. Then the $\beta$-composition of $M$ and $\set{N_{x}}$ denoted $M^{*}=M_{\beta}[\mathcal{N}]$ is a metric on the vertex set $\set{(x,u):x\in M,u\in N_{x}}$ such that distances in $M^{*}$ are
$d_{M^{*}}((x,u),(y,v)):=\begin{cases}d_{N_{x}}(u,v)&x=y\\ \beta\cdot\Delta\cdot d_{M}(x,y)&else\end{cases}.$
We provide a visualization of this composition in Figure 2. Bartal et al. [13] show that for any $M$ and $\mathcal{N}$, $M_{\beta}[\mathcal{N}]$ defines a proper metric if $\beta\geq 1/2$. While it’s easy to factor a composition back into its original components if $\beta$ is quite large, it’s less clear how to do so when $\beta=1/2$.
$N_{x}$$N_{y}$$N_{z}$$\beta\Delta\cdot d_{M}(x,y)$$\beta\Delta\cdot d_{M}(y,z)$$\beta\Delta\cdot d_{M}(x,z)$$\Delta:=\max_{w\in\{x,y,z\}}(diameter(N_{w}))$
Figure 2: A visual example of metric composition where $M$ is some metric on the set $\set{x,y,z}$.
The highest distortion required to embed some metric in $\set{M}\cup\mathcal{N}$ into HSTs is a lower bound on the minimum distortion to embed $M_{\beta}[\mathcal{N}]$ into HSTs, and in Theorem 3.2 we show that in fact this quantity is an upper bound as well (up to constant factors).
Theorem 3.2:
Let $(M,d_{M})$ be a metric with distortion at most $c$ for probabilistically embedding into HSTs, and let the metrics in $\mathcal{N}=\set{(N_{x},d_{N_{x}})}_{x\in M}$ also probabilistically embed with distortion at most $c$. Then there is a probabilistic embedding of $M_{\beta}[\mathcal{N}]$ into HSTs with distortion at most $O(c)$.
The proof of the theorem is deferred to Appendix A.1.
3.2 Buy-at-Bulk
In the (uniform) buy-at-bulk problem, [^4] We assume the metric version of the problem rather than the graph version, as Awerbuch and Azar [6] note that they are equivalent. we are given a metric space $\mathcal{X}$ as well as a set of demand pairs $(s_{i},t_{i})$, each with a demand $d_{i}$ and a sub-additive cost function $g$. (We will assume wlog that $g(1)=1$, as this is always possible by rescaling.) The goal is to select a single path $P_{i}$ for each demand pair in a way that minimizes $\sum_{e}\ell_{e}g(\sum_{i:e\in P_{i}}d_{i})$, where $\ell_{e}$ is the length of edge $e$ and all the sum is over all edges. This has natural applications in many areas of operation research.
A key result of Awerbuch and Azar [6] states that if $\mathcal{D}$ is a probabilistic tree embedding of metric $(X,d)$ with distortion $c$, then the optimal cost of the buy-at-bulk problem on $(X,d)$ is at most a factor of $c$ less than the expected value of the optimal solution to the same problem on the embedded metric. The buy-at-bulk problem is trivial on trees, where there is only one path between each pair of nodes. Thus, using the results of Munagala et al. [39] we can obtain an expected approximation ratio $O(c)$, where $c$ is the optimal distortion for probabilistically embedding the input into HSTs.
In this subsection, we will work to improve this approximation ratio for certain metrics by solving the problem separately on a set of requests only involving a set of points with a low distortion HST embedding and then on the remaining set of requests, which we seek to ensure has relatively small total cost. We formalize this idea in Algorithm 1.555Algorithm 1 actually gives a more general framework for solving the buy-at-bulk or dial-a-ride problem, and as a result it asks for an oracle that solves the underlying problem on HSTs. For the buy-at-bulk problem on trees, the only feasible solution is optimal, so $\mathcal{A}$ is trivial. To outline this approach, we must first define some terms that will help us set up a “weighted outlier” problem on the input metric space.
In particular, consider a buy-at-bulk instance $(\mathcal{X},\set{(s_{i},t_{i},d_{i})}_{i=1}^{m})$. Let $OPT_{i}$ be the cost of the optimal solution to the buy-at-bulk instance $(\mathcal{X},(s_{i},t_{i},d_{i}))$ (i.e. the same instance but with only demand pair $i$). This is easy to solve optimally, as we can simply pick a shortest $s_{i}$-$t_{i}$ path and purchase $d_{i}$ capacity on that path. We will consider $OPT_{i}$ to be the “cost” of demand pair $i$. In order to decide the “cost” of removing a node $x$ from $\mathcal{X}$, we will sum the costs of all demand pairs that involve $x$. That is, let $R_{x}:=\set{i:x\in\set{s_{i},t_{i}}}$, and we will set the weight of $x$ to be $w(x):=\sum_{i\in R_{x}}OPT_{i}$. Further, for any $K\subseteq X$, let $w(K)=\sum_{x\in K}w(x)$. This weighting of the points of $\mathcal{X}$ sets up a weighted outlier problem for $\mathcal{X}$.
For a given metric $\mathcal{X}$ with weights $w$ and any choice of $c\geq 1$, we will define $\text{OutlierCost}(c):=\min_{K\subseteq X;(X\setminus K)\xhookrightarrow{c}_{p}\mathcal{H}}w(K)$ to be the cost of the cheapest set we can leave out while still embedding the remaining metric into HSTs with at most $c$ distortion. We will also be interested in the size of such a cheap set, so we define $\mathcal{K}(c):=|\arg\min_{K\subseteq X;(X\setminus K)\xhookrightarrow{c}_{p}\mathcal{H}}w(K)|$. In Theorem 3.3, we give a bound on the cost of Algorithm 1. Note that for some constant $\eta$, if $c=\eta\log n$, [^6] In this section, we let $\eta$ be a constant such that for all $n$-point metric spaces $X$, $X\xhookrightarrow{\eta\log n}_{p}\mathcal{H}$. (Such a constant is guaranteed to exist by [28].) we have $\text{OutlierCost}(c)=0$, so Algorithm 1 never obtains worse than an $O(\log n)$ approximation.
Algorithm 1 Approximation algorithm for dial-a-ride or buy-at-bulk
Input: metric $\mathcal{X}$, request set $\set{(s_{i},t_{i})}_{i=1}^{m}$, approximation algorithm $\mathcal{A}$ for buy-at-bulk or dial-a-ride on HSTs, integer $\ell$ if dial-a-ride and demands $\set{d_{i}}$ if buy-at-bulk
Output: A buy-at-bulk or dial-a-ride solution
1:for $x=1$ to $n$ do
2:  $R_{x}\leftarrow\set{i:x\in\set{s_{i},t_{i}}}$
3:  Let $OPT_{i}$ be the cost of an optimal solution that only fulfills only the $i$th request
4:  $w_{x}\leftarrow\sum_{i\in R_{x}}OPT_{i}$
5:for $i=0$ to $\log\log n$ do
6:  Let $c_{i}\leftarrow\eta\log n/2^{i}$ $\triangleright$ here $\eta$ is the appropriate constant from [28]
7:  Run the weighted version of Algorithm 4 with input distortion $c_{i}$ and $\epsilon=1$
8:  Receive an outlier set $K_{i}$ of cost $w^{i}$ and an embedding $\alpha_{i}$ of distortion $2c_{i}$
9:  Embed $X\setminus K_{i}$ into HSTs
10:  Use $\mathcal{A}$ to solve the relevant problem on requests $R_{1}:=\set{j:\set{s_{j},t_{j}}\cap K_{i}=\emptyset}$ and nodes in $X\setminus K_{i}$
11:  Let $sol_{i}^{1}$ be the solution obtained
12:  Let $R_{2}:=\set{j:\set{s_{j},t_{j}}\cap K_{i}\neq\emptyset}=R\setminus R_{1}$
13:  Find a minimum distortion embedding of $X_{i}=K_{i}\cup\set{x:\exists y\in K_{i}\text{ s.t. }\set{x,y}\in R}$ into HSTs
14:  Use $\mathcal{A}$ to find a solution to the problem with node set $X_{i}$ and request set $R_{2}$
15:  Let $sol_{i}^{2}$ be the solution obtained
16:  Find the “naive” solution to the problem with node set $X_{i}$ and request set $R_{2}$
17:  Let $sol_{i}^{3}$ be the solution obtained
18:  Let $sol_{i}$ be the solution obtained by combining solution $sol_{i}^{1}$ with either $sol_{i}^{2}$ or $sol_{i}^{3}$, whichever is cheaper
19:Return the solution of lowest cost
Theorem 3.3:
Let $(\mathcal{X},R=\set{(s_{i},t_{i},d_{i})}_{i=1}^{m})$ be an instance of the buy-at-bulk problem with an overall solution of cost at most $OPT$.
Then Algorithm 1 on this instance (and with the trivial oracle $\mathcal{A}$ for buy-at-bulk) returns a solution of cost at most
[…]
In fact, for any $c$ directly considered by the algorithm in line 6, the algorithm’s cost is bounded by
$O({c\cdot OPT+\text{OutlierCost}(c)\cdot\log^{2}(\mathcal{K}(c))}$.
We defer the formal proof to Appendix A.2.
An example where the approximation ratio improves
Let $n$ be such that $\log n$ is an even integer dividing $n$. Let $M$ be an HST metric of size $n/\log n$ and diameter $\ell\geq\log^{4}n$, and let $\mathcal{N}=\set{N_{x}}_{x\in M}$ be a set of HST metrics, each of size $\log n$. Let $\mathcal{N}^{\prime}=\set{N_{x}^{\prime}}_{x\in M}$ be the same as $\mathcal{N}$, but for one $x^{*}\in M$, replace $N_{x^{*}}$ with the metric of an unweighted expander graph that does not embed well into HSTs. Let $M^{*}:=M_{1/2}[\mathcal{N}^{\prime}]$ be a metric of size $n$.
Consider an instance of the buy-at-bulk problem on $M^{*}$ defined as follows
  • For each $x\in M$, divide the nodes of $N_{x}^{\prime}$ into pairs $(u,v)$, and make $((x,u),(x,v))$ a demand pair in $M^{*}$ with demand $1$.
  • Let $x_{1},x_{2}\neq x^{*}$ be a pair of nodes in $M$ with distance $\ell$. (We assume $x^{*}$ is chosen such that such a pair exists.) Select some $u\in N_{x_{1}}$ and $v\in N_{x_{2}}$ and add $((x_{1},u),(x_{2},v))$ as a demand pair with demand $1$.
Note that $M^{*}$ requires distortion $\Omega(\log\log n)$ to probabilistically embed into HSTs (due to the presence of an expander graph of size $\log n$ as a submetric), so if we run the algorithm of Munagala et al. [39] and buy paths on the resulting tree, we could potentially pay as much as $\Omega(\log\log n\cdot OPT)$, depending on which distances were distorted.
Now consider the result of running Algorithm 1 on this instance. Note that if we remove from $M^{*}$ the nodes corresponding to $N_{x^{*}}^{\prime}$, we get a metric that embeds into HSTs with constant distortion, by Theorem 3.2. We also know that the outlier cost of the nodes corresponding to $N_{x^{*}}^{\prime}$ is at most $\log^{2}n$, as there are $O(\log n)$ demand pairs associated with these nodes; each such pair has demand $1$; nodes in $N_{x^{*}}^{\prime}$ only participate in demand pairs with other nodes in $N_{x^{*}}^{\prime}$; and the diameter of $N_{x^{*}}^{\prime}$ is at most $\log n$ (due to being the path metric of an unweighted graph on $\log n$ nodes). Thus, Algorithm 1 gives us a solution of cost at most $O(OPT+\log^{4}n)=O(OPT)$, where we used the fact that $OPT\geq\ell\geq\log^{4}n$, as there is some demand pair with distance at least $\ell$.
3.3 Dial-A-Ride
In the dial-a-ride problem, we have a van of capacity $\ell$ and we are given a metric space $(X,d)$ as well as a set of transportation requests $\set{(s_{i},t_{i})}_{i=1}^{m}$ from $m$ people such that person $i$ wants the van to transport them from location $s_{i}$ to location $t_{i}$. The goal of the problem is to find a schedule of pickups and drop-offs for the van such that: [^7] Technically we should also require that the tour start and end at some specified “depot” location, but solving this relaxed version of the problem and then adding on beginning and ending drops from/to the depot only worsens the overall approximation ratio by a constant factor (as each of the new trips we add has cost at most $OPT$), so we will discuss this simplified version instead.
  1. 1.
    The number of passengers in the van is the number that have been picked up minus the number that have been dropped off, and it is always between 0 and $\ell$,
  2. 2.
    For any customer $i$, that customer is picked up at their starting location $s_{i}$ and dropped off later at their destination $t_{i}$, and
  3. 3.
    The cost of the solution is the sum of distances traveled by the van between each consecutive stop on its schedule.
This problem is NP-hard, as it generalizes the Traveling Salesman Problem [17]. Note that if we serve only a single request $i$, the problem is easy and incurs cost $d(s_{i},t_{i})$ in the optimal solution. Thus, computing $OPT_{i}$ for each request $i$ is efficient, and we can run Algorithm 1 for this problem.
However, we will need to use a more complicated algorithm $\mathcal{A}$ to solve the dial-a-ride problem on HSTs than we did for buy-at-bulk. In particular, we will consider an algorithm $\mathcal{A}$ that gets approximation ratio $r_{n,m,\ell}$ for dial-a-ride instances on HSTs with $n$ nodes, $m$ requests, and $\ell$ van capacity. In Theorem 3.4, we consider the approximation ratio achieved by Algorithm 1 with such an input $\mathcal{A}$. Again note that for $c=\eta\cdot\log n$, $\text{OutlierCost}(c)=0$, so the worst-case approximation ratio is $O(\log n)$.
Theorem 3.4:
Let $\mathcal{A}$ be an approximation algorithm for the dial-a-ride problem on HSTs that achieves approximation ratio $r_{n,m,\ell}$ for instances of $n$ nodes, $m$ requests, and $\ell$ van capacity.
Let $(\mathcal{X},\set{(s_{i},t_{i})}_{i=1}^{m},\ell)$ be a dial-a-ride input such that $OPT$ is the cost of an optimal solution for this instance, $\Delta$ is the diameter of $\mathcal{X}$, and $\text{OutlierCost}(c)$ and $\mathcal{K}(c)$ are as in Section 3.2. Then Algorithm 1 achieves a feasible solution with expected cost at most
[…]
In fact, for any $c$ directly considered by the algorithm in line 6, the algorithm’s cost is bounded by
$O({c\cdot OPT+\text{OutlierCost}(c)\cdot\log^{2}(\mathcal{K}(c))}$.
We prove this theorem in Appendix A.3. Note that this algorithm’s performance is generally better on instances whose optimal cost is much bigger than $\Delta$ (such as instances in which the total number of requests is large compared to $\ell\Delta$).
Charikar et al. [18] give an algorithm $\mathcal{A}$ for dial-a-ride on HSTs with $r_{n,m,\ell}=O(\sqrt{\ell})$. Thus, we get an algorithm that produces a solution such that for any $c\geq 1$, the solution’s cost is at most $O(\sqrt{\ell}\cdot c\cdot OPT+\Delta\cdot\text{OutlierCost}(c)\log^{2}n)$. If each location is involved in at most $t$ requests (i.e. the demand is “spread out”), then the solution also has cost at most $O(\sqrt{\ell}\cdot(c+\log(t\cdot\mathcal{K}(c)))\cdot OPT)$ (discussed further in Appendix A.3).
An example where the approximation ratio improves
We can use the metric in Figure 1 as an example where the approximation ratio for this problem improves. In particular, assume $n$ is even and $\log n$ is an integer and consider an instance of dial-a-ride where the nodes of this metric are partitioned into pairs such that each pair serves as one dial-a-ride request. Because this metric requires $\Omega(\log\log n)$ distortion to embed into HSTs, the cost guarantee of Munagala et al. [39] combined with Charikar and Raghavachari [17] on this metric is at best $O(\sqrt{\ell}\log\log n\cdot OPT)$. However, $G_{\log n}$ forms an outlier set of this metric with cost $O(\log^{2}n)$, as there are at most $2\log n$ requests involving nodes in $G_{\log n}$ and the diameter of the graph is $O(\log n)$. The remaining graph without $G_{\log n}$ embeds into HSTs with distortion $O(1)$. Thus, using the algorithm of Charikar and Raghavachari [17], Algorithm 1 achieves a solution of cost $O(\sqrt{\ell}\cdot OPT+\log^{5}n)$. However, we note that $OPT\geq n$, as any optimal solution must visit every node in the graph at least once, so the cost is $O(\sqrt{\ell}\cdot OPT)$.
3.4 Minimum cost communication tree with outliers
In the Minimum Cost Communication Tree (MCCT) problem, we are given a finite metric $\mathcal{X}=(X,d)$ and weights $r_{ij}$ for all pairs of nodes $i,j$ and we are asked for a non-contracting embedding $\alpha$ of $X$ into a tree such that $\sum_{i,j\in X}r_{ij}\cdot d_{\alpha}(i,j)$ is minimized.
This problem is motivated by evolutionary biology, in which a scientist may attempt to reconstruct a tree of relationships between multiple organisms, and she comes to the problem with some known minimum distances between them, derived through a method such as DNA sequence analysis. In such a setting, the $r_{ij}$ are an additional tunable parameter that can allow the scientist to indicate pairs she considers “more likely” to be related, by giving them a bigger $r_{ij}$ value. Alternatively, we could imagine a scenario where we are looking for a tree that minimizes the average communication cost of a set of messages, where nodes $i$ and $j$ send $r_{ij}$ messages to each other and the communication cost of a message is the distance it travels. As noted by Wu et al. [46], this problem can be approximated in expectation to within a factor of $O(\log n)$ by simply embedding randomly into an HST.
Directly applying the algorithm of Munagala et al. [39] with the observations of Wu et al. [46], we have a randomized algorithm that gets an $O(c(\mathcal{X}))$-approximation for the MCCT problem, where $c(\mathcal{X})$ is the minimum distortion needed to probabilistically embed $\mathcal{X}$ into HSTs.
It is quite natural to consider an outlier version of the MCCT problem. In the evolutionary biology setting, we could consider the case that we are given some noisy data points, or points that don’t really belong in the set being analyzed. In the message-sending setting, we could consider a situation where there is a budget of some number of messages whose communication costs we allow to be arbitrarily large.
One approach to this type of outlier problem may be to look for a $k$ outlier tree embedding of minimum distortion, but the set of outliers that minimizes the embedding distortion does not necessarily also minimize the MCCT solution cost. [^8] For example, points that are very distant from all other points provide a large discount when they are discarded, though it turns out that using this observation to construct a greedy strategy does not work here either, even when the input is a tree metric. We may expect that the outlier MCCT problem is easier to solve on tree metrics than on general metrics, so we could instead try the following approach: embed the input metric into a tree, and then try to find the best outliers for the resulting tree metric. In fact, if no outliers are allowed, MCCT is trivial on tree metrics. Further, we show in Appendix A.4 that if the input is a tree metric, we can obtain a bicriteria approximation to the more general $k$-outlier MCCT problem. In particular, if the optimal cost of a $k$ outlier solution for a given instance of a tree metric is $OPT_{k}$, then we give an algorithm that finds a solution with at most $3k$ outliers and at most $24OPT_{k}$ cost.
Consider an instance $\mathcal{X}$ of the MCCT problem, and let $OPT_{k}$ be the value of an optimal $k$-outlier solution for $\mathcal{X}$. Let $c_{k}(\mathcal{X})$ be the minimum distortion required for probabilistically embedding all but at most $k$ points in $\mathcal{X}$ into HSTs, and let $c(\mathcal{X})=c_{0}(\mathcal{X})$. We show in Appendix A.4 that if we combine the above bicriteria approximation with the Munagala et al. [39] algorithm for HST embeddings, we get a result that uses at most $3k$ outliers and has cost at most $O((1+\epsilon)\cdot c(\mathcal{X})\cdot OPT_{k})$. Meanwhile, if we replace this embedding algorithm with the outlier embedding algorithm from this paper, we get a solution with at most $O(\frac{k}{\epsilon}\log^{2}k)$ outliers and $O((1+\epsilon)\cdot c_{k}(\mathcal{X})\cdot OPT_{k})$ cost, where $\epsilon$ is a tunable parameter larger than 0. Thus, for some metrics we are able to greatly improve the approximation factor on the total cost, while removing a polylogarithmically larger number of outliers.
4 Probabilistic nested compositions of embeddings
In this section we will develop an approach for constructing probabilistic strong nested compositions into arbitrary classes of metrics. Our approach is based on the nested composition technique of Chawla and Sheridan [20] for $\ell_{1}$ embeddings but strengthens that technique by introducing a general procedure for merging two metrics.
Recall that we are given a metric space $(X,d_{X})$, a target class of metrics $\mathcal{Y}$, a subset $S$ with $k=|X\setminus S|$, and an assortment of probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$. Our goal is to produce a single probabilistic embedding for the entire set $X$ that has small expansion. Chawla and Sheridan consider the deterministic case where we start with a deterministic embedding $\alpha_{S}$ from $S$ to $\mathcal{Y}$. At a high level, their algorithm works as follows. [^9] We were notified after the publication of [20] that the Lipschitz extension algorithm of that paper (a building block of the nested embedding algorithm presented here) is essentially the same as the rounding algorithm presented in Section 2 of Calinescu et al. [14], with the following changes: (1) [14] use a semimetric obtained from an LP solution to the 0-extension problem everywhere that we use $d$, (2) [14] pick a uniformly random permutation over $S$ (which they refer to as terminals) rather than taking a random permutation over $\gamma^{-1}(K)$ as we do here (note that |S|=nk|S|=n-k and $|\gamma^{-1}(K)|\leq k$), and (3) [14] pick $b$ from the range $[1,2)$. Notably, if their algorithm is modified to use a random permutation of $\gamma^{-1}(K)$ rather than $S$, a slight modification of their analysis obtains the some of the same results that we do, with tighter constants. It partitions the set of “outliers” $K=X\setminus S$ into multiple sets $K_{1},K_{2},\ldots$ in a certain random fashion. It also identifies for each set $K_{i}$ a “close” vertex $\gamma_{i}\in S$. In the composition embedding, it then maps every vertex in $K_{i}$ to the image of $\gamma_{i}$ under $\alpha_{S}$.
There are two challenges that we need to address in order to obtain strong compositions for probabilistic embeddings. First, we need to ensure that the composition does not contract distances. In [20]’s Algorithm 4.1, although the expansion for every pair of points can be bounded for this composition, distances can get contracted by arbitrarily large factors – indeed pairs of vertices in the same $K_{i}$ end up colocated under the composition. Our algorithm proceeds in a fashion similar to that of [20]. We randomly partition $K$ into sets $K_{1},K_{2},\ldots$ as before, and identify a “close” vertex $\gamma_{i}\in S$ for each such set. Then, we consider the embedding in the given assortment that maps $K_{i}\cup\{\gamma_{i}\}$ into $\mathcal{Y}$. We “merge” this embedding with the embedding $\alpha_{S}$ in such a fashion that neither embedding by itself gets distorted any further. This allows us to bound contraction in a similar way to Algorithm 5.1 from [20], though we frame it more generally. We then iterate in this manner over all of the components in the partition of $K$.
The second challenge is to combine the per-pair $\alpha_{S}$ expansion guarantee into a expected distortion guarantee when we draw an $\alpha_{S}$ from the given probabilistic embedding $\mathcal{D}$. The key property that allows us to do so is the independence of the procedure for partitioning $K$ and the structure of $\alpha_{S}$. We call this property obliviousness and will revisit it more formally in Appendix D.
We will now identify properties of the merging procedure that we require in order to obtain guarantees for the final composition.
Definition 4.1 (Perfect merge):
Let $Z_{1},Z_{2}\subseteq X$ with $Z_{1}\cap Z_{2}=\{v\}$. Let $\alpha_{1}:Z_{1}\rightarrow Y_{1}$ and $\alpha_{2}:Z_{2}\rightarrow Y_{2}$ be deterministic expanding embeddings, for $Y_{1},Y_{2}\in\mathcal{Y}$. We say that an embedding $\alpha:Z_{1}\cup Z_{2}\rightarrow Y$ for some $Y\in\mathcal{Y}$ is a perfect merge of $\alpha_{1}$ and $\alpha_{2}$ if
  • For all $x,y\in Z_{1}$, $d_{\alpha}(x,y)=d_{\alpha_{1}}(x,y)$.
  • For all $x,y\in Z_{2}$, $d_{\alpha}(x,y)=d_{\alpha_{2}}(x,y)$.
  • For $x\in Z_{1}\setminus\set{v}$ and $y\in Z_{2}\setminus\set{v}$, we have $d_{\alpha}(x,y)\geq\max\set{d_{\alpha_{1}}(x,v),d_{\alpha_{2}}(v,y)}$.
We say that a procedure Merge is a perfect merge function for embedding $X$ into $\mathcal{Y}$ if given any pair $\alpha_{1}$ and $\alpha_{2}$ of embeddings from subsets of $X$ into $\mathcal{Y}$ satisfying the conditions above, it produces a perfect merge of the pair.
Now we present Algorithm 2 (Nested Composition) that given an assortment of probabilistic embeddings on subsets of $X$, produces a new embedding that is drawn from a strongly nested composition of the input embeddings. Note that Algorithm 2 is a generalization of Algorithm 5.1 from [20]. We highlight in red the changes we have made. In that algorithm, the input PerfectMerge function was merely concatenation of vectors and the target metric was $\ell_{1}$. We will develop a different PerfectMerge merge function for embeddings into HSTs in Section 5.
Algorithm 2 Nested Composition: Algorithm for finding a composition of probabilistic nested embeddings
Input:
  • Metric space $(X,d)$, subset $S\subseteq X$, $k=|X\setminus K|$.
  • Assortment of embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ where $\mathcal{D}_{S}:S\rightarrow\mathcal{Y}$ and $\mathcal{D}_{K^{\prime}}:K^{\prime}\rightarrow\mathcal{Y}$ for $K^{\prime}\subset X$.
  • A function PerfectMerge that takes in two deterministic embeddings $\alpha:Z\rightarrow\mathcal{Y}$ and $\alpha^{\prime}:Z^{\prime}\rightarrow\mathcal{Y}$ for $Z,Z^{\prime}\subset X$ and $|Z\cap Z^{\prime}|=1$, and returns an embedding $\alpha^{\prime\prime}:Z\cup Z^{\prime}\rightarrow\mathcal{Y}$.
Output: A randomly sampled embedding $\alpha:X\rightarrow\mathcal{Y}$.
[…]
2: Sample $\alpha_{S}\leftarrow\mathcal{D}_{S}$
3:Define a function $\gamma:K\rightarrow S$ such that $\gamma(u)\in\arg\min_{v\in S}d(u,v)$. $\triangleright$ i.e. $\gamma(u)$ is one of $u$’s closest neighbors in $S$
4:Select $b$ uniformly at random from the range $[2,4]$
5:Select a uniformly random permutation $\pi:K\rightarrow[k]$ of the vertices in $K$
[…]
7:for $i=1$ to $k$ do
8:   $u_{i}\leftarrow\pi^{-1}(i)$
9:   $K_{i}\leftarrow\{v\in K^{\prime}\ |\ d(v,u_{i})\leq b\cdot d(v,\gamma(v))\}$ $\triangleright$ Let the “center” of $K_{i}$ be $u_{i}$
10:   Sample $\alpha_{i}\leftarrow\mathcal{D}_{K_{i}\cup\set{\gamma(u_{i})}}$
11:   $\alpha\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$
12:  $K^{\prime}\leftarrow K^{\prime}\setminus K_{i}$
13:Output $\alpha$
The main result of this section is as follows.
Theorem 4.2:
Let PerfectMerge be a perfect merge function for embeddings from $(X,d)$ into a given class $\mathcal{Y}$. Then, given an assortment of probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ and the procedure PerfectMerge, Algorithm 2 returns a sample from a strong probabilistic $(3,g)$-nested composition with $g(c_{S},c_{X},k)=O((c_{S}+c_{X}(k))\cdot H_{k})$, where $H_{k}$ is the $k$th harmonic number. Here we use $c_{X}(k)$ to denote the maximum distortion of $\mathcal{D}_{K^{\prime}}$ over all subsets $K^{\prime}$ with $|K^{\prime}|\leq k+1$.
We break down the proof of the theorem into two parts: bounding the expansion of the embedding (Lemma 4.3) and its contraction (Lemma 4.4).
Lemma 4.3:
Let PerfectMerge be a perfect merge function for embeddings from $(X,d)$ into a given class $\mathcal{Y}$. Then, given an assortment of expanding probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ and the procedure PerfectMerge, Algorithm 2 returns a random embedding $\alpha$ from $X$ to $\mathcal{Y}$ such that for all $x,y\in X$,
$\mathop{E}_{\alpha}[d_{\alpha}(x,y)]$ $\leq(125\,c_{S}+116\,c_{X}(k))\cdot H_{k}\cdot d(x,y).$
Further, if $x,y\in S$, then $\mathop{E}_{\alpha}[d_{\alpha}(x,y)]\leq c_{S}\cdot d(x,y)$.
Lemma 4.4:
Let PerfectMerge be a perfect merge function for embeddings from $(X,d)$ into a given class $\mathcal{Y}$. Then, given an assortment of expanding probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ and the procedure PerfectMerge, Algorithm 2 returns an embedding $\alpha$ from $X$ to $\mathcal{Y}$ such that for all $x,y\in X$, $d_{\alpha}(x,y)\geq d(x,y)/3$.
The proof of the expansion lemma (Lemma 4.3) follows a very similar argument to that given for Algorithm 5.1 in [20]. The main difference is that it allows the input embedding $\mathcal{D}_{s}$ to be probabilistic, and uses the fact that each $\alpha_{S}$ drawn from this distribution is independent of the partition $K_{i}$ as well as the embeddings $\alpha_{i}$. For completeness, we include the full proof in Appendix B.
We will now prove Lemma 4.4. An essential consequence of using a perfect merge function with Algorithm 2 is that when bounding expansion or contraction for a pair $x,y$, we need only bound this value at the point in the algorithm where they both appear in the domain of $\alpha$ for the first time. At any subsequent iteration, as $\alpha$ undergoes further merge steps, the first condition in Definition 4.1 ensures that the distance between $x$ and $y$ stays the same. We summarize this observation in the following fact.
Fact 1:
Let $x,y\in X$. Consider the first iteration of Algorithm 2 in which both $x$ and $y$ are in the domain of $\alpha$, and let $\hat{d}$ be $d_{\alpha}(x,y)$ at that time step. Then at every subsequent iteration of the algorithm, $d_{\alpha}(x,y)=\hat{d}$.
We now proceed to bound the contraction. Note that we do get a small amount of contraction, so to actually obtain an expanding embedding, we need to be able to “scale up” the set of points we map to while still staying in $\mathcal{Y}$.
Proof of Lemma 4.4 We will actually prove a slightly stronger claim about contraction: for $x,y\in X$, we have
  1. 1.
    If $x$ and $y$ are in $S$, then $d_{\alpha}(x,y)\geq d(x,y)$.
  2. 2.
    If at least one of $x,y$ is in $S$, then $d_{\alpha}(x,y)\geq d(x,y)/2$.
  3. 3.
    For all $x,y$, $d_{\alpha}(x,y)\geq d(x,y)/3$.
As established by Fact 1, we need only focus on the time when $x$ and $y$ first appear in the domain of $\alpha$.
  1. 1.
    Consider $x,y\in S$. Note $\alpha$ is first defined as $\alpha_{S}\leftarrow\mathcal{D}_{S}$, so at this point $x$ and $y$ are both in its domain and $d_{\alpha}(x,y)=d_{\alpha_{S}}(x,y)\geq d(x,y)$, by the fact that all embeddings in $\mathcal{D}_{S}$ are expanding.
  2. 2.
    Now consider $x\in S,y\in X\setminus S$. We show that no matter the choice of $\set{K_{j}}$ (the partition of the outliers), the claim holds. In particular, consider some choice of $\set{K_{j}}$ with $y\in K_{i}$ for some $i$. Then let $\alpha_{i}$ be in the support of $\mathcal{D}_{K_{i}}$, and define $\alpha^{\prime}\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$, where $\alpha$ is the embedding at the start of this iteration in the algorithm. Also let $\gamma(u_{i})$ be the vertex common to the two embeddings that are being merged. Then we get from Definition 4.1
    $d_{\alpha^{\prime}}(x,y)$ $\geq\max\set{d_{\alpha}(x,\gamma(u_{i})),d_{\alpha_{i}}(\gamma(u_{i}),y)}$
    $\geq\max\set{d(x,\gamma(u_{i})),d(\gamma(u_{i}),y)}$
    $\geq d(x,y)/2$
    where the first line is because PerfectMerge is a perfect merge function, the second line is by the first part of this claim and the fact that $\alpha_{i}$ is expanding, and the third line is by applying the triangle inequality on $d(x,y)$.
  3. 3.
    Finally, consider $x,y\in X\setminus S$. Again, fix a choice of the partition $\set{K_{j}}$ from the algorithm. If $x,y\in K_{i}$ for the same $i$, let $\alpha_{i}$ be in the support of $\mathcal{D}_{K_{i}}$ and let $\alpha^{\prime}\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$. Then $d_{\alpha^{\prime}}(x,y)=d_{\alpha_{i}}(x,y)\geq d(x,y)$, where we used the fact that we have a perfect merge function and $\alpha_{i}$ is expanding.
    Thus, we are left with the case that $x\in K_{i},y\in K_{j}$ for $i\neq j$. Wlog assume $i<j$. Then $x$ and $y$ first appear in the domain of $\alpha$ at iteration $j$ of the algorithm. Let $\alpha$ be the embedding before iteration $j$ and let $\alpha^{\prime}$ be the embedding that results from the merge in iteration $j$. Let $\alpha_{j}$ be drawn from the support of $\mathcal{D}_{K_{j}}$. Then, recalling that $x$ is in the domain of $\alpha$ prior to the merge, we get
    $d_{\alpha^{\prime}}(x,y)$ $\geq\max\set{d_{\alpha}(x,\gamma(u_{j})),d_{\alpha_{j}}(\gamma(u_{j}),y)}$
    $\geq\max\set{d(x,\gamma(u_{j}))/2,d(\gamma(u_{j}),y)}$
    $\geq d(x,y)/3,$
    where the first line is because we have a perfect merge function; the second line is by the second case of this claim, the fact that $\gamma(u_{j})\in S$, and the fact that all embeddings in $\mathcal{D}_{K_{j}}$ are expanding; the third is by the triangle inequality applied to $d(x,y)$. $\square$
Theorem 4.2 follows almost immediately from Lemmas 4.3 and 4.4, with one minor change. Technically, we defined a nested composition to be deterministically non-contracting, and here we have only shown that contraction is deterministically upper bounded by $3$. However, if the metrics in $\mathcal{Y}$ can be “scaled up” and stay in $\mathcal{Y}$ (such as in HSTs), we can obtain the proper definition by scaling the expansion up by an extra factor of $3$.
5 Nested and outlier embeddings for HSTs
In this section, we will develop an algorithm for constructing probabilistic nested embeddings when the host class of metrics is hierarchically separated trees (HSTs). We will then use this construction to develop an algorithm for bicriteria approximations of outlier embeddings into HSTs, which is the main result of this paper.
5.1 Compositions of nested embeddings for HSTs
Recall that in Section 4 we developed an algorithm for constructing nested embeddings that requires a subroutine for a perfect merge of embeddings into the host metric. We begin by developing such a subroutine for HSTs in Algorithm 3. The algorithm takes in two $\beta$-HSTs for a fixed ratio $\beta$ and returns another $\beta$-HST.
Algorithm 3 $\texttt{MergeHST}(\alpha_{1},\alpha_{2})$ Algorithm for merging two $\beta$-HSTs
Input: Embeddings $\alpha_{1}:Z_{1}\rightarrow T_{1}$, $\alpha_{2}:Z_{2}\rightarrow T_{2}$ for $\beta$-HSTs $T_{1},T_{2}$, where $|Z_{1}\cap Z_{2}|=1$. We assume without loss of generality that the height of $T_{1}$ is at least as large as the height of $T_{2}$.
Output: An embedding of $Z_{1}\cup Z_{2}$ into a single $\beta$-HST
1:Let $u=Z_{1}\cap Z_{2}$
2:Create a new HST $T$ that is initially a copy of $T_{1}$
[…]
4:for $i=0$ to the height of $T_{1}$ do
5:  for the subtree of each child of $\alpha_{2}(u)$’s ancestor at height $i$ in $T_{2}$ do
6:   if the subtree does not contain $\alpha_{2}(u)$ then $\triangleright$ That is, if the subtree is one we have not yet merged into $T_{1}$
7:     Make a copy of the subtree and make it a child of $\alpha_{1}(u)$’s ancestor at level $i$ in $T$
8:     for any $x\in Z_{2}\setminus\set{u}$ such that $\alpha_{2}(x)$ is in the subtree at this step do
9:      Set $\alpha(x)$ to be the copy of $\alpha_{2}(x)$ created in $T$ at this step
10:Return $\alpha$
Figure 3: This figure depicts what happens in the algorithm MergeHST. In particular, $u$ is the node common to both HSTs on the left side of the figure. The triangles represent subtrees, and the letters below the triangles indicate the presence of the corresponding node as a leaf of that subtree. Note that after merging trees, the least common ancestor of $x$ and $y$ is that of $x$ and $u$ from the first tree, as this was higher than that of $y$ and $u$. Further, the least common ancestor of $x^{\prime}$ and $y^{\prime}$ is at the height of the least common ancestor of $u$ and $y^{\prime}$ in the second tree, as that is higher than the least common ancestor of $u$ and $x^{\prime}$ in the first tree. This is the foundational idea behind why $d_{\alpha}(x,y)\geq\max\set{d_{\alpha_{1}}(u,x),d_{\alpha_{2}}(u,y)}$ holds when $\alpha$ is the output embedding for MergeHST on inputs $\alpha_{1}$ and $\alpha_{2}$.
We show in Lemma 5.1 that MergeHST is perfect. Figure 3 illustrates how the merge function works. Given two HSTs that share a common leaf vertex, the function aligns the two HSTs along the path from the shared leaf to the respective roots, and merges the intermediate vertices along the way. By the manner in which we have defined $\beta$-HSTs, the lengths of the edges along the common paths align correctly with each other, enabling the guarantees we require.
Lemma 5.1:
Algorithm 3 is a perfect merge function for embeddings into the space of $\beta$-HSTs.
Proof.
Let $x\in Z_{1}$ and $y\in Z_{2}$ with $Z_{1}\cap Z_{2}=\set{u}$ and let $\alpha_{1}:Z_{1}\rightarrow T_{1}$, $\alpha_{2}:Z_{2}\rightarrow T_{2}$, where both $T_{1}$ and $T_{2}$ are $\beta$-HSTs with elements of $Z_{1}$ and $Z_{2}$ mapped to their leaves respectively. Let $\alpha\leftarrow\texttt{MergeHST}(\alpha_{1},\alpha_{2})$ be the resulting $\beta$-HST embedding from $Z_{1}\cup Z_{2}$ into $T$.
We first argue that $T$ is indeed a $\beta$-HST. Consider a node at height $i$ in $T$. If this node was originally in $T_{1}$, then any subtree that may have been added as a subtree of this node has height exactly i1i-1 (by the fact that they are copies of subtrees of the children of $\alpha_{2}(u)$’s ancestor at height $i$), so this node is still at height $i$ and is still labeled properly. Now consider the case that this is a copy of a node from $T_{2}$. If the original was at height $i$ in $T_{2}$, then this copy is also at height $i$, as its subtree is identical to that in $T_{2}$. Therefore, the node still retains the correct label.
We will show each of the perfect merge criteria separately. Let $LCA(x,y)$ be the least common ancestor of $\alpha(x)$ and $\alpha(y)$ in the final merged tree $T$.
  • First consider the case that $x,y\in Z_{1}$. Note that $T$ is identical to $T_{1}$ except with some extra children added to some internal nodes, and $\alpha$ maps $x$ and $y$ to exactly the same leaves as under $\alpha_{1}$, so their distance is unchanged.
  • Now consider $x,y\in Z_{2}$. Let $i^{*}$ be the height of the least common ancestor of $\alpha_{2}(x)$ and $\alpha_{2}(y)$ in $T_{2}$. Let $S_{i}(x)$ be the subtree of $T_{2}$ of height $i$ that contains $x$. Note that by definition of $i^{*}$, $S_{i^{*}-1}(x)\neq S_{i^{*}-1}(y)$ (which also holds for all $i<i^{*}$).
    First, consider the case that $S_{i^{*}-1}(x)\neq S_{i^{*}-1}(u)$ and $S_{i^{*}-1}(y)\neq S_{i^{*}-1}(u)$. Then the root of $S_{i^{*}-1}(x)$ and root of $S_{i^{*}-1}(y)$ are both made children of $\alpha(u)$’s ancestor at height $i^{*}$. Thus, their least common ancestor is still at level $i^{*}$ and their distance is unchanged.
    Next, consider the case that $S_{i^{*}-1}(x)=S_{i^{*}-1}(u)$ and $S_{i^{*}-1}(y)\neq S_{i^{*}-1}(u)$. (Note that at most one can be equal to $S_{i^{*}-1}(u)$ so we may assume this without loss of generality.) The first ancestor $\alpha(x)$ and $\alpha(y)$ could have in common is $\alpha(u)$’s ancestor at level $i^{*}$, as $\alpha(x)$’s first ancestor from $T_{1}$ is below this level (and thus all ancestors above this level are also from $T_{1}$), and $\alpha(y)$’s lowest ancestor in common with $u$ is at level $i^{*}$. Further, they do have this ancestor in common by construction, so the least common ancestor height under $\alpha$ is the same as under $\alpha_{2}$. Thus, we get $d_{\alpha}(x,y)=d_{\alpha_{2}}(x,y)$.
  • Finally consider the case that $x\in Z_{1}\setminus\set{u}$ and $y\in Z_{2}\setminus\set{u}$. Let $LCA(a,b)$ be the least common ancestor of $\alpha(a)$ and $\alpha(b)$ in $T$. Note that by definition of HSTs, for any set of nodes $\set{a,b,c,d}$, we have $d_{\alpha}(a,b)\geq d_{\alpha}(c,d)$ if and only if $LCA(a,b)$ is at or above the height of $LCA(c,d)$.
    We will show $d_{\alpha}(x,y)\geq d_{\alpha}(x,u)=d_{\alpha_{1}}(x,u)$ and $d_{\alpha}(x,y)\geq d_{\alpha}(y,u)=d_{\alpha_{2}}(y,u)$ separately.
    • $d_{\alpha}(x,y)\geq d_{\alpha}(x,u)$: Note that $LCA(x,y)$ is an original member of $T_{1}$, as nodes in $T_{1}$ only have ancestors from $T_{1}$. Further, by construction all ancestors of $\alpha(y)$ that are from $T_{1}$ are also ancestors of $\alpha(u)$. Thus, $LCA(x,y)$ must also be an ancestor of $\alpha(u)$ and it must be at or above the height of $LCA(x,u)$.
    • $d_{\alpha}(x,y)\geq d_{\alpha}(y,u)$: Note that $LCA(y,u)$ is $\alpha(y)$’s lowest height ancestor that was an original member of $T_{1}$, as all ancestors below that height are from a copied subtree from $T_{2}$. Further, since $\alpha(x)$ only has ancestors that were original members of $T_{1}$, $LCA(x,y)$ must be at or above the height of LCA(y,u)LCA(y,u).∎
We now combine all of the components of a nested composition to get our main result for this section. Using the following HST embedding result of [28] allows us to bound the $c_{X}$ term in the final distortion.
Theorem 5.2 (Theorem 1 from [28]):
For any finite metric of size $n$, there exists an efficient algorithm that samples from a probabilistic embedding $\mathcal{D}$ of that metric into HSTs, with distortion at most $96\ln n$.
This implies the existence of an efficiently sample-able embedding $\mathcal{D}_{K}\xhookrightarrow{O(\log|K|)}\mathcal{H}$ for any subset $K$ of a given metric. Thus, Theorem 5.2, Theorem 4.2, and Lemma 5.1 combine to show that probabilistic embeddings into HSTs with distortion $c_{S}$ can be extended to nested embeddings with expansion at most $4$ on pairs of nodes in $S$ and expansion at most $\zeta\cdot(c_{S}\log k+\log^{2}k)$ on all pairs of nodes, for some constant $\zeta$, as stated in Theorem 2.4.101010We use $4$ instead of $3$ for the rescaling, since HSTs as we have defined them here require distances to be a power of $2$.
5.2 Outlier embeddings for HSTs
Now, given the existence of nested embeddings into HSTs, we turn to approximating outlier embeddings into HSTs. In [20], this type of approximation is obtained for $\ell_{2}$ by using a semidefinite program for finding the minimum distortion $\ell_{2}$ embedding and introducing new “outlier variables” and rounding. We will employ a similar technique here using the linear program and associated rounding algorithm of [39] to approximate the optimal distortion for probabilistic HST embeddings.
Below we present our modified version of this LP with the changes highlighted in red. Let $\Delta$ be the diameter of $(X,d)$ and let $M=\set{2^{0},2,2^{2},\ldots,2^{\lceil\log\Delta\rceil}}$. Let $\zeta$ be the constant factor from Theorem 2.4 and let $B_{i}(r):=\set{j\in X:d(i,j)\leq r}$. This LP is parameterized by inputs $((X,d),k,c)$.
$\min{\color[rgb]{0.65,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.87}{0.35}\pgfsys@color@cmyk@fill{0}{0.85}{0.87}{0.35}\sum_{i=1}^{n}\delta_{i}}$ (HST LP)
$\forall j,j^{\prime}$ $\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}\leq{\color[rgb]{0.65,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.87}{0.35}\pgfsys@color@cmyk@fill{0}{0.85}{0.87}{0.35}(4+\zeta\cdot(\delta_{j}+\delta_{j^{\prime}})\cdot\log^{2}k)}\cdot c\cdot d(j,j^{\prime})$ (4)
$\forall i,r,(j,j^{\prime}\in B_{i}(r))\ \ \ \ \$$\min(x_{i,j}^{r},x_{ij^{\prime}}^{r})\geq z_{ijj^{\prime}}^{r}$ (5)
$\forall j,j^{\prime},i,r$ $\sum_{i:j,j^{\prime}\in B_{i}(r)}z_{ijj^{\prime}}^{r}\geq 1-\gamma_{jj^{\prime}}^{r}$ (6)
$\forall j,r$ $\sum_{i:j\in B_{i}(r)}x_{ij}^{r}=1$ (7)
$\forall j,j^{\prime},(r<d(j,j^{\prime}))\ \ \ \ \$$\gamma_{jj^{\prime}}^{r}=1$ (8)
$\forall j,j^{\prime},r\ \ \ \$$\gamma_{jj^{\prime}}^{r}\leq 1$ (9)
$\forall i,j,j^{\prime},r\ \ \ \ \$$\gamma_{jj^{\prime}}^{r}x_{ij}^{r},z_{ijj^{\prime}}^{r},{\color[rgb]{0.65,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.87}{0.35}\pgfsys@color@cmyk@fill{0}{0.85}{0.87}{0.35}\delta_{i}}\geq 0;{\color[rgb]{0.65,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.65,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.87}{0.35}\pgfsys@color@cmyk@fill{0}{0.85}{0.87}{0.35}\delta_{i}\leq 1}$ (10)
Note that the original version of this LP in [39] set $c$ to be a variable and minimized its value as the objective function. Here, $c$ is a parameter that is a part of our input, rather than a variable. We have introduced new variables $\delta_{i}$ for each $i\in X$ that we use as an indicator for whether or not $i$ is an outlier in the embedding. Intuitively, we imagine assigning a “representative” at each node of the HST such that the representative of a leaf node is the leaf node itself, and the representative of any other node is the same as one of its children’s representatives. Then we can think of $z_{ijj^{\prime}}^{r}$ as the probability that $i$ is the “representative” of $j$ and $j^{\prime}$ at level $r$, $\gamma_{jj^{\prime}}^{r}$ as the probability that $j$ and $j^{\prime}$ are separated at level $r$, and $x_{ij}^{r}$ as the probability that $i$ is the representative of $j$ at level $r$. We will discuss this further when we bound the optimal objective value for this LP.
In [39], constraint 4 is written as $\forall j,j^{\prime},\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}\leq c\cdot d(j,j^{\prime})$ for a variable $c$. Intuitively, the left hand side of the constraint represents the expected distance between $j$ and $j^{\prime}$ that is allowed in the embedding represented by this LP. We have modified this right hand side by setting $c$ to be a parameter of the problem rather than a variable, and by allowing an extra factor of $4$ for the distortion of all pairs of nodes, as well as an additive $O((\delta_{j}+\delta_{j^{\prime}})c\log^{2}k)$ factor. (Note that $k$ is technically unknown, but because it comes from a small range, we can run this program with different values of $k$ until we reach the smallest $k^{\prime}$ that yields an optimal LP solution with value at most $k^{\prime}$, which we will show is no more than the optimal value for $k$. See Algorithm 4.) Further, we will use the rounding algorithm given by [39] without changes. Note that this algorithm will generate some embedding on the entire set $(X,d)$. To account for this, we will select some threshold $\delta^{*}$ and remove all nodes $i$ with $\delta_{i}\geq\delta^{*}$ from the embedding. We argue that the resulting embedding has no contraction and only bounded expansion, and the number of outliers generated is bounded by $k/\delta^{*}$.
Next, we briefly outline the proof techniques of [39] and how they apply to our result. For completeness, the specific rounding algorithm for the embedding on $X$ will appear in Appendix C, but it is unaltered from that given in [39].
5.2.1 Bounding the optimal LP solution
We first need to argue that the optimal solution to this LP has value at most $k$.
Lemma 5.3:
The optimal solution to HST LP has value at most $k$.
This lemma is true for the same reasons as given in [39]. [^11] Note that in order to obtain Lemma 5.3, we need to have that the final embedding has no contraction, even on outliers, as we need to be able to guarantee that any node $j$ is represented at level $r$ by a node $i$ in $B_{j}(r)$. If we allow contraction for outliers, this may not be the case, and $j$ may be represented at each level by something that is actually quite far away from it. The primary change we have made is that $c$ is now a constant, and in finding a feasible solution to the LP, we need to assign values to the $\delta_{i}$ in addition to assigning the other variables as in [39]. Note that if we assign $\delta_{i}$ to be $1$ if $i$ is an outlier, then the right hand side of constraint 4 is an upper bound on the distortion for any pair of nodes under the nested composition of embeddings by Theorem 2.4. Note in particular that constraints 6 through 8 require that the underlying embedding be expanding on all pairs, even when one or both nodes may be an outlier. Thus, bounding the optimal solution to this LP is why it was essential to find nested compositions of embeddings for HSTs, rather than just Lipschitz extensions.
Proof of Lemma 5.3.
Let $\mathcal{D}_{S}$ be a probabilistic embedding on some $S\subseteq X$ into HSTs with distortion $c$ such that $K:=|X\setminus S|$ has size at most $k$. By Theorem 2.4, there exists a probabilistic embedding $\mathcal{D}$ that is non-contracting, has distortion at most $4c$ on pairs of nodes in $S$, and has overall distortion at most $\zeta\cdot(c\log k+\log^{2}k)$ for some constant $\zeta$. Set the variables $z_{ijj^{\prime}}^{r},\gamma_{jj^{\prime}}^{r},$ and $x_{ij}^{r}$ as in [39]. (That is, $z_{ijj^{\prime}}^{r}$ is the probability that $i$ is the “representative” of $j$ and $j^{\prime}$ at level $r$, $\gamma_{jj^{\prime}}^{r}$ is the probability that $j$ and $j^{\prime}$ are separated at level $r$, and $x_{ij}^{r}$ is the probability that $i$ is the representative of $j$ at level $r$.) Additionally, define $\delta_{i}$ be $1$ if $i$ is an outlier and 0 otherwise. Clearly this solution has value $k$, so we need only show it satisfies the constraints.
Constraints 5 through 10 are satisfied for identical reasoning to that given in [39]. Constraint 4 is also satisfied for the same reasons as the equivalent constraint in [39], but we briefly justify it here as well since we changed the constraint. We defer justification of the other constraints’ satisfaction to Appendix C.
To see why constraint 4 is satisfied, let $\mu_{jj^{\prime}}^{r}$ be the probability that $j$ and $j^{\prime}$ are a distance exactly $r$ apart in the embedding sampled from $\mathcal{D}$. Note that if $j$ and $j^{\prime}$ are separated at level $r$, then their distance in that embedding is at least $r^{*}>r$. Thus, $\gamma_{jj^{\prime}}^{r}$ is the probability that the embedding drawn from the distribution places $j$ and $j^{\prime}$ at a distance strictly greater than $r$ apart, which means $\gamma_{jj^{\prime}}^{r}=\sum_{r^{*}\in M:r^{*}>r}\mu_{jj^{\prime}}^{r^{*}}$. Thus we get
$\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}$ $=\sum_{r\in M}r\cdot\sum_{r^{*}\in M:r^{*}>r}\mu_{jj^{\prime}}^{r^{*}}$
$=\sum_{r^{*}\in M}\mu_{jj^{\prime}}^{r^{*}}\sum_{r\in M:r^{*}>r}r$
$\leq\sum_{r^{*}\in M}r^{*}\mu_{jj^{\prime}}^{r^{*}}$
$=\mathop{E}_{\alpha\sim\mathcal{D}}\left[d_{\alpha}(j,j^{\prime})\right].$
The first line is by what we argued above; the second is by reordering the sums; the third is by the fact that $\sum_{r\in M:r<r^{*}}r\leq r^{*}$ (as all values in $M$ are powers of $2$); the fourth line is by definition of $\mu_{jj^{\prime}}^{r}$ as the probability that $j$ and $j^{\prime}$ are a distance exactly $r$ apart. Note that if at least one of $j,j^{\prime}$ is an outlier, then at least one of $\delta_{j}$ and $\delta_{j^{\prime}}$ is $1$, so the right hand side of constraint 4 is at least $\zeta\cdot c\cdot\log^{2}k\cdot d(j,j^{\prime})$, which is an upper bound on the expected distance between $j$ and $j^{\prime}$ under $\mathcal{D}$. If neither $j$ nor $j^{\prime}$ is an outlier, then the right hand side of constraint 4 is $4c\cdot d(j,j^{\prime})$, which is an upper bound on their expected distance under $\mathcal{D}$. ∎
5.2.2 Rounding the LP
We formally describe a rounding algorithm for the LP in Algorithm 4. This algorithm largely relies on that of [39], but it takes the extra step of excluding from the domain of the output embedding all nodes in $X$ with optimal $\delta_{j}$ value above some threshold $\delta^{*}$. These excluded nodes form the outlier set for this embedding. Note that the outlier set is deterministic, as it is decided by a simple thresholding of the deterministic LP solution.
Algorithm 4 Algorithm for approximating outlier sets for HST embeddings
Input: Metric space $(X,d)$, target distortion $c$, approximation factor $\epsilon>0$
Output: An embedding $\alpha:S\rightarrow T$ for some HST $T$ and some $S\subseteq X$
1:for $k=1$ to $n$ do
2:  Let $LP_{k}$ be HST LP for $(X,d),c,k$ $\triangleright$ For weighted outlier sets, use $\min\sum_{i\in X}w_{i}\delta_{i}$ as the objective
3:  Find an optimal solution to $LP_{k}$ of value $v_{k}$
4:  Use the rounding algorithm of [39] (Algorithms 6 and 7) to sample an embedding $\alpha_{k}:X\rightarrow T$ for HST $T$
5:  $\delta^{*}\leftarrow\frac{\epsilon}{16\zeta\cdot\log^{2}k}$
6:  $K_{k}\leftarrow\set{j\in X:\delta_{j}\geq\delta^{*}\text{ in the optimal LP solution}}$
7:Let $k^{*}$ be the smallest value of $k$ with $v_{k}\leq k$ $\triangleright$ For weighted outlier sets, let $k^{*}$ minimize $v_{k}\log^{2}k$
8:Return $\alpha_{k^{*}}$ and $K_{k^{*}}$
In order to show Algorithm 4 does in fact sample from a probabilistic $(O(\frac{k}{\epsilon}\log^{2}k),(32+\epsilon)c)$-outlier embedding, we will reference the following lemmas from [39].
Lemma 5.4 (counterpart of Lemma 10 from [39]):
The embedding produced by the rounding Algorithms 6 and 7 is non-contracting.
Lemma 5.5 (counterpart of Lemma 11 from [39]):
If $\delta_{j}$ is the optimal LP value of the variable $\delta_{j}$ and if $\alpha$ is sampled by Algorithms 6 and 7, we have $\mathop{E}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8\cdot(4+\zeta\cdot\log^{2}k\cdot(\delta_{j}+\delta_{j^{\prime}}))\cdot c\cdot d(j,j^{\prime})$
Note that Lemma 5.4 can be proven identically to in [39], as the relevant algorithms and constraints involved in that proof are unchanged in our version of that LP, though we will briefly discuss this in Appendix C.
Proof outline of Lemma 5.5.
The original lemma of [39] stated that $\mathop{E}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8q$, where $q$ is the optimal LP solution value for their distortion variable. (Recall that for us, the distortion variable is replaced with a constant and we focus on our new outlier variables.) To prove this, they used only constraints 5 through 10 to show that $\mathop{E}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8\cdot\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}$ and then applied their version of constraint 4 to get the final result. As Constraints 5 through 10 and the rounding algorithm remain unchanged in our algorithm, the statement $\mathop{E}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8\cdot\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}$ still holds. Thus, by our slightly modified version of Constraint 4, we get $\mathop{E}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8\cdot\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}\leq 8\cdot(4+\zeta\cdot\log^{2}k\cdot(\delta_{j}+\delta_{j^{\prime}}))\cdot c\cdot d(j,j^{\prime})$, as stated in the lemma. ∎
Now we can prove Theorem 2.5, which says that we can efficiently sample from an $(O(\frac{k}{\epsilon}\log^{2}k),(32+\epsilon)c)$ embedding.
Proof of Theorem 2.5.
We know from Lemma 5.4 that the sampled embedding from Algorithm 4 is non-contracting. To consider its expected expansion, we need only consider pairs of nodes $j,j^{\prime}$ with $\delta_{j},\delta_{j^{\prime}}\leq\delta^{*}$. Then by Lemma 5.5, the expected distance between such a $j$ and $j^{\prime}$ is at most
$(32+8\zeta\cdot\log^{2}k\cdot 2\delta^{*})\cdot c\cdot d(j,j^{\prime})\leq(32+\epsilon)\cdot c\cdot d(j,j^{\prime}),$
by the choice of $\delta^{*}$ in Algorithm 4.
Additionally, note that the set of outliers produced by Algorithm 4 is deterministic and has size at most $k/\delta^{*}=O(\frac{k}{\epsilon}\log^{2}k)$, as Lemma 5.1 tells us that the optimal solution to HST LP has value at most $k$. ∎
5.2.3 A note on the runtime
Notably, Algorithm 6 (Algorithm 1 from [39]) runs in expected polynomial time, not deterministic polynomial time. Following the same reasoning as [39], we can select some additional error term $\epsilon^{\prime}$ and a sufficiently large constant $c^{\prime}$ and terminate the while loop after at most $c^{\prime}\cdot n\ln\frac{\Delta}{\epsilon^{\prime}}$ steps. If the algorithm would have otherwise exceeded that number of steps, we just output an HST in which all distances are $2^{\lceil\log\Delta\rceil}$. Using sufficient concentration bounds, the rounding scheme incurs a distortion at most $\epsilon^{\prime}$ higher than in the original form. If we want the distortion of the final embedding to be at most $(32+\epsilon)c$, we can select $\epsilon/2$ for $\epsilon^{\prime}$ in this section and for the approximation factor in choosing a threshold $\delta^{*}$. Note that because of this choice, the runtime of the algorithm depends on $1/\epsilon$ and $\Delta$ (specifically growing with $\ln(\frac{\Delta}{\epsilon})$), unlike the $\ell_{2}$ outlier embedding algorithm in [20].
5.3 Weighted outlier sets
In the weighted outlier set problem, we are given $\mathcal{X}=(X,d)$ and $c$, as well as a set of weights $w$ on the nodes, and our goal is to find a subset $K\subseteq X$ such that $X\setminus K\xhookrightarrow{c}_{p}\mathcal{H}$ and $w(K)$ is minimized. So far we have considered the unweighted version of this problem, in which $w(i)=1$ for all $i\in X$. In this subsection, we will discuss the slight changes needed to apply the results of Section 5 to the weighted outlier setting and thus obtain a bicriteria approximation for this problem.
The weights of the nodes affect the cost of leaving those nodes out, but they have no effect on how good the resulting embedding may be. This means that if we attempt to modify HST LP for the weighted setting, the weights should only influence the objective function. To solve this problem, we will use Algorithm 4 with the changes noted in red for weighted outlier embeddings. In the weighted setting, we are not necessarily concerned with simply minimizing the outlier set size, so intuitively we should select the LP solution corresponding to the $k$ that minimizes the objective value function for that LP. However, in order to account for the extra $\log^{2}k$ approximation factor incurred by our algorithm, we will instead look for a value of $k$ that minimizes the objective value of its LP solution, times $\log^{2}k$. The necessary changes are noted as comments in Algorithm 4. Below we give the proof of Corollary 2.6.
Proof of Corollary 2.6.
Let $K$ be an outlier set of $\mathcal{X}$ that has size at most $\mathcal{K}(c)$, has weight at most $\text{OutlierCost}(c)$, and has $X\setminus K\xhookrightarrow{c}_{p}\mathcal{H}$. Define $k:=|K|$.
Let $k^{*}$ be the value chosen by the algorithm to minimize $v_{k^{\prime}}\log^{2}{k^{\prime}}$. Then
$v_{k^{*}}\cdot\log^{2}(k^{*})$ $\leq v_{k}\cdot\log^{2}k$
$\leq\text{OutlierCost}(c)\cdot\log^{2}\mathcal{K}(c),$
where the first line is by definition of $k^{*}$ and the second line is by the fact that $\delta_{i}=\begin{cases}1&i\in K\\ 0&else\end{cases}$ is part of a valid solution to HST LP with parameter $k$, and this solution has objective value at most $\text{OutlierCost}(c)$. (Note that this is part of a valid solution because only the objective function has changed, so all arguments with respect to feasibility of such a solution still go through.)
Now we consider the rounding step of the algorithm. We know that by selecting all $i$ with $\delta_{i}\geq\frac{\epsilon}{\zeta\cdot\log^{2}k^{*}}$ in the LP solution corresponding to $k^{*}$, we obtain a solution with at most $(32+\epsilon)c$ distortion on the non-outliers (again following identical reasoning to that in the unweighted case). Thus, we are just left with bounding the weight of the outlier set. We have that the total cost is at most $v_{k^{*}}\cdot\frac{\zeta\cdot\log^{2}(k^{*})}{\epsilon}$ as a result of our rounding, as $v_{k^{*}}$ is the value of the objective function under this solution. This implies an outlier set of weight at most $O(\frac{1}{\epsilon}\cdot\text{OutlierCost}(c)\cdot\log^{2}\mathcal{K}(c))$. ∎
Appendix A More information on applications
A.1 Bad metrics
Proof of Claim 3.2.
Let $\mathcal{D}_{M}:M\xhookrightarrow{c}_{p}\mathcal{H}$ be a probabilistic embedding of $(M,d_{M})$ into HSTs and let $\mathcal{D}_{x}:N_{x}\xhookrightarrow{c}_{p}\mathcal{H}$ be a set of probabilistic embeddings of the $N_{x}$ into HSTs. Let $\mathcal{D}:M_{\beta}[\mathcal{N}]\rightarrow\mathcal{H}$ be the probabilistic embedding whose distribution is defined by drawing $\alpha_{x}$ from each $\mathcal{D}_{x}$ and $\alpha_{M}$ from $\mathcal{D}_{M}$ and outputting the result of Algorithm 5 (from this section) on those embeddings. We will show that $\mathcal{D}$ is non-contracting and $\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}((x,u),(y,v))]=O(c)\cdot d_{M^{*}}((x,u),(y,v))$, where $M^{*}=M_{\beta}[\mathcal{N}]$. Intuitively, Algorithm 5 just applies the MergeHST algorithm from Section 5 to compose the input metrics. In order to create a “merge point” for that algorithm, we add node $x\in M$ to its corresponding set $N_{x}$ and we just set its distance from everything to be the maximum diameter of any $N_{x^{\prime}}$. This will ensure the triangle inequality is still satisfied on all pairs, and it makes it easy to adjust the embedding on $N_{x}$ to a new embedding of distortion at most $c$.
Algorithm 5 Metric HST composition: Algorithm for finding an HST embedding for a composition of metrics
Input: A metric $(M,d_{M})$ and set of metrics $\mathcal{N}=\set{(N_{x},d_{N_{x}})}_{x\in M}$, HST embeddings $\alpha_{M}$ and $\set{\alpha_{x}}$ on all of the input metrics, parameter $\beta\geq 1/2$; the labels of nodes in $M$ should be disjoint from the labels of nodes in $N_{x}$ and the minimum distance in $M$ should be exactly 11
Output: An embedding of $M_{\beta}[\mathcal{N}]$ into HSTs
1:Let $\Delta$ be the maximum diameter of any $N_{x}$ $\triangleright$ We assume wlog that the height of all HSTs mapped to by some $\alpha_{x}$ is $\lceil\log c\Delta\rceil$}
2:for $x\in M$ do
3:  $N_{x}^{\prime}\leftarrow N_{x}\cup\set{x}$
4:  For $y,z\in N_{x}$, $d_{x}^{\prime}(y,z):=d_{N_{x}}(y,z)$
5:  For $y\in N_{x}$, $d_{x}^{\prime}(x,y):=\Delta$
6:  Let $\alpha_{x}^{\prime}:N_{x}^{\prime}\rightarrow\mathcal{H}$ be the same as $\alpha_{x}$ but place $x$ at maximum distance from all other nodes
7:  That is, add a new path from the root to the leaf level, and map $x$ to the leaf of that path
8:Initially let $\alpha$ be a $2^{\lceil\log\beta\Delta\rceil}$ rescaling of $\alpha_{M}$, but rename each $x\in M$ to $(x,x)$
9:for $x\in M$ do
10:  $\alpha\leftarrow\texttt{MergeHST}(\alpha,\alpha_{x}^{\prime})$
11:Return $\alpha|_{M_{\beta}[\mathcal{N}]\setminus\set{(x,x):x\in M}}$
Let $N_{x}^{\prime}$ and $d_{x}^{\prime}$ be as in lines 3-5 of Algorithm 5 and note that $d_{x}^{\prime}$ satisfies the triangle inequality on all pairs in $N_{x}^{\prime}$, so it does define a metric. Our assumption on the height of the embeddings $\alpha_{x}$ is also valid, as the maximum embedding distance is at most $c\Delta$ for any such embedding, and we can always make an HST taller by adding a parent for the root. Further, because of our height assumption, we get that for all $u\in N_{x}$ $\Delta\leq d_{\alpha_{x}}(x,u)=2^{\lceil\log c\Delta\rceil}\leq 2c\Delta=2c\cdot d_{x}^{\prime}(x,u)$.
Consider a pair of nodes $(x,u),(y,v)\in M^{*}$ (where $u\in N_{x}$ and $v\in N_{y}$), and we will bound their expected distance under $\mathcal{D}$. We can break this into the cases where $x$ does or does not equal $y$.
  • First consider the case that $x=y$. We have $d_{M^{*}}((x,u),(y,v))=d_{N_{x}}(u,v)$. Further, the perfect merge properties of MergeHST ensure that for the final $\alpha$, $d_{\alpha}((x,u),(y,v))=d_{\alpha_{x}}(u,v)$. Thus, the expected distortion of between $u$ and $v$ under $\mathcal{D}_{N_{x}}$ is also the expected distortion between $u$ and $v$ under $\mathcal{D}$.
  • Now consider $(x,u),(y,v)$ for $x\neq y$. We have $d_{M^{*}}((x,u),(y,v))=\beta\Delta d_{M}(x,y)$, and we know
    $\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}((x,u),(x,v))]$ $\leq\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}((x,u),(x,x))+d_{\alpha}((x,x),(y,y))+d_{\alpha}((y,y),(y,v))]$
    $\leq 4c\cdot\Delta+\beta\Delta\mathop{E}_{\alpha_{M}\sim\mathcal{D}_{M}}[d_{\alpha_{M}}(x,y)]$
    $\leq 4c\cdot\Delta+c\cdot\beta\Delta d_{M}(x,y)$
    $\leq 5c\cdot\beta\Delta d_{M}(x,y),$
    where the first inequality is by the triangle inequality; the second is by linearity of expectation, the fact that $N_{x}$ has diameter at most $\Delta$, and the perfect merge properties of MergeHST; the third is by definition of $\mathcal{D}_{M}$, and the fourth is by the fact that $d_{M}(x,y)\geq 1$.
Thus, we are left with showing at all embeddings in the distribution are non-contracting. We again break this into cases.
For any pair of nodes $(x,u),(x,v)$, non-contraction comes automatically from non-contraction of embeddings in $\mathcal{D}_{x}$ combined with the perfect merge properties of MergeHST.
Now consider a pair of nodes $(x,u),(y,v)$ with $x\neq y$, and assume wlog that $x$ preceded $y$ in the ordering used by Algorithm 5. We will show that once $x$ has been processed, $d_{\alpha}((x,u),(y,y))\geq\Delta\beta d_{M}(x,y)$ for all $u\in N_{x}$ and all $y\in M$, and from there we will be able to show the desired result.
Consider when $x$ is first processed. Prior to the merge, $d_{\alpha}((x,x),(y,y))\geq\Delta\beta d_{M}(x,y)$. The perfect merge properties of MergeHST ensure that $d_{\alpha}((x,u),(y,y))\geq d_{\alpha^{\prime}}((x,x),(y,y))\geq\Delta\beta d_{M}(x,y)$, where $\alpha^{\prime}$ is the embedding prior to this merge.
Now consider when the embedding $\alpha_{y}^{\prime}$ is merged into the overall embedding. The perfect merge properties of MergeHST tell us that if $\alpha^{\prime}$ is the overall embedding before this merge step, then $d_{\alpha}((x,u),(y,v))\geq d_{\alpha^{\prime}}((y,y),(x,u))\geq\beta\Delta d_{M}(x,y)$, as we showed earlier. ∎
A.2 Buy-at-Bulk
In the proofs of Theorems 3.3 and 3.4, we will show that the given bound holds for all choices of $c_{i}$ considered by the algorithm. Here we briefly justify why this is sufficient to prove the overall statements, which minimize over all $c$ not just the $c_{i}$ considered by the algorithm. First, for all $c\geq\eta\log n=c_{0}$, $\text{OutlierCost}(c)=0$, so we need only justify this for $c<\eta\log n$. Let $c$ be the value that minimizes $c\cdot OPT+\text{OutlierCost}(c)\log^{2}(n)$, and note that any $c_{i}>c$ has optimal outlier set value at most equal to that of $c$. As the algorithm will consider some value of $c_{i}$ that is bigger than $c$ and at most $\eta$ times its value (if $c\geq\eta/2$, it will in fact consider a choice of $c_{i}$ that is at most twice this value), we will end up with the desired approximation.
Proof of Theorem 3.3.
We show that if index $i$ is selected as the optimal solution by the algorithm, then the cost of the solution is at most $O(c_{i}\cdot OPT+\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i})))$. Let $K_{i}$ be the outlier set from line 8 of Algorithm 1.
Let $S_{1}$ be the set of pairs with neither endpoint in $K_{i}$ and let $S_{2}$ be the remaining pairs. We know that the cost of the solution we obtain for pairs in $S_{1}$ is at most $c_{i}\cdot OPT$, as the tree embedding incurs distortion $c_{i}$ and we get an optimal solution for the resulting tree.
Now consider the cost associated with our purchase of edges for pairs in $S_{2}$. The total cost of purchasing for each demand pair independently is at most $\sum_{i\in S_{2}}OPT_{i}$. (As the cost function is sub-additive, buying more capacity on a single edge never costs more than splitting that capacity into two parts and buying them separately.) We know that $\sum_{i\in S_{2}}OPT_{i}=O(\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i})))$ by Algorithm 4’s guarantees on the outlier set cost. ∎
A.3 Dial-A-Ride
Proof of Theorem 3.4.
We will show a slightly stronger version of the claim by showing that for each $i$, the cost associated with the solution for $c_{i}$ is bounded by $O(r_{n,m,\ell}\cdot c_{i}\cdot OPT+\text{OutlierCost}(c_{i})\log^{2}\mathcal{K}(c_{i})\cdot\Delta)$ and also by $O(r_{n,m,\ell}\cdot c_{i}+r_{2|R_{2}^{i}|,m,\ell}\log(|R_{2}^{i}|)\cdot OPT)$, where $R_{2}^{i}$ is the set of requests that intersect $K_{i}$ in iteration $i$ of the algorithm. (Only the first of these two bounds will be necessary to show the claim.) For the rest of this proof we will fix a choice of $i$ and consider the solution the algorithm obtains for that choice of $i$ and we refer to $R_{2}^{i}$ as $R_{2}$.
Let $K_{i}$ be the outlier set from line 8 of Algorithm 1, let $R_{1}$ be the set of requests $(s_{j},t_{j})$ that contain a point in $K_{i}$, and recall that $R_{2}$ is the set of remaining requests. Note that each solution the algorithm finds is feasible, as we handle all requests in $R_{1}$ according to some feasible solution given by $\mathcal{A}$, and then we handle all requests in $R_{2}$ according to another feasible solution from $\mathcal{A}$ or by a naive algorithm for this set of requests.
We will divide the cost of the algorithm’s solution into three sections: the cost of servicing requests in $R_{1}$, the cost of servicing requests in $R_{2}$, and the cost of moving the van from the last drop-off location in the $R_{1}$ solution to the first pickup location in the $R_{2}$ solution.
  1. 1.
    First consider the cost of the solution we obtain for servicing requests in $R_{1}$. Our expected cost is at most $r_{n,m,\ell}$ times the optimum solution on the tree this solution is embedded in, and the expected distortion of that embedding is bounded by $2c_{i}$ (as we potentially lose a factor of $2$ by setting $\epsilon=1$). Thus, the total cost of this part of the solution is bounded by $2r_{n,m,\ell}\cdot c_{i}\cdot OPT$ (noting that the optimal solution for $R_{1}$ is no more than the optimal solution for $R$).
  2. 2.
    Now consider the cost of the second part of the solution. From the analysis of the weighted version of Algorithm 4, we know that the total cost of the outlier set is at most $O(\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i}))$ This means
    $\sum_{i\in R_{2}}d(s_{i},t_{i})$ $\leq\sum_{x\in K_{i^{*}}}\sum_{i\in R_{x}}d(s_{i},t_{i})$
    $\leq\sum_{x\in K_{i^{*}}}w_{x}$
    $\leq O(\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i})).$
    The first line comes immediately from the definition of $R_{x}$ (the set of requests involving $x$) and $R_{2}$; the second comes from the definition of $w_{x}$; and the third comes from our bound on the outlier set value of Algorithm 4.
    Consider a naive algorithm that simply picks an arbitrary order on the requests and services the pick-ups and drop-offs one at a time in that order. The cost of the $j$th pickup and drop-off is $d(s_{j},t_{j})$, and the cost of moving from $t_{j}$ to $s_{j+1}$ is $d(t_{j},s_{j+1})\leq\Delta$. In total we service at most $\text{OutlierCost}(c_{i})$ requests in $R_{2}$, as we assume each request has a cost at least $1$. Thus, the total cost incurred on this part of the solution is bounded by $O(\text{OutlierCost}(c_{i})\log^{2}\mathcal{K}(c_{i})+\text{OutlierCost}(c_{i})\cdot\Delta)$.
    Additionally, we know that the optimal cost on this set of requests is bounded by $OPT$, and the number of points we need to embed to get a solution is at most $2|R_{2}|$. Thus, we can use a standard Fakcharoenphol et al. [28] embedding on the set of points involved in this request and we get a solution of value at most $O(r_{2|R_{2}|,m,\ell}\log|R_{2}|OPT)$.
  3. 3.
    Finally, consider the cost of the algorithm moving from the last drop-off spot $t_{j}$ in $R_{1}$ to the first pickup spot $s_{j^{\prime}}$ in $R_{2}$. This total cost is $d(t_{j},s_{j^{\prime}})$, but this is upper bounded by $OPT$, as at some point any optimal algorithm must visit both locations. ∎
A.4 Minimum communication cost tree
In this section, we discuss a linear program-based approximation algorithm for the outlier version of the MCCT problem on trees. [^12] An intuitive approach here might be to greedily remove the node $i$ maximizing $\sum_{j}d(i,j)$, but this does not get a finite approximation ratio, even for trees. As we have already argued in Section 3.4, this can combine this with our outlier approximation algorithm to get a wide variety of potential approximations whose quality depends on the structure of the underlying metric. We will expand on that fact in this section.
Let $(T,d)$ be a tree metric, and let $k$ be an integer. Then we define variables $x_{ij}$ for each pair of nodes $ij$ and variables $\delta_{i}$ for each variable $i$. We can use them to make the following linear program.
$\min\sum_{i,j\in T}x_{ij}\cdot r_{ij}\cdot d(i,j)\quad s.t.$ (MCCT Outlier LP)
$\forall i,j\in T:\quad$ $\delta_{i}+\delta_{j}+x_{ij}\geq 1$
$\sum_{i\in T}\delta_{i}\leq k$
$\forall i,j\in T:\quad$ $\delta_{i},x_{ij}\in[0,1]$
Claim A.1:
Let $(T,d)$ be a tree metric such that there exists a non-contracting embedding $\alpha:T\rightarrow\mathcal{H}$ with a set $K\subseteq T$, $|K|\leq k$, and $\sum_{i,j\in T\setminus K}r_{ij}\cdot d_{\alpha}(i,j)\leq D$.
Then the following procedure finds an embedding $\alpha^{\prime}$ and subset $K^{\prime}\subseteq T$ with $|K^{\prime}|\leq 3k$ and $\sum_{i,j\in T\setminus K^{\prime}}d_{\alpha^{\prime}}(i,j)\leq 24D$.
  1. 1.
    Solve MCCT Outlier LP and let $x_{ij},\delta_{j}$ be the resulting values
  2. 2.
    Let $K^{\prime}=\set{i:\delta_{i}\geq 1/3}$
  3. 3.
    Embed $T\setminus K^{\prime}$ into a tree $T^{\prime}$ with distortion at most $8$, using the algorithm of Gupta [31]
Proof.
First, MCCT Outlier LP has a solution of value at most $D$. In particular, let $K$ be as in the claim statement and consider setting $\delta_{i}=\begin{cases}1&i\in K\\ 0&else\end{cases}$ and $x_{ij}=\begin{cases}1&\delta_{i}=\delta_{j}=0\\ 0&else\end{cases}$. As there are only $k$ outliers, the second constraint is clearly satisfied, and we have defined $x_{ij}$ such that the other constraints are also satisfied. Further, $x_{ij}$ is 0 if either $i$ or $j$ is an outlier in $K$, so the objective function value is just $\sum_{i,j\in T\setminus K}r_{ij}d(i,j)$, which by the claim statement is bounded by $D$.
Let $K^{\prime}$ be as in the claim. As the $\delta_{i}$ sum to at most $k$, we know $|K^{\prime}|\leq 3k$. Consider any $x_{ij}$ with $i,j\notin K^{\prime}$ We have $x_{ij}\geq 1-\delta_{i}-\delta_{j}\geq 1/3$ by how we chose $K^{\prime}$. The cost associated with a given pair $ij$ of non-outliers in our solution is $r_{ij}d(i,j)$, but that same pair contributed at least $r_{ij}d(i,j)/3$ to the optimal objective function value, so our final cost is at most $3$ times the objective, which is at most $3D$. At this point, we have an isometric (distortion $1$) embedding of the metric induced by $T\setminus K^{\prime}$ into the original metric $T$ itself. Thus, we can view $K^{\prime}$ as a set of Steiner points in our original metric.
Finally we consider the loss incurred applying the algorithm of Gupta [31]. This creates a tree embedding that removes Steiner points and incurs distortion at most $8$, so the cost of this embedding on $T\setminus K^{\prime}$ is at most $24D$, as desired. This step is what prevents us from applying this algorithm directly to general metrics. ∎
Note additionally that if $(T,d)$ is an HST metric, then we can “remove” the bad points we’ve found by just deleting the leaves corresponding to those nodes, and as a result we do not incur the extra factor of $8$ on the distortion.
Converting to an approximation on general metrics
In Section 3.4, we argued that we can use the outlier embeddings algorithm from this paper to obtain a solution with at most $O((1+\epsilon)c_{k}(\mathcal{X})(OPT_{k}))$ cost and $O(\frac{k}{\epsilon}\log^{2}k)$ outliers, where $c_{k}(\mathcal{X})$ is the minimum distortion for embedding $\mathcal{X}$ into HSTs with at most $k$ outliers and $OPT_{k}$ is the cost of the optimal solution that is allowed to select $k$ outliers. We will discuss this in additional detail here.
In particular, assume that an optimal algorithm for the $k$-outlier MCCT problem removes a subset $K^{*}\subseteq X$. Let $S=X\setminus K^{*}$. The cost of this algorithm, which is $OPT_{k}$, is at least $\sum_{i,j\in S}r_{ij}d(i,j)$. Let $OPT_{\alpha,k}$ be the cost of the optimal $k$-outlier solution on the metric $d_{\alpha}$ (i.e. the optimal solution after the embedding $\alpha$ is applied). We get that if $\mathcal{D}$ is a probabilistic embedding with expected distortion at most $c$, then $\mathop{E}_{\alpha\sim\mathcal{D}}[OPT_{\alpha,k}]\leq E_{\alpha}[\sum_{i,j\in S}r_{ij}d_{\alpha}(i,j)]\leq c\sum_{i,j\in S}r_{ij}d(i,j)\leq 8c\cdot OPT_{k}$, where we used the fact that the optimal cost for the $k$-outlier problem on $d_{\alpha}$ is bounded by the cost of embedding all points but $K^{*}$ into a tree, and we applied the Steiner point removals results of Gupta [31]. Consider probabilistically embedding a given metric into trees using $\mathcal{D}$ and then applying the outlier approximation algorithm from Claim A.1. For an input metric $d_{\alpha}$, this algorithm produces a solution with at most $3k$ outliers and cost at most $3OPT_{\alpha,k}$. Thus, the expected cost of this solution is at most $\mathop{E}_{\alpha\sim\mathcal{D}}[3OPT_{\alpha,k}]\leq 24c\cdot OPT_{k}$, where $c$ is the distortion of embedding $\mathcal{D}$. This is enough to obtain the results we claimed in Section 3.4.
The outlier embedding results from this paper let us improve the factor $c$ that we can obtain for such an approximation if we are willing to increase the number of outliers. Note also that because $\epsilon$ is a tunable parameter in our algorithm, it does not necessarily need to be a constant, so if we are more interested in minimizing the number of outliers, we can also obtain a solution of cost at most $O(\log^{2}k\cdot c_{k}(\mathcal{X})\cdot OPT_{k})$ with only $O(k)$ outliers.
Appendix B Proof of Lemma 4.3
In this appendix, we will prove Lemma 4.3. Notably, this proof follows the same lines as the proof of Lemma 5.2 in [20]. We only repeat it here for completeness and to clarify or simplify a few places in the lemma proof where terms were stated in a way that specific to $\ell_{1}$ embeddings or deterministic embeddings. A modification of the expansion proof in Section 2 of Calinescu et al. [14] could similarly help bound expansion for this algorithm.
We will begin by rewriting a new version of Lemma 5.2 from [20]. Note that parts a, c, d, and e together imply Lemma 4.3.
Lemma B.1:
[…]
  1. (a)
    If $x,y\in S$, then ED^*[d_α(x,y)]c_Sd(x,y).
  2. (b)
    If $x,y\in K_{i}$ for some $K_{i}$ at step 9 of the algorithm, then let $\set{K_{j}}$ be the full partition of $X\setminus S$ by the algorithm, and we get ED^*[d_α(x,y)—{K_j}]c_X(k)d(x,y).
  3. (c)
    Let $y\in X\setminus S$ and let $x\in S$. Then ED^*[d_α(x,y)] (10c_S+9c_X(k))d(x,y).
  4. (d)
    If $x,y\in X\setminus S$ and $d(x,\gamma(x))\leq 2\cdot d(x,y)$ for $\gamma$ as defined in line 2 of the algorithm, then ED^*[d_α(x,y)][50c_S+45c_X(k)]d(x,y).
  5. (e)
    If $x,y\in X\setminus S$ and $d(x,\gamma(x)),d(y,\gamma(y))>2\cdot d(x,y)$ for $\gamma$ as defined in line 2 of the algorithm, then
    ED^*[d_α(x,y)] (125c_S+116c_X(k))H_k d(x,y)
The proof of each part of this lemma will largely follow that of [20], with small changes to account for the input probability distributions of subsets of metrics rather than a single embedding on the entire metric (as well as randomness in the embedding on the good subset), and to account for the fact that we are dealing with a general merge algorithm and host metric, rather than specifically concatenation and $\ell_{1}$. We have also simplified the proof in a few places.
Proof of Lemma B.1a.
Note that for our initial setting of $\alpha$, we set it to be exactly $\alpha_{S}$. Thus, the distance that $x$ and $y$ incur under $\alpha_{S}$ is exactly that incurred under the final $\alpha$. Thus by Fact 1
$\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]=\mathop{E}_{\alpha_{S}\sim\mathcal{D}_{S}}[d_{\alpha_{S}}(x,y)]\leq c_{S}\cdot d(x,y),$
where the last line is by definition of $c_{S}$. ∎
Proof of Lemma B.1b.
We have $x,y\in K_{i}$, so $x$ and $y$ are first added to the domain of $\alpha$ in the $i$th iteration of Step 11 in the algorithm. At this point, we set $\alpha\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$, where $x$ and $y$ are in the domain of $\alpha_{i}$ and $\alpha_{i}$ is sampled from $\mathcal{D}_{K_{i}}$. Due to Fact 1, the distance between $x$ and $y$ under $\alpha$ is unchanged after all future steps, and due to the fact we have a PerfectMerge function, their distance under $\alpha$ at this step is exactly their distance under $\alpha_{i}$. Thus, by Fact 1 at this step we get
$\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|K_{i}]=\mathop{E}_{\alpha_{i}\sim\mathcal{D}_{K_{i}}}[d_{\alpha_{i}}(x,y)]\leq c_{X}(k)\cdot d(x,y),$
where we use the bound on distortion of $\mathcal{D}_{K_{i}}$. ∎
Proof of Lemma B.1c.
We begin by bounding $\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}$], the expectation conditioned on the partition of $\set{K_{j}}$ of $X\setminus S$. Let $y\in K_{i}$ for $K_{i}\in\set{K_{j}}$. We first focus on bounding the expected distortion between $y$ and $\gamma(u_{i})$ (the nearest neighbor of $y$’s cluster’s center).
$\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(\gamma(u_{i}),y)|\set{K_{j}}$]] $=\mathop{E}_{\alpha_{i}\sim\mathcal{D}_{K_{i}}}[d_{\alpha_{i}}(\gamma(u_{i}),y)]$
$\leq c_{X}(k)\cdot d(\gamma(u_{i}),y),$
where the first line is by the fact that when $y$ first joins the domain of $\alpha$, it is as a result of $\texttt{PerfectMerge}(\alpha,\alpha_{i})$ for some $\alpha$ where $\gamma(u_{i})$ is in the domain of both $\alpha$ and $\alpha_{i}$. Since PerfectMerge is a perfect merge function, we get that the distance between $\gamma(u_{i})$ and $y$ is the same as before the merge. The second line is by the expected distortion of $\mathcal{D}_{K_{i}}$.
Now consider a more general $x\in S$. We get
$\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(x,y)|\set{K_{j}}$] $\leq\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(\gamma(u_{i}),x)|\set{K_{j}}]+\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(\gamma(u_{i}),y)|\set{K_{j}}$]
$\leq c_{S}\cdot d(\gamma(u_{i}),x)+c_{X}(k)\cdot d(\gamma(u_{i}),y)$
$\leq c_{S}\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot d(\gamma(u_{i}),y)\text{ (triangle inequality)}$
$\leq c_{S}\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot(d(u_{i},\gamma(u_{i}))+d(u_{i},y))\text{ (triangle inequality)}$
$\leq c_{S}\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot(d(u_{i},x)+4d(x,y))$
$\leq(5c_{S}+4c_{X}(k))\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot(d(u_{i},y)+d(x,y))\text{ (triangle inequality)}$
$\leq(6c_{S}+5c_{X}(k))\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot 4d(x,y)$
$=(10c_{S}+9c_{X}(k))\cdot d(x,y).$
The first line here is by the triangle inequality on $\alpha$. The second line is by Lemma B.1a and the previous case of this lemma. Lines three, four, and six are by the triangle inequality on $(X,d)$, and the last line is by rearranging. Line five is by the fact that $d(u_{i},\gamma(u_{i}))\leq d(u_{i},x)$, as $\gamma(u_{i})$ is a closest node in $S$ to $u_{i}$ and $x\in S$. The fifth and seventh lines also use the fact that $d(u_{i},y)\leq 4d(y,\gamma(y))\leq 4d(y,x)$, as well as using the definition of $\gamma(x)$ and the upper bound on $b$ in the algorithm.
Now we show that this inequality still holds when we do not condition on $\set{K_{j}}$, and we note that the law of total expectation automatically gives us the result.
$\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]$ $=\mathop{E}_{\set{K_{j}}}[\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}$]]
$\leq\mathop{E}_{\set{K_{j}}}[(10c_{S}+9c_{X}(k))\cdot d(x,y)]$
$=(10c_{S}+9c_{X}(k))\cdot d(x,y).\qed$
Proof of Lemma B.1d.
We have
$\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]$ $\leq\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,\gamma(x))]+\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(y,\gamma(x))]$
$\leq(10c_{S}+9c_{X}(k))\cdot d(x,\gamma(x))+(10c_{S}+9c_{X}(k))\cdot d(y,\gamma(x))$
$\leq(20c_{S}+18c_{X}(k))\cdot d(x,\gamma(x))+(10c_{S}+9c_{X}(k))\cdot d(x,y)$
$\leq(50c_{S}+45c_{X}(k))\cdot d(x,y),$
where the first line is by the triangle inequality on $Y(\alpha)$ and linearity of expectation, the second is by Lemma B.1c, the third is by the triangle inequality on $(X,d)$, and the fifth is by the condition of this lemma that $d(x,\gamma(x))\leq 2d(x,y)$. ∎
Before we prove the last part of the Lemma, we will discuss some of the terminology and results from the proof of Lemma 4.1(d) from [20]. This will help abbreviate our proof of Lemma B.1d in this paper.
First, we say that $x,y\in X\setminus S$ are “split” if $x\in K_{i},y\in K_{j}$ for $i\neq j$ from Line 9 of Algorithm 2. Further, we say “$u$ splits $x$ and yy” if $u$ is the first node under the ordering $\pi$ such that exactly one of $x$ and $y$ is in $K_{\pi(u)}$. For a given pair $x,y\in X\setminus S$, we define $\beta_{u,\set{x,y}}:=\min\set{\frac{d(x,u)}{d(x,\gamma(x))},\frac{d(y,u)}{d(y,\gamma(y))}}$. Consider a fixed ordering of the $u\in X\setminus S$ by increasing $\beta_{u,\set{x,y}}$. Then denote the index of $u$ in this ordering as $\texttt{index}_{\set{x,y}}(u)$. From [20] and the fact that partitions in this algorithm are picked the same way as in that paper’s Algorithm 4.1, we get the following claim.
Claim B.2 (from [20]):
For a pair $x,y\in X\setminus K$ and a node $u\in X\setminus K$, let $x_{u}$ be $\arg\min_{x,y}\set{d(x,u),d(x,y)}$. Then in Algorithm 2, we have $\Pr[x\text{ and }y\text{ are split by }u]\leq 5\frac{d(x,y)}{d(x_{u},\gamma(x_{u}))}\cdot\frac{1}{\texttt{index}_{\set{x,y}}(u)}$.
Proof of Lemma B.1e.
Finally, we are left with pairs $x,y\in X\setminus S$ with $d(x,y)\leq d(x,\gamma(x))/2,d(y,\gamma(y))/2$. Consider a set of $\set{K_{j}}$ that does split $x$ and $y$. We get
$\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}$] $\leq\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,\gamma(x))|\set{K_{j}}]+\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(\gamma(x),y)|\set{K_{j}}$]
$\leq(10c_{S}+9c_{X}(k))\cdot d(x,\gamma(x))+(10c_{S}+9c_{X}(k))\cdot d(y,\gamma(x))$
$\leq(20c_{S}+18c_{X}(k))\cdot d(x,\gamma(x))+(10c_{S}+9c_{X}(k))\cdot d(x,y)$
$\leq(25c_{S}+23c_{X}(k))\cdot d(x,\gamma(x)),$
where the first line is by linearity of expectation and the triangle inequality on $Y(\alpha)$, the second line is by Lemma B.1c, the third is by the triangle inequality on $(X,d)$, and the fourth is by the condition of this lemma that $d(x,y)\leq d(x,\gamma(x))/2$ (and rounding up when dividing odd numbers by $2$). Note that by symmetry, we also get $\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}]\leq(25c_{S}+23c_{X}(k))\cdot d(y,\gamma(y))$.
These inequalities hold for every choice of $\set{K_{j}}$ that separates $x$ and $y$. Let’s consider the expected distance overall for $x$ and $y$, and let $H_{k}$ be the $k$th harmonic number. We get
[…]
$\leq c_{X}(k)d(x,y)+\sum_{\set{K_{j}}\text{ that splits }x,y}\Pr[\set{K_{j}}]\cdot(25c_{S}+23c_{X}(k))\cdot\min\set{d(x,\gamma(x)),d(y,\gamma(y))}$
$=c_{X}(k)d(x,y)+(25c_{S}+23c_{X}(k))\cdot\min\set{d(x,\gamma(x)),d(y,\gamma(y))}\sum_{u\in X\setminus S}\Pr[u\text{ splits }x,y]$
$\leq c_{X}(k)d(x,y)+(25c_{S}+23c_{X}(k))\cdot\min\set{d(x,\gamma(x)),d(y,\gamma(y))}\cdot\sum_{u\in X\setminus S}5\frac{d(x,y)}{d(x_{u},\gamma(x_{u}))}\cdot\frac{1}{\texttt{index}_{\set{x,y}}(u)}$
$\leq c_{X}(k)d(x,y)+5(25c_{S}+23c_{X}(k))\cdot d(x,y)\cdot\sum_{u\in X\setminus S}\frac{1}{\texttt{index}_{\set{x,y}}(u)}$
$\leq H_{k}\cdot(125c_{S}+116c_{X}(k))\cdot d(x,y),$
[…]
Appendix C Rounding algorithms and additional details for Section 5
In this appendix, we reproduce the rounding algorithms of [39] and formally show that the proofs of [39] apply to our modified version of their linear program.
Algorithm 6 (Algorithm 1 from [39]): Random selection of partitions for the nodes in $X$
1:for $r\in M$ in decreasing order do
2:  $S\leftarrow X$; $P_{i}^{r}\leftarrow\emptyset$ for all $i\in X$
3:  while $S\neq\emptyset$ do
4:   Choose a center $i\in X$ uniformly at random independent of past choices
5:   Choose $\ell_{i}^{r}\in[0,1]$ uniformly at random independently of past choices
6:   For each $j\in S\cap B_{i}(r)$, if $x_{ij}^{r}\geq\ell_{i}^{r}$, assign $j$ to $P_{i}^{r}$ and remove $j$ from $S$.
Algorithm 7 (Algorithm 2 from [39]): Embedding $(X,d)$ (with diameter $\Delta$) into an HST given randomly chosen partitions $P_{i}^{r}$ from Algorithm 6
1:Place a root node $w$ at the highest level with $\eta_{w}\leftarrow 2\cdot 2^{\lfloor\log\Delta\rfloor}$ and set $S_{w}\leftarrow X$
2:for $r\in M$ in decreasing order do
3:  for each node $w$ at the previous (parent) level with set $S_{w}$ do
4:   for each $i$ with $P_{i}^{r}\cap S_{w}\neq\emptyset$ do
5:     Place a child node $v$ with set $S_{v}\leftarrow P_{i}^{r}\cap S_{w}$ and $\eta_{v}\leftarrow 2r$
Continued proof of Lemma 5.3.
We show that the proposed LP solution satisfies the constraints we did not already address in Section 5. (This reasoning is identical to that in [39].)
  1. 1.
    Constraint 5 is trivially satisfied, as the probability that both $j$ and $j^{\prime}$ have $i$ as a representative at level $r$ is at most the probability that either of them individually has this representative.
  2. 2.
    Constraint 6, the inequality is satisfied for the same reason as in [39], as the embedding is non-contractive, so every node has a representative within a distance $r$ of it at level $r$.
  3. 3.
    Constraint 7 is satisfied for the same reason as in the proof of [39] (i.e. by non-contraction of the embedding, we have $x_{ij}^{r}=0$ if $j\notin B_{i}(r)$ so the left side is the probability that $j$ has some representative at level $r$, which happens with probability 11).
  4. 4.
    Constraint 8 is satisfied by non-contraction of all embeddings, and constraints 9 and 10 are satisfied because all probabilities are in the range $[0,1]$.
Proof of Lemma 5.4.
As Algorithms 6 and 7 are unchanged from [39], we can use the same reasoning. In particular, for any pair of nodes $j,j^{\prime}$, if $j$ and $j^{\prime}$ are in the same part of the partition at the level corresponding to $r$, then there is some $i$ that is a distance at most $r$ from each of them. This means that the first level at which they share a parent is the level corresponding to $r$, and since we double all distances at the end of Algorithm 7 by scaling everything up by $2$, this implies non-contraction. ∎
Appendix D Oblivious probabilistic Lipschitz extensions
While the primary focus of this paper is on strong nested embeddings and their applications to finding outlier HST embeddings, in this section we will also briefly mention how an “oblivious” Lipschitz extension algorithm for deterministic embeddings can also serve as a Lipschitz extension algorithm for probabilistic embeddings. In particular, we consider the case that we have some “oblivious” algorithm that takes in a deterministic embedding $\alpha:S\xhookrightarrow{c_{S}}Y$ for $S\subseteq X$ and samples from a probabilistic Lipschitz extension of that embedding. We will show here that if we have a probabilistic embedding $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$ instead, we can sample any embedding from $\mathcal{D}_{S}$ and run the existing Lipschitz extension algorithm on it in order to obtain a sample from a Lipschitz extension of $\mathcal{D}_{S}$. Notably, the original Lipschitz extension (and corresponding algorithm) must be oblivious, in that for each $x\notin S$, the extension always maps $x$ to the same spot as some $u\in S$, and the choice of such $u$ is completely independent of the input embedding $\alpha_{S}$.
Let us first define the notion of oblivious Lipschitz extensions.
Definition D.1:
A probabilistic Lipschitz extension of a deterministic embedding with extension factor $L$ is a function $f$ that takes in a deterministic embedding $\alpha_{S}:S\xhookrightarrow{c_{S}}Y$ for some space $Y$ and $S\subseteq X$ and outputs a probabilistic embedding $\mathcal{D}:X\xhookrightarrow{}_{p}Y$ such that for all $\alpha$ in the support of $\mathcal{D}$ and all $u\in S$, $\alpha(u)=\alpha_{S}(u)$; and for all $x,y\in X$ $\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(x,y)]\leq L\cdot c_{S}\cdot d(x,y)$.
Further, we say that this extension is oblivious if for all $x\in X\setminus S$, $\alpha(x)\in\set{\alpha(u):u\in S}$ with probability $1$, and for all $u\in S$, $\Pr_{\alpha\sim f(\alpha_{S})}[\alpha(x)=\alpha(u)]$ is the same for all inputs $\alpha_{S}$.
Notably, Lipschitz extensions $\alpha$ on deterministic embeddings $\alpha_{S}$ require that $\alpha(u)=\alpha_{S}(u)$ for all $u\in S$ with probability $1$. In the randomized setting, $u$ is not mapped to a fixed place in a metric space, so we need to clarify what we mean by the concept of a Lipschitz embedding. In particular, we will require not only that the distribution of $u$ under the new embedding be the same as under the old embedding, but also that the joint distribution of the embedding on all points in $S$ is unchanged by the new embedding. Note because of this, we get that the output embedding has the same expected distortion on any pair of nodes as under the original embedding, but in fact this is a much stricter requirement.
We define Lipschitz extensions of probabilistic embeddings below.
Definition D.2:
Let $(X,d_{X})$ be a metric space, let $\mathcal{Y}$ be a set of metric spaces, and $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$ for $S\subseteq X$. Then $\mathcal{D}:X\xhookrightarrow{}_{p}\mathcal{Y}$ is a Lipschitz extension of $\mathcal{D}_{S}$ with extension factor $L$ if:
  1. 1.
    The distribution of $(\alpha(x_{i}))_{x_{i}\in S}$ is unchanged under $\mathcal{D}$ compared to $\mathcal{D}_{S}$;
  2. 2.
    For all $x,y\in X$,
    $\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(x,y)]$ $\leq L\cdot c_{S}\cdot d_{X}(x,y).$
    In other words, the expected expansion under $\mathcal{D}$ should be bounded by $L\cdot c_{S}$.
We now argue that if there are oblivious probabilistic Lipschitz extensions of deterministic embeddings, they imply the existence of Lipschitz extensions of probabilistic embeddings.
Theorem D.3:
Let $f$ be function that produces oblivious probabilistic Lipschitz extensions of embeddings $\alpha_{S}:S\xhookrightarrow{c_{S}}Y$ for any $Y\in\mathcal{Y}\cup\set{S}$ with extension factor $L$.
Then consider $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$ and let $g$ be an algorithm that samples $\alpha_{S}\sim\mathcal{D}_{S}$ and outputs $f(\alpha_{S})$. Let $\mathcal{D}^{*}$ be the distribution produced by running $g(\mathcal{D}_{S})$ and sampling an embedding from the resulting distribution. Then $\mathcal{D}^{*}$ is a Lipschitz extension of $\mathcal{D}_{S}$ with extension factor $L$.
Proof.
Note that we can sample from $\mathcal{D}^{*}$ by sampling $\alpha_{S}$ from $\mathcal{D}_{S}$ and then sampling from $f(\alpha_{S})$. Further, for any $\alpha$ in the support of $f(\alpha_{S})$, $(\alpha(u))_{u\in S}$ is fixed to be exactly $(\alpha_{S}(u))_{u\in S}$. Thus, $\mathcal{D}^{*}$ has the same distribution on $S$ as does $\mathcal{D}_{S}$.
Now we bound expansion for an arbitrary pair of nodes $x,y\in X\setminus S$. Define $K_{u}$ to be the random set equal to $\set{x\in X\setminus S:\alpha(x)=\alpha_{S}(u)}$ when $\alpha$ is drawn from $f(\alpha_{S})$. Note that for $u\neq v\in S$, $\alpha_{S}(u)\neq\alpha_{S}(v)$ because $\alpha_{S}$ is expanding. This implies that the random sets $K_{u}$ form a random partition of $X\setminus S$.
Further, we get that the probability of any particular partition $\set{K_{u}}$ occurring is the same for all distributions in the image of $f$, as $f$ produces an oblivious Lipschitz extension and thus the input to $f$ cannot affect these probabilities.
Let $u_{x}$ be the random variable corresponding to the $u\in S$ such that $\alpha(x)=\alpha_{S}(u)$. Then we get
$\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]$ $=\mathop{E}_{\set{K_{u}}}[\mathop{E}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{u}}$]]
$=\mathop{E}_{\set{K_{u}}}[\mathop{E}_{\alpha_{S}\sim\mathcal{D}_{S}}[d_{\alpha_{S}}(u_{x},u_{y})|\set{K_{u}}$]]
$\leq\mathop{E}_{\set{K_{u}}}[c_{S}\cdot d(u_{x},u_{y})]$
$\leq c_{S}\cdot L\cdot d(x,y).$
The first line comes from the law of total probability. The second line comes from the definition of $u_{x}$ and $u_{y}$. The third line comes from the fact that if $\set{K_{u}}$ is fixed, then $u_{x}$ and $u_{y}$ are fixed, and the fact that the expected distortion between any pair of nodes in $S$ is $c_{S}$ under $\alpha_{S}$. Note that we were able to take the expectation over $\set{K_{u}}$ without setting the choice of $\alpha_{S}\sim\mathcal{D}_{S}$ in the first line due to the obliviousness of the Lipschitz extension.
To see correctness of the fourth line, let $\iota$ be the identity map from $S$ to $S$. Note that
$\mathop{E}_{\set{K_{u}}}[d(u_{x},u_{y})]=\mathop{E}_{\alpha\sim f(\iota)}[d_{\alpha}(x,y)]$, and $f(\iota)$ is a probabilistic Lipschitz extension of the identity function with extension factor $L$. (Further, $\iota$ is in the domain of $f$ because it maps $S$ to $S\in\mathcal{Y}\cup\set{S}$.) ∎
We note that it is essential that the probabilistic Lipschitz extension of the deterministic embedding used in the statement of the above lemma be oblivious. If this is not the case, then sampling from $\mathcal{D}_{S}$ and then applying the extension to the resulting embedding could create a distribution with very high expansion on pairs that are not in $S$. In particular, the worst expansion between any pair of nodes in the drawn embedding can be quite high with high probability, even if the average expansion for any given pair is small. The definition of a Lipschitz extension of a deterministic embedding only requires that the final expansion in the new embedding be bounded with respect to the worst case expansion for any pair in the original embedding. Thus, if we try to apply the arguments in Lemma D.3 to a non-oblivious extension, we may not obtain the same result.
Further, we note that Algorithm 4.1 in [20] in fact takes in any embedding $\alpha_{S}:S\xhookrightarrow{c_{S}}Y$ for any metric space $Y$ and samples from an oblivious probabilistic Lipschitz extension of $\alpha_{S}$ with extension factor $125H_{k}$, where $k:=|X\setminus S|$. Thus, Theorem D.3 implies that for any $\mathcal{Y}$ and any $\mathcal{D}_{S}\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$, there exists a Lipschitz extension $\mathcal{D}$ of $\mathcal{D}_{S}$ with extension factor $125H_{k}$.
Appendix E Ultrametrics
We can also consider how our results affect embeddings into another popular class of metrics called ultrametrics, which are defined as follows.
Definition E.1:
A metric $(X,d)$ is an ultrametric if for all $x,y,z\in X$, $d(x,y)\leq\max\set{d(x,z),d(z,y)}$.
Note that HSTs form a subset of the set of ultrametrics. Further, because HSTs are a subset of all ultrametrics and any ultrametric embeds into an HST with distortion at most $2$ [13], we also obtain the following corollary to Theorem 2.5.
Corollary E.2:
Given a metric $(X,d)$ and target distortion $c$, let $k$ be the optimal outlier set size for probabilistically embedding $(X,d)$ into ultrametrics. Then for any constant $\epsilon>0$, there exists an algorithm that efficiently samples from a probabilistic embedding of some $S\subseteq X$ into ultrametrics with $|X\setminus S|\leq O(\frac{1}{\epsilon}k\log^{2}k)$ and distortion at most $(64+\epsilon)c$.