### v2

# Nested and outlier embeddings into trees

Abstract
In this paper, we consider outlier embeddings into HSTs. In particular, for metric $(X,d)$, let $k$ be the size of the smallest subset of $X$ such that all but that subset (the “outlier set”) can be probabilistically embedded into the space of HSTs with expected distortion at most $c$. Our primary result is showing that there exists an efficient algorithm that takes in $(X,d)$ and a target distortion $c$ and samples from a probabilistic embedding with at most $O(\frac{k}{\epsilon}\log^{2}k)$ outliers and distortion at most $(32+\epsilon)c$, for any $\epsilon>0$. In order to facilitate our results, we show how to find good nested embeddings into HSTs and combine this with an approximation algorithm of Munagala et al. [31] to obtain our results.
1 Introduction
Many algorithmic solutions to metric problems begin by embedding the given metric into a hierarchically separated tree (HST) and then solving the problem on that more structured space. The distortion (the multiplicatively worst change in pairwise distance) of the embedding affects the final approximation factor of such algorithms. Randomized algorithms that optimize the expected value of the objective function can use probabilistic embeddings, which bound the expected distortion for each pair of nodes. A series of works [4, 10, 11, 23] showed that probabilistic embeddings can embed any $n$-point metric into trees with expected distortion $O(\log n)$.
While uniform distortion bounds are helpful in understanding the general effectiveness of using metric embeddings to solve a problem, from a practical perspective, we may ask how well a particular given metric $(X,d)$ can embed into trees. Indeed, for some metrics the optimal distortion is $\Omega(\log n)$ but for others it may be much smaller, e.g. $O(1)$, providing better algorithmic results.
In this paper we investigate instance-optimal distortion for probabilistic embeddings into trees by exploring the role of outliers. For a finite metric $(X,d)$, a $(k,c)$-outlier embedding is a map from “most of” $X$, namely all but $k$ vertices that we call outliers, into another metric space $Y$ such that the distortion over the non-outliers is bounded by $c$. Intuitively, we can view the outliers as noisy or adversarial data that may hurt the quality of the instance-optimal embedding from $X$ into $Y$. The goal of the outlier embedding is to identify and remove these outliers upfront so as to obtain a much better embedding over the remaining points. Algorithmically, given a target distortion $c$, we want to find the smallest $k$ such that the removal of $k$ outliers enables embedding the rest of the metric with distortion at most $c$. In this paper, we focus on outlier embeddings in a probabilistic setting, and give bicriteria approximations that provide tradeoffs between outlier set size and distortion. Our main result is as follows.
Theorem 1.1
Let $(X,d)$ be any metric that admits a $(k,c)$ probabilistic outlier embedding into HSTs. Then for any constant $\epsilon>0$, there exists an algorithm that efficiently samples from a probabilistic embedding into HSTs of some $S\subseteq X$ with $|X\setminus S|\leq O(\frac{1}{\epsilon}k\log^{2}k)$ and expected distortion at most $(32+\epsilon)c$.
Technical tools and challenges
Munagala et al. [31] provide an approximation algorithm for the instance-optimal probabilistic embedding into HSTs that rounds a natural linear programming (LP) relaxation. We extend this LP to incorporate outliers. Proving the feasibility of our extended LP requires the concept of nested embeddings, that was first developed by Chawla and Sheridan [19] in the context of deterministic embeddings into normed metrics. We extend this notion to probabilistic HST embeddings, thereby guiding the choice of distortion constraints in the LP; and then provide a rounding similar to that in [31].
Informally, a nested embedding takes as input an embedding on a subset $S\subseteq X$ with low distortion $c_{S}$, as well as a set of embeddings on all subsets of $X$ with size at most $|X\setminus S|+1$, each with (larger) distortion $c_{X}$; it then finds a single new embedding on the entire set $X$ such that distortion is close to $c_{S}$ for pairs of points in $S$ and is only a polylogarithmic factor worse than $c_{X}$ over other pairs. Extending Chawla and Sheridan [19]’s technique to HSTs poses two key technical challenges that we will now discuss.
Nested embeddings into HSTs with bounded contraction.
Chawla and Sheridan [19] approximate optimal outlier sets by rounding an SDP solution. The SDP has designated “outlier” variables $\delta_{i}$ such that when $\delta_{i}$ is set to $1$, the expansion and contraction conditions for the distance between $i$ and any other node are relaxed; in fact, the contraction can be unbounded. Thus, [19] need only show existence of nested embeddings that bound expansion, not contraction.
By contrast, while our modified version of the LP from Munagala et al. [31] also includes outlier variables $\delta_{i}$, a large $\delta_{i}$ does not admit contraction between nodes, so bounding the optimal LP solution requires nested embeddings that bound both expansion and contraction. In particular, Munagala et al.’s LP essentially sets up a metric labeling problem with the goal of assigning each vertex to a representative at geometrically increasing levels. Each vertex $v$ is assigned a representative at each level $i$ that is at a distance no more than $2^{i}$ from $v$. Importantly, the level-wise labeling constraints cannot be weakened for fractional outliers – each vertex, whether or not it’s an outlier, must still be assigned at level $i$ to a representative within distance $2^{i}$ from it. Even slightly weakening this constraint can potentially introduce errors at all levels of the hierarchy, leading to very large distortion (depending on $|X|$ rather than $k$ and $c$).
In contrast to the main outlier embedding results of Chawla and Sheridan [19], our setting calls for a careful alignment and stitching of the two embeddings that does not contract any distances. We call this procedure a “merge” and characterize the conditions under which it returns a correct nested embedding. Chawla and Sheridan implicitly give such a merge function for $\ell_{1}$ embeddings, and we show that such a merge function exists for HST embeddings.
Probabilistic nested embedding.
The second technical challenge is in extending the notion of nested embeddings to probabilistic embeddings. This is important in the context of tree embeddings because probabilistic tree embeddings can obtain exponentially smaller distortion than deterministic tree embeddings.
Consider taking a probabilistic embedding of some subset $S$ of a given metric $(X,d)$ into a host space $Y$ and extending it to a probabilistic embedding of the entire metric $X$ into $Y$. One potential way of leveraging a deterministic nested embedding is to draw a random embedding $\alpha$ mapping $S$ to $Y$, extend this embedding, and return the extension. What can we say about the distortion of this random extension? We are guaranteed distortion at most $\mathbb{E}_{\alpha}[\max_{u,v\in X}(d_{\alpha}(u,v)/d(u,v))]$, but this can be much larger than the expected distortion of the original embedding, $\max_{u,v\in X}(\mathbb{E}_{\alpha}[d_{\alpha}(u,v)]/d(u,v))$. To avoid this issue, we show that our nested embeddings into HSTs satisfy a certain obliviousness property, and that this property lets us obtain a final distortion guarantee that depends on $\max_{u,v\in X}(\mathbb{E}_{\alpha}[d_{\alpha}(u,v)]/d(u,v))$.
Related work
Embedding into HSTs has been studied primarily with the objective of minimizing “uniform” distortion. See e.g. [4, 10, 11, 16, 23]. In our work, we will use as a subroutine the algorithm of Fakcharoenphol et al. [23], which for any $n$ point metric samples from an $O(\log n)$ distortion probabilistic embedding into HSTs. Instance-optimal distortion has been studied for target spaces such as constant dimensional Euclidean space [8, 13, 22, 29, 35], the line [29, 24, 7, 33], trees [20, 9], and ultrametrics [3, 31]. The goal in such embeddings is to minimize the distortion needed for the given specific $(X,d)$, which in some cases may be smaller than the best bound achievable for all metrics of that size.
The notion of outlier embeddings was initially defined by Sidiropolous et al. [36] when they studied deterministic embeddings into ultrametrics, trees, and constant-dimensional $\ell_{2}$ space. They consider approximating deterministic, isometric outlier embeddings (where the distortion $c$ is 1) and bicriteria approximations with additive distortion. The structural properties used for isometric approximations don’t apply to non-isometric embeddings, and techniques for bounding additive distortion tend to be quite different from those for multiplicative distortion, the focus of this work. Moreover, additive distortion does not provide approximation guarantees for algorithmic applications.
Chubarian and Sidiropolous [21] continued the work on outlier embeddings and studied non-isometric outlier embeddings of unweighted graph metrics into the line. Later, Chawla and Sheridan [19] studied deterministic outlier embeddings into $\ell_{2}$ space with unbounded dimension. The techniques of [19] are similar to the those of Calinescu et al. [14] for the 0-extension problem. The primary differences are: (1) [14] use a semimetric obtained from an LP solution to the 0-extension problem everywhere that [19] use the input metric $d$, and (2) [14] pick a uniformly random permutation over the set $S$ of nodes in the good embedding (which they refer to as terminals), but [19] take a random permutation over the set of $X\setminus S$’s closest neighbors in $S$. We will follow the analysis of [19], but a modified version of the analysis in [14] provides another approach.
Another related notion is embeddings with slack, in which the goal is to minimize the number of edges that are distorted by more than some target factor [1, 15, 27]. Outlier embeddings are also closely related to metric Ramsey theory [12, 30]. Given inputs $n$ and $c$, metric Ramsey theory asks: What is the largest $m$ such that all metrics of size $n$ have a submetric of size $m$ that embeds into some target space with distortion at most $c$? In this paper we are primarily focused on instance-optimal solutions rather than universal bounds that apply to all metrics, and we seek to approximate the size of the outlier set rather than the size of the non-outlier set.
Our notion of nested embeddings first appears in [19], but many other types of embedding compositions or extensions have been considered. One particular example (which we discuss further in Appendix 0.C) is the notion of a Lipschitz extension [26, 32, 14], which seeks to leave non-outliers unaltered from the original embedding and ensure the expansion (but not necessarily the contraction) of every pair of nodes is bounded. Other works study a variety of different embedding compositions [2, 5, 17, 34, 28]. These notions largely differ from ours in that they emphasize bounding the worst distortion for any pair in the entire metric, while our notion focuses on bounding the worst distortion between any pair subject to a “special set” having even stricter distortion constraints.
2 Definitions and results
Expanding embeddings and distortion
A metric space is a set of points $X$ with distances $d_{X}(\cdot,\cdot)$ between them that obey the triangle inequality. That is, for all $x,y,z\in X$, $d_{X}(x,z)\leq d_{X}(x,y)+d_{X}(y,z)$. We assume without loss of generality (via scaling) that in the metrics we consider, the minimum distance between distinct pairs of points is at least $1$. If $\alpha:X\rightarrow Y$ is a map from $(X,d_{X})$ to $(Y,d_{Y})$, we write $d_{\alpha}(x,y):=d_{Y}(\alpha(x),\alpha(y))$ to denote the distance between the images of two points $x,y\in X$ under this map.
$\mathcal{D}$ is a probabilistic embedding of $(X,d_{X})$ into $\mathcal{Y}$ if for each function $\alpha$ in the support of $\mathcal{D}$, $\alpha:X\rightarrow Y(\alpha)$ for some $(Y(\alpha),d_{Y(\alpha)})\in\mathcal{Y}$. This embedding has distortion $c$ if for all $u,v\in X$ we have:
$d_{\alpha}(u,v)$ $\geq d(u,v)\quad\quad\forall\alpha\in\operatorname{support}(\mathcal{D})$
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq c\cdot d(u,v).$
Observe that every embedding in the support of $\mathcal{D}$ is expanding, [^1] Embeddings can also be defined without the expansion requirement, but we will stick to expanding embeddings in this paper. but pairwise distortion is bounded in expectation over the randomness in $\mathcal{D}$. If the support of $\mathcal{D}$ contains a single embedding, it is a deterministic embedding.
We use the notation $\alpha:X\xhookrightarrow{c}Y$ to denote a deterministic embedding with distortion $c$ from $(X,d_{X})$ into $(Y,d_{Y})$ and $\mathcal{D}:X\xhookrightarrow{c}_{p}\mathcal{Y}$ to denote a probabilistic expanding embedding with distortion $c$ from $(X,d_{X})$ into $\mathcal{Y}$.
Outlier embeddings
If $\mathcal{D}$ is a probabilistic embedding from $(X,d_{X})$ into $\mathcal{Y}$ and $K\subseteq X$, define $\mathcal{D}_{X\setminus K}$ to be the distribution identical to $\mathcal{D}$ but with all functions $\alpha$ in the support restricted to the set $X\setminus K$ (i.e. we can sample from $\mathcal{D}_{X\setminus K}$ by sampling from $\mathcal{D}$ and restricting the domain of the function drawn).
Definition 1
A probabilistic embedding $\mathcal{D}$ of $(X,d_{X})$ into $\mathcal{Y}$ is a probabilistic $(k,c)$-outlier embedding if there exists $K\subseteq X$ such that $|K|\leq k$ and $\mathcal{D}_{X\setminus K}$ is a probabilistic embedding of $(X\setminus K,d_{X})$ into $\mathcal{Y}$ with distortion $c$.
Note that in Definition 1, there is a deterministic subset $K$ whose distances to other nodes (and each other) we don’t care about. The randomness is only on the embedding for the remaining nodes in $X\setminus K$. One motivation for outlier embeddings is as a way to stop noise or adversarial data from preventing good embeddings, and we believe that making the outlier set itself deterministic is consistent with this motivation. In Section 5 we also discuss an application that naturally lends itself to finding a deterministic outlier set.
Nested embeddings
As mentioned in earlier, we extend the notion of nested embeddings from Chawla and Sheridan [19] to the probabilistic setting.
Definition 2(Probabilistic composition of nested embeddings, modified from [19])
Let $(X,d_{X})$ be a metric space with $S\subset X$, $k:=|X\setminus S|$, and $\mathcal{Y}$ a class of metrics. Consider an assortment of embeddings: $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$, and for all subsets $K^{\prime}\subseteq X$ with $|K^{\prime}|\leq k+1$, $\mathcal{D}_{K^{\prime}}:K^{\prime}\xhookrightarrow{c_{X}}_{p}\mathcal{Y}$.
An $(\ell,g)$-nested composition of the assortment $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$, where $\ell$ is a constant and $g:[1,\infty)^{2}\times\mathbb{N}\rightarrow[1,\infty)$, is an algorithm that given the assortment returns a probabilistic embedding $\mathcal{D}:X\xhookrightarrow{}_{p}\mathcal{Y}$ such that:
$\text{for all }u,v\in S,\text{ and for all }\alpha\in\operatorname{support}(\mathcal{D}),$ $d_{\alpha}(u,v)$ $\geq d(u,v)$ (1)
$\text{for all }u,v\in S,$ $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq\ell\cdot c_{S}\cdot d(u,v),$ (2)
$\text{and, for all }u,v\in X,$ $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq g(c_{S},c_{X},k)\cdot d(u,v).$ (3)
Hierarchically Separated Trees (HSTs)
Definition 3
A metric $(X,d)$ is a $\beta$-hierarchically separated tree ($\beta$-HST) if there is a map from $X$ to the leaves of a tree such that the following hold:
  • Each node $u$ in the tree is associated with some value $\eta_{u}$, and $\eta_{v}=\beta\cdot\eta_{u}$ whenever $u$ is a child of vv. [^2] In this paper, we will use the term $\beta$-HSTs to refer to what other sources sometimes call exact $\beta$-HSTs. These sources typically define non-exact $\beta$-HSTs as only requiring $\eta_{v}\geq\beta\cdot\eta_{u}$. Further, for all leaves $v$, $\eta_{v}=1/\beta$.
  • For two leaves $x,y\in X$ with least common ancestor $z$, $d(x,y)=\eta_{z}$.
For simplicity, we will primarily focus on $2$-HSTs, and in referring to HSTs, we mean $2$-HSTs unless otherwise noted. Note that focusing on $\beta$-HSTs for a different constant $\beta$ at worst changes results by a constant factor.
Main results
To achieve our main outlier embedding result (Theorem 1.1), we show in Section 3 that HSTs admit nested embeddings.
Theorem 2.1(HSTs have nested embeddings)
Given a probabilistic embedding $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}$ HSTs, there exists an embedding $\mathcal{D}:X\xhookrightarrow{}_{p}\mathcal{H}$ with distortion at most $4c_{S}$ on pairs of nodes in $S$ and distortion at most $\zeta\cdot(c_{S}\log k+\log^{2}k)$ on other pairs, where $k=|X\setminus S|$ and $\zeta$ is a constant.
Further, we can efficiently sample from the distribution $\mathcal{D}$ if we can efficiently sample from $\mathcal{D}_{S}$. We will also briefly discuss an extension of this theorem to ultrametrics in Appendix 0.D.
In Section 4, we use Theorem 2.1 and a modification of the LP from Munagala et al. [31] to prove Theorem 1.1. Our LP-based techniques also allow us to obtain the following result for the weighted variant of the problem. [^3] In the weighted outlier set problem we are given weights $w_{i}$ on all nodes $i$ in the metric space. The cost of a subset $K$ of the metric space is $w(K):=\sum_{i\in K}w_{i}$, and the goal is to find an outlier embedding with some target distortion such that the cost of the outlier set is minimized.
Corollary 1
Given a metric $(X,d)$, node weights $w$, and target distortion $c$, let $w^{*}$ be the optimal outlier set cost for probabilistically embedding $(X,d)$ into HSTs with $c$ distortion and let $k$ be the size of such an optimal outlier set. Then for any constant $\epsilon>0$, there exists an algorithm that efficiently samples from a probabilistic embedding into HSTs of some $S\subseteq X$ with $w(X\setminus S)\leq O(\frac{1}{\epsilon}w^{*}\log^{2}k)$ and expected distortion at most $(32+\epsilon)c$.
3 Probabilistic nested embeddings
In this section we prove Theorem 2.1 by first establishing a set of properties for the existence of good probabilistic nested embeddings. We then show how to achieve these properties for HST embeddings Recall that we are given a metric space $(X,d_{X})$, a target class of metrics $\mathcal{Y}$, a subset $S$ with $k=|X\setminus S|$, and an assortment of probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$. Our goal is to produce a single probabilistic embedding for the entire set $X$ that has small expansion.
Our approach is based on the nested embedding technique of Algorithms 4.1 and 5.1 in Chawla and Sheridan [19] for $\ell_{1}$ embeddings but strengthens that technique by introducing a general procedure for merging embeddings. There are two challenges that we need to address in order to adapt these algorithms to our setting. First, we need to ensure that the final composition does not contract distances. In [19]’s Algorithm 4.1, although the expansion for every pair of points can be bounded for this composition, distances can get contracted by arbitrarily large factors – indeed some vertices end up colocated under the composition. As in [19], we randomly partition $K=X\setminus S$ into sets $K_{1},K_{2},\ldots$, and identify a “close” vertex $\gamma_{i}\in S$ for each such set. Then, we consider the embedding in the given assortment that maps $K_{i}\cup\{\gamma_{i}\}$ into $\mathcal{Y}$. We “merge” this embedding with the embedding $\alpha_{S}$ in such a fashion that neither embedding by itself gets distorted any further. This allows us to bound contraction in a similar way to Algorithm 5.1 from [19], though we frame it more generally. We then iterate in this manner over all of the components in the partition of $K$.
The second challenge is to combine the per-pair $\alpha_{S}$ expansion guarantee into a expected distortion guarantee when we draw an $\alpha_{S}$ from the given probabilistic embedding $\mathcal{D}_{S}$. The key property that allows us to do so is the independence of the procedure for partitioning $K$ and the structure of $\alpha_{S}$. We call this property obliviousness and will revisit it more formally in Appendix 0.C.
We will now identify properties of the merging procedure that we require in order to obtain guarantees for the final composition.
Definition 4(Perfect merge)
Let $Z_{1},Z_{2}\subseteq X$ with $Z_{1}\cap Z_{2}=\{v\}$. Let $\alpha_{1}:Z_{1}\rightarrow Y_{1}$ and $\alpha_{2}:Z_{2}\rightarrow Y_{2}$ be deterministic expanding embeddings, for $Y_{1},Y_{2}\in\mathcal{Y}$. We say that a procedure Merge is a perfect merge function for embedding $X$ into $\mathcal{Y}$ if given any pair $\alpha_{1}$ and $\alpha_{2}$ of embeddings from subsets of $X$ into $\mathcal{Y}$ satisfying the conditions above, it produces a an embedding $\alpha:Z_{1}\cup Z_{2}\rightarrow Y$ for some $Y\in\mathcal{Y}$ such that:
  1. 1.
    For all $x,y\in Z_{1}$, $d_{\alpha}(x,y)=d_{\alpha_{1}}(x,y)$.
  2. 2.
    For all $x,y\in Z_{2}$, $d_{\alpha}(x,y)=d_{\alpha_{2}}(x,y)$.
  3. 3.
    For $x\in Z_{1}\setminus\set{v}$ and $y\in Z_{2}\setminus\set{v}$, we have $d_{\alpha}(x,y)\geq\max\set{d_{\alpha_{1}}(x,v),d_{\alpha_{2}}(v,y)}$.
Now we present Algorithm 1 that, given an assortment of probabilistic embeddings on subsets of $X$, produces a new embedding that is drawn from a strongly nested composition of the input embeddings. We highlight in red where Algorithm 1 differs from Algorithm 5.1 in [19]. In [19], the input PerfectMerge function was merely concatenation of vectors in the $\ell_{1}$ metric. We will develop a different PerfectMerge merge function for embeddings into HSTs in this section.
Algorithm 1 Nested Composition: Algorithm for finding a composition of probabilistic nested embeddings
Input: Metric space $(X,d)$, subset $S\subseteq X$, $k=|X\setminus K|$; Assortment of embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ where $\mathcal{D}_{S}:S\rightarrow\mathcal{Y}$ and $\mathcal{D}_{K^{\prime}}:K^{\prime}\rightarrow\mathcal{Y}$ for $K^{\prime}\subset X$; A function PerfectMerge that takes in two deterministic embeddings $\alpha:Z\rightarrow\mathcal{Y}$ and $\alpha^{\prime}:Z^{\prime}\rightarrow\mathcal{Y}$ for $Z,Z^{\prime}\subset X$ and $|Z\cap Z^{\prime}|=1$, and returns an embedding $\alpha^{\prime\prime}:Z\cup Z^{\prime}\rightarrow\mathcal{Y}$.
Output: A randomly sampled embedding $\alpha:X\rightarrow\mathcal{Y}$.
[…]
2: Sample $\alpha_{S}\leftarrow\mathcal{D}_{S}$
3:Define a function $\gamma:K\rightarrow S$ such that $\gamma(u)\in\arg\min_{v\in S}d(u,v)$. $\triangleright$ i.e. $\gamma(u)$ is one of $u$’s closest neighbors in $S$
4:Select $b$ uniformly at random from the range $[2,4]$
5:Select a uniformly random permutation $\pi:K\rightarrow[k]$ of the vertices in $K$
[…]
7:for $i=1$ to $k$ do
8:   $u_{i}\leftarrow\pi^{-1}(i)$
9:   $K_{i}\leftarrow\{v\in K^{\prime}\ |\ d(v,u_{i})\leq b\cdot d(v,\gamma(v))\}$ $\triangleright$ Let the “center" of $K_{i}$ be $u_{i}$
10:   Sample $\alpha_{i}\leftarrow\mathcal{D}_{K_{i}\cup\set{\gamma(u_{i})}}$
11:   $\alpha\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$
12:  $K^{\prime}\leftarrow K^{\prime}\setminus K_{i}$
13:Output $\alpha$
Theorem 3.1
Let PerfectMerge be a perfect merge function for embeddings from $(X,d)$ into a given class $\mathcal{Y}$. Then, given an assortment of probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ and the procedure PerfectMerge, Algorithm 1 returns a sample from a strong probabilistic $(3,g)$-nested composition with $g(c_{S},c_{X},k)=O((c_{S}+c_{X}(k))\cdot H_{k})$, where $H_{k}$ is the $k$th harmonic number and $c_{X}(k)$ is an upper bound on the distortion for any of the $\mathcal{D}_{K^{\prime}}$ in the input.
We break down the proof of the theorem into two parts: bounding the expansion of the embedding (Lemma 1) and its contraction (Lemma 2).
Lemma 1
Let PerfectMerge be a perfect merge function for embeddings from $(X,d)$ into a given class $\mathcal{Y}$. Then, given an assortment of expanding probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ and the procedure PerfectMerge, Algorithm 1 returns a random embedding $\alpha$ from $X$ to $\mathcal{Y}$ such that for all $x,y\in X$,
$\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(x,y)]$ $\leq(125\,c_{S}+116\,c_{X}(k))\cdot H_{k}\cdot d(x,y).$
Further, if $x,y\in S$, then $\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(x,y)]\leq c_{S}\cdot d(x,y)$.
Lemma 2
Let PerfectMerge be a perfect merge function for embeddings from $(X,d)$ into a given class $\mathcal{Y}$. Then, given an assortment of expanding probabilistic embeddings $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$ and the procedure PerfectMerge, Algorithm 1 returns an embedding $\alpha$ from $X$ to $\mathcal{Y}$ such that for all $x,y\in X$, $d_{\alpha}(x,y)\geq d(x,y)/3$.
We formally prove each of these lemmas in Appendix 0.A.1 and 0.A.2, respectively. Lemma 1 follows essentially the same proof technique as that of Algorithm 5.1 in [19], though we pay special attention to the independence of the $\alpha_{S}$ we draw from $\mathcal{D}_{S}$ and the random partitioning of $K$. To intuitively understand the contraction bound, note that when some node $v\in K_{i}$ is added to the domain of $\alpha$, there is no contraction between $v$ and other nodes in $K_{i}\cup\set{\gamma(u_{i})}$, as $\mathcal{D}_{K_{i}\cup\set{\gamma(u_{i})}}$ is non-contracting and we use a perfect merge function to combine this with $\alpha$. Further, for any node $u$ already in the domain of $\alpha$ when $v$ was added, property 3 of perfect merge functions ensures $d_{\alpha}(u,v)\geq\max\set{d_{\alpha}(u,\gamma(u_{i})),d_{\alpha}(v,\gamma(u_{i}))}$. Applying the triangle inequality and appropriate contraction bounds on $\alpha$ prior to this point gives us the result. Notably, Lemma 2 does allow a small amount of contraction, but if metrics in $\mathcal{Y}$ can be “scaled up” and stay in $\mathcal{Y}$ (such as when $\mathcal{Y}$ is the set of HSTs), we immediately get Theorem 3.1 from Lemmas 1 and 2.
Nested embeddings for HSTs
Algorithm 2 $\texttt{MergeHST}(\alpha_{1},\alpha_{2})$ Algorithm for merging two $\beta$-HSTs
Input: Embeddings $\alpha_{1}:Z_{1}\rightarrow T_{1}$, $\alpha_{2}:Z_{2}\rightarrow T_{2}$ for $\beta$-HSTs $T_{1},T_{2}$, where $|Z_{1}\cap Z_{2}|=1$. We assume without loss of generality that the height of $T_{1}$ is at least as large as the height of $T_{2}$.
Output: An embedding of $Z_{1}\cup Z_{2}$ into a single $\beta$-HST
1:Let $u=Z_{1}\cap Z_{2}$
2:Create a new HST $T$ that is initially a copy of $T_{1}$
[…]
4:for $i=0$ to the height of $T_{1}$ do
5:  for the subtree of each child of $\alpha_{2}(u)$’s ancestor at height $i$ in $T_{2}$ do
6:   if the subtree does not contain $\alpha_{2}(u)$ then $\triangleright$ That is, if the subtree is one we have not yet merged into $T_{1}$
7:     Make a copy of the subtree and make it a child of $\alpha_{1}(u)$’s ancestor at level $i$ in $T$
8:     for any $x\in Z_{2}\setminus\set{u}$ such that $\alpha_{2}(x)$ is in the subtree at this step do
9:      Set $\alpha(x)$ to be the copy of $\alpha_{2}(x)$ created in $T$ at this step
10:Return $\alpha$
We now present a perfect merge function for HSTs. Algorithm 2 takes in two $\beta$-HSTs for a fixed ratio $\beta$ and returns another $\beta$-HST. We show in Lemma 3 that MergeHST is perfect.
Lemma 3
Algorithm 2 is a perfect merge function for embeddings into $\beta$-HSTs.
Figure 1: This figure depicts what happens in the algorithm MergeHST. In particular, $u$ is the node common to both HSTs on the left side of the figure. The triangles represent subtrees, and the letters below the triangles indicate the presence of the corresponding node as a leaf of that subtree. Note that after merging trees, the least common ancestor of $x$ and $y$ is that of $x$ and $u$ from the first tree, as this was higher than that of $y$ and $u$. This is the foundational idea behind why $d_{\alpha}(x,y)\geq\max\set{d_{\alpha_{1}}(u,x),d_{\alpha_{2}}(u,y)}$ holds when $\alpha$ is the output embedding for MergeHST on inputs $\alpha_{1}$ and $\alpha_{2}$.
Figure 1 illustrates how MergeHST works. Given two HSTs that share a common leaf vertex, the function aligns the two HSTs along the path from the shared leaf to the respective roots, and merges the intermediate vertices along the way. By the manner in which we have defined $\beta$-HSTs, the lengths of the edges along the common paths align correctly with each other, enabling the guarantees we require. We defer the correctness proof for this lemma to Appendix 0.A.3. Note that Theorem 3.1 and Lemma 3 plus the fact that all $n$-point metrics probabilistically embed into HSTs with $O(\log n)$ distortion [23] imply Theorem 2.1.
4 Outlier embeddings for HSTs
Now, given the existence of nested embeddings into HSTs, we turn to approximating outlier embeddings into HSTs. In [19], this type of approximation is obtained for $\ell_{2}$ by introducing new “outlier variables” into a semidefinite program for a minimum distortion $\ell_{2}$ embedding and rounding. We will employ a similar technique here using the linear program and associated rounding algorithm of [31] to approximate the optimal distortion for probabilistic HST embeddings.
Let $\Delta$ be the diameter of $(X,d)$ and let $M=\set{2^{0},2,2^{2},\ldots,2^{\lceil\log\Delta\rceil}}$. Let $\zeta$ be the constant factor from Theorem 2.1 and let $B_{i}(r):=\set{j\in X:d(i,j)\leq r}$. We define (HST LP), which is parameterized by inputs $((X,d),k,c)$. The changes from [31] are highlighted in red.
Table 1 gives an intuitive description of the variables in the LP. Note that in our LP $c$ is a parameter, but in [31] it is the variable they are minimizing. The biggest change from [31] is that we have introduced new variables $\delta_{i}$ for each $i\in X$ that we use as an indicator for whether $i$ is an outlier in the embedding.
$\min{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\sum_{i=1}^{n}\delta_{i}}$ (HST LP)
$\forall j,j^{\prime}\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}$ $\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(4+\zeta\cdot(\delta_{j}+\delta_{j^{\prime}})\cdot\log^{2}k)}\cdot c\cdot d(j,j^{\prime})$ (4)
$\forall i,r,(j,j^{\prime}\in B_{i}(r))\ \ \ \ \ \min(x_{i,j}^{r},x_{ij^{\prime}}^{r})$ $\geq z_{ijj^{\prime}}^{r}$ (5)
$\forall j,j^{\prime},i,r\sum_{i:j,j^{\prime}\in B_{i}(r)}z_{ijj^{\prime}}^{r}$ $\geq 1-\gamma_{jj^{\prime}}^{r}$ (6)
$\forall j,r\sum_{i:j\in B_{i}(r)}x_{ij}^{r}$ $=1$ (7)
$\forall j,j^{\prime},(r<d(j,j^{\prime}))\ \ \ \ \ \gamma_{jj^{\prime}}^{r}$ $=1$ (8)
$\forall j,j^{\prime},r\ \ \ \ \gamma_{jj^{\prime}}^{r}$ $\leq 1$ (9)
$\forall i,j,j^{\prime},r\ \ \ \ \ \gamma_{jj^{\prime}}^{r}x_{ij}^{r},z_{ijj^{\prime}}^{r},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\delta_{i}}$ $\geq 0;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\delta_{i}\leq 1}$ (10)
variable intuitive meaning
$\delta_{i}$ indicator for whether $i$ is an outlier
$z_{ijj^{\prime}}^{r}$ the probability that $i$ is the representative of $j$ and $j^{\prime}$ at level $r$
$\gamma_{jj^{\prime}}^{r}$ the probability that $j$ and $j^{\prime}$ are separated at level $r$
$x_{ij}^{r}$ the probability that $i$ is the representative of $j$ at level $r$
Table 1: We imagine assigning a “representative” at each node of the HST such that the representative of a leaf node is the leaf node itself, and the representative of any other node is the same as one of its children’s representatives. This table describes the intuitive meaning of each variable in the LP (which is largely the same as in [31]).
Given an optimal solution to this linear program, we need to define an outlier embedding with low distortion and few outliers. To do so, we will first use the rounding algorithm of [31] to find an embedding on the entire metric, and then we take as outliers all vertices $i$ such that $\delta_{i}\geq\delta^{*}$. In Appendix 0.B we formally prove that this rounding algorithm works and gives us Theorem 1.1 as a result. In a similar manner to $\ell_{2}$ outlier embeddings in [19], we use the existence of nested embeddings to bound the optimal LP solution and thus obtain our final result.
Further, if we modify (HST LP) by changing the objective function to $\sum_{i}w_{i}\cdot\delta_{i}$ for some set of weights $w_{i}$, we can use the same technique to obtain a bicriteria approximation for the weighted variant of the outlier embedding problem, where the cost of making $i$ an outlier is $w_{i}$. This result is expressed as Corollary 1, and we prove it in Appendix 0.B.3.
5 Applications
In the Minimum Cost Communication Tree (MCCT) problem, we are given a finite metric $\mathcal{X}=(X,d)$ and weights $r_{ij}$ for all pairs of nodes $i,j$ and we are asked for a non-contracting embedding $\alpha$ of $X$ into a tree with vertex set $X$ such that $\sum_{i,j\in X}r_{ij}\cdot d_{\alpha}(i,j)$ is minimized. This problem is motivated by evolutionary biology, in which a scientist may attempt to reconstruct a tree of relationships between multiple organisms, and she comes to the problem with some known minimum distances between them, derived through a method such as DNA sequence analysis. In such a setting, the $r_{ij}$ are an additional tunable parameter that can allow the scientist to indicate pairs she considers “more likely” to be related, by giving them a bigger $r_{ij}$ value. As noted by Wu et al. [37], this problem can be approximated in expectation to within a factor of $O(\log n)$ by simply embedding randomly into an HST and converting it to a tree with vertex set $X$. Directly applying the algorithm of Munagala et al. [31] with the observations of Wu et al. [37], we have a randomized algorithm that gets an $O(c(\mathcal{X}))$-approximation for the MCCT problem, where $c(\mathcal{X})$ is the minimum distortion needed to probabilistically embed $\mathcal{X}$ into HSTs.
It is quite natural to consider an outlier version of the MCCT problem. In the evolutionary biology setting, we could consider the case that we are given some noisy data points, or points that don’t really belong in the set being analyzed.
One approach to this type of outlier problem may be to look for a $k$ outlier tree embedding of minimum distortion, but the set of outliers that minimizes the embedding distortion does not necessarily also minimize the MCCT solution cost. We may expect that the outlier MCCT problem is easier to solve on tree metrics than on general metrics, so we could instead try the following approach: embed the input metric into a tree, and then try to find the best outliers for the resulting tree metric. In fact, if no outliers are allowed, MCCT is trivial on tree metrics. Further, in Appendix 0.E we give a bicriteria approximation to the $k$-outlier MCCT problem when the input is a tree metric. In particular, if the optimal cost of a $k$ outlier solution for a given instance of a tree metric is $OPT_{k}$, then we give an algorithm that finds a solution with at most $3k$ outliers and at most $24OPT_{k}$ cost. Let $c_{k}(\mathcal{X})$ be the minimum distortion required for probabilistically embedding all but at most $k$ points in $\mathcal{X}$ into HSTs. We can use the outlier embedding algorithm from this paper to embed any input metric into a tree with outliers and then apply the approximation algorithm for outlier MCCT on trees, obtaining a solution with at most $O(\frac{k}{\epsilon}\log^{2}k)$ outliers and expected cost $O((1+\epsilon)\cdot c_{k}(\mathcal{X})\cdot OPT_{k})$. By contrast, directly using the embedding algorithm of [31] with the tree approximation algorithm from Appendix 0.E instead gives a solution with at most $3k$ outliers and $O((1+\epsilon)\cdot c_{0}(\mathcal{X})\cdot OPT_{k})$ expected cost, so we are unable to save on the cost of the optimal solution if there is a large subset that embeds well into HSTs but the entire metric does not.
{credits}
5.0.1 Acknowledgements
This research was funded in part by NSF awards CCF-2217069 and CCF-2225259. This research was funded in part by JPMorgan Chase & Co. Any views or opinions expressed herein are solely those of the authors listed, and may differ from the views and opinions expressed by JPMorgan Chase & Co. or its affiliates. This material is not a product of the Research Department of J.P. Morgan Securities LLC. This material should not be construed as an individual recommendation for any particular client and is not intended as a recommendation of particular securities, financial instruments or strategies for a particular client. This material does not constitute a solicitation or offer in any jurisdiction.
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Omitted proofs from Section 3
In this appendix we give formal proofs that were omitted in Section 3. We first prove Lemma 1 (bounding expansion) and Lemma 2 (bounding contraction) in Sections 0.A.1 and 0.A.2. Finally, we prove Lemma 3 in Section 0.A.3.
An essential consequence of using a perfect merge function with Algorithm 1 is that when bounding expansion or contraction for a pair $x,y$, we need only bound this value at the point in the algorithm where they both appear in the domain of $\alpha$ for the first time. At any subsequent iteration, as $\alpha$ undergoes further merge steps, the first condition in Definition 4 ensures that the distance between $x$ and $y$ stays the same. We summarize this observation in the following fact.
Fact 1
Let $x,y\in X$. Consider the first iteration of Algorithm 1 in which both $x$ and $y$ are in the domain of $\alpha$, and let $\hat{d}$ be $d_{\alpha}(x,y)$ at that time step. Then at every subsequent iteration of the algorithm, $d_{\alpha}(x,y)=\hat{d}$.
0.A.1 Proof of Lemma 1
Notably, this proof follows the same lines as the proof of Lemma 5.2 in [19]. We only repeat it here for completeness and to clarify or simplify a few places in the lemma proof where terms were stated in a way that specific to $\ell_{1}$ embeddings or deterministic embeddings. A modification of the expansion proof in Section 2 of Calinescu et al. [14] could similarly help bound expansion for this algorithm. We will begin by giving a new version of Lemma 5.2 from [19]. Note that parts a, c, d, and e together imply Lemma 1.
Lemma 4
[…]
  1. (a)
    If $x,y\in S$, then ED^*[d_α(x,y)]c_Sd(x,y).
  2. (b)
    If $x,y\in K_{i}$ for some $K_{i}$ at step 9 of the algorithm, then let $\set{K_{j}}$ be the full partition of $X\setminus S$ by the algorithm, and we get ED^*[d_α(x,y)|{K_j}]c_X(k)d(x,y).
  3. (c)
    If $y\in X\setminus S$ and $x\in S$, then ED^*[d_α(x,y)] (10c_S+9c_X(k))d(x,y).
  4. (d)
    If $x,y\in X\setminus S$ and $d(x,\gamma(x))\leq 2\cdot d(x,y)$ for $\gamma$ as defined in line 2 of the algorithm, then ED^*[d_α(x,y)][50c_S+45c_X(k)]d(x,y).
  5. (e)
    If $x,y\in X\setminus S$ and $d(x,\gamma(x)),d(y,\gamma(y))>2\cdot d(x,y)$ for $\gamma$ as defined in line 2 of the algorithm, then
    ED^*[d_α(x,y)] (125c_S+116c_X(k))H_k d(x,y)
The proof of each part of this lemma will largely follow that of [19], with small changes to account for the input probability distributions of subsets of metrics rather than a single embedding on the entire metric (as well as randomness in the embedding on the good subset), and to account for the fact that we are dealing with a general merge algorithm and host metric, rather than specifically concatenation and $\ell_{1}$. We have also simplified the proof in a few places.
Proof of Lemma 4a. Note that our initial setting of $\alpha$ is exactly $\alpha_{S}$. Thus, the distance that $x$ and $y$ incur under $\alpha_{S}$ is exactly that incurred under the final $\alpha$. Thus by Fact 1
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]=\mathop{\mathbb{E}}_{\alpha_{S}\sim\mathcal{D}_{S}}[d_{\alpha_{S}}(x,y)]\leq c_{S}\cdot d(x,y),$
where the last line is by definition of $c_{S}$. ∎
Proof of Lemma 4b. We have $x,y\in K_{i}$, so $x$ and $y$ are first added to the domain of $\alpha$ in the $i$th iteration of Step 11 in the algorithm. At this point, we set $\alpha\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$, where $x$ and $y$ are in the domain of $\alpha_{i}$ and $\alpha_{i}$ is sampled from $\mathcal{D}_{K_{i}}$. Due to Fact 1, the distance between $x$ and $y$ under $\alpha$ is unchanged after all future steps, and due to the fact we have a PerfectMerge function, their distance under $\alpha$ at this step is exactly their distance under $\alpha_{i}$. Thus, by Fact 1 at this step we get
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}]=\mathop{\mathbb{E}}_{\alpha_{i}\sim\mathcal{D}_{K_{i}}}[d_{\alpha_{i}}(x,y)]\leq c_{X}(k)\cdot d(x,y),$
where we use the bound on distortion of $\mathcal{D}_{K_{i}}$. ∎
Proof of Lemma 4c. We begin by bounding $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}$], the expectation conditioned on the partition of $\set{K_{j}}$ of $X\setminus S$. Let $y\in K_{i}$ for $K_{i}\in\set{K_{j}}$. We first focus on bounding the expected distortion between $y$ and $\gamma(u_{i})$ (the nearest neighbor of $y$’s cluster’s center).
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(\gamma(u_{i}),y)|\set{K_{j}}$]] $=\mathop{\mathbb{E}}_{\alpha_{i}\sim\mathcal{D}_{K_{i}}}[d_{\alpha_{i}}(\gamma(u_{i}),y)]$
$\leq c_{X}(k)\cdot d(\gamma(u_{i}),y),$
where the first line is by the fact that when $y$ first joins the domain of $\alpha$, it is as a result of $\texttt{PerfectMerge}(\alpha,\alpha_{i})$ for some $\alpha$ where $\gamma(u_{i})$ is in the domain of both $\alpha$ and $\alpha_{i}$. Since PerfectMerge is a perfect merge function, we get that the distance between $\gamma(u_{i})$ and $y$ is the same as before the merge. The second line is by the expected distortion of $\mathcal{D}_{K_{i}}$.
Now consider a more general $x\in S$. We get
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(x,y)|\set{K_{j}}$] $\leq\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(\gamma(u_{i}),x)|\set{K_{j}}]+\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(\gamma(u_{i}),y)|\set{K_{j}}$]
$\leq c_{S}\cdot d(\gamma(u_{i}),x)+c_{X}(k)\cdot d(\gamma(u_{i}),y)$
$\leq c_{S}\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot d(\gamma(u_{i}),y)\text{ (triangle inequality)}$
$\leq c_{S}\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot(d(u_{i},\gamma(u_{i}))+d(u_{i},y))\text{ (triangle inequality)}$
$\leq c_{S}\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot(d(u_{i},x)+4d(x,y))$
$\leq(5c_{S}+4c_{X}(k))\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot(d(u_{i},y)+d(x,y))\text{ (triangle inequality)}$
$\leq(6c_{S}+5c_{X}(k))\cdot d(x,y)+(c_{S}+c_{X}(k))\cdot 4d(x,y)$
$=(10c_{S}+9c_{X}(k))\cdot d(x,y).$
The first line here is by the triangle inequality on $\alpha$. The second line is by Lemma 4a and the previous case of this lemma. Lines three, four, and six are by the triangle inequality on $(X,d)$, and the last line is by rearranging. Line five is by the fact that $d(u_{i},\gamma(u_{i}))\leq d(u_{i},x)$, as $\gamma(u_{i})$ is a closest node in $S$ to $u_{i}$ and $x\in S$. The fifth and seventh lines also use the fact that $d(u_{i},y)\leq 4d(y,\gamma(y))\leq 4d(y,x)$, as well as the definition of $\gamma(x)$ and the upper bound on $b$ in the algorithm.
Now we show that this inequality still holds when we do not condition on $\set{K_{j}}$, and we note that the law of total expectation immediately gives us the result.
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]$ $=\mathop{\mathbb{E}}_{\set{K_{j}}}[\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}$]]
$\leq\mathop{\mathbb{E}}_{\set{K_{j}}}[(10c_{S}+9c_{X}(k))\cdot d(x,y)]$
$=(10c_{S}+9c_{X}(k))\cdot d(x,y).$
Proof of of Lemma 4d. We have
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]$ $\leq\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,\gamma(x))]+\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(y,\gamma(x))]$
$\leq(10c_{S}+9c_{X}(k))\cdot d(x,\gamma(x))+(10c_{S}+9c_{X}(k))\cdot d(y,\gamma(x))$
$\leq(20c_{S}+18c_{X}(k))\cdot d(x,\gamma(x))+(10c_{S}+9c_{X}(k))\cdot d(x,y)$
$\leq(50c_{S}+45c_{X}(k))\cdot d(x,y),$
where the first line is by the triangle inequality on $Y(\alpha)$ and linearity of expectation, the second is by Lemma 4c, the third is by the triangle inequality on $(X,d)$, and the fifth is by the condition of this lemma that $d(x,\gamma(x))\leq 2d(x,y)$. ∎
Before we prove the last part of the Lemma, we will discuss some of the terminology and results from the proof of Lemma 4.1(d) from [19]. This will help abbreviate our proof of Lemma 4d in this paper.
First, we say that $x,y\in X\setminus S$ are “split” if $x\in K_{i},y\in K_{j}$ for $i\neq j$ from Line 9 of Algorithm 1. Further, we say “$u$ splits $x$ and yy” if $u$ is the first node under the ordering $\pi$ such that exactly one of $x$ and $y$ is in $K_{\pi(u)}$. For a given pair $x,y\in X\setminus S$, we define $\beta_{u,\set{x,y}}:=\min\set{\frac{d(x,u)}{d(x,\gamma(x))},\frac{d(y,u)}{d(y,\gamma(y))}}$. Consider a fixed ordering of the $u\in X\setminus S$ by increasing $\beta_{u,\set{x,y}}$. Then denote the index of $u$ in this ordering as $\texttt{index}_{\set{x,y}}(u)$. From [19] and the fact that partitions in this algorithm are picked the same way as in that paper’s Algorithm 4.1, we get the following claim.
Claim 1 (from [19])
For a pair $x,y\in X\setminus K$ and a node $u\in X\setminus K$, let $x_{u}$ be $\arg\min_{x,y}\set{d(x,u),d(x,y)}$. Then in Algorithm 1, we have $\Pr[x\text{ and }y\text{ are split by }u]\leq 5\frac{d(x,y)}{d(x_{u},\gamma(x_{u}))}\cdot\frac{1}{\texttt{index}_{\set{x,y}}(u)}$.
Proof of Lemma 4e. Finally, we are left with pairs $x,y\in X\setminus S$ with $d(x,y)\leq d(x,\gamma(x))/2,d(y,\gamma(y))/2$. Consider a set of $\set{K_{j}}$ that does split $x$ and $y$. We get
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}$] $\leq\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,\gamma(x))|\set{K_{j}}]+\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(\gamma(x),y)|\set{K_{j}}$]
$\leq(10c_{S}+9c_{X}(k))\cdot d(x,\gamma(x))+(10c_{S}+9c_{X}(k))\cdot d(y,\gamma(x))$
$\leq(20c_{S}+18c_{X}(k))\cdot d(x,\gamma(x))+(10c_{S}+9c_{X}(k))\cdot d(x,y)$
$\leq(25c_{S}+23c_{X}(k))\cdot d(x,\gamma(x)),$
where the first line is by linearity of expectation and the triangle inequality on $Y(\alpha)$, the second line is by Lemma 4c, the third is by the triangle inequality on $(X,d)$, and the fourth is by the condition of this lemma that $d(x,y)\leq d(x,\gamma(x))/2$ (and rounding up when dividing odd numbers by $2$). Note that by symmetry, we also get $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}]\leq(25c_{S}+23c_{X}(k))\cdot d(y,\gamma(y))$.
These inequalities hold for every choice of $\set{K_{j}}$ that separates $x$ and $y$. Let’s consider the expected distance overall for $x$ and $y$, and let $H_{k}$ be the $k$th harmonic number. We get
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]$ $=\Pr[x,y\text{ are not split}]\cdot\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|x,y\text{ are not split}]$
$\ \ \ \ \ +\sum_{\set{K_{j}}\text{ that splits }x,y}\Pr[\set{K_{j}}]\cdot\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{j}}$]
$\leq c_{X}(k)d(x,y)$
$\ \ \ \ \ +\sum_{\set{K_{j}}\text{ that splits }x,y}\Pr[\set{K_{j}}]\cdot(25c_{S}+23c_{X}(k))\cdot\min\set{d(x,\gamma(x)),d(y,\gamma(y))}$
$=c_{X}(k)d(x,y)+(25c_{S}+23c_{X}(k))\cdot\min\set{d(x,\gamma(x)),d(y,\gamma(y))}\sum_{u\in X\setminus S}\Pr[u\text{ splits }x,y]$
$\leq c_{X}(k)d(x,y)$
$\ \ \ \ \ +(25c_{S}+23c_{X}(k))\cdot\min\set{d(x,\gamma(x)),d(y,\gamma(y))}\cdot\sum_{u\in X\setminus S}5\frac{d(x,y)}{d(x_{u},\gamma(x_{u}))}\cdot\frac{1}{\texttt{index}_{\set{x,y}}(u)}$
$\leq c_{X}(k)d(x,y)+5(25c_{S}+23c_{X}(k))\cdot d(x,y)\cdot\sum_{u\in X\setminus S}\frac{1}{\texttt{index}_{\set{x,y}}(u)}$
$\leq H_{k}\cdot(125c_{S}+116c_{X}(k))\cdot d(x,y),$
[…]
0.A.2 Proof of Lemma 2
We will now prove Lemma 2. We now proceed to bound the contraction. Note that we do get a small amount of contraction, so to actually obtain an expanding embedding, we need to be able to “scale up” the set of points we map to while still staying in $\mathcal{Y}$.
Proof of Lemma 2. We will actually prove a slightly stronger claim about contraction: for $x,y\in X$, we have
  1. 1.
    If $x$ and $y$ are in $S$, then $d_{\alpha}(x,y)\geq d(x,y)$.
  2. 2.
    If at least one of $x,y$ is in $S$, then $d_{\alpha}(x,y)\geq d(x,y)/2$.
  3. 3.
    For all $x,y$, $d_{\alpha}(x,y)\geq d(x,y)/3$.
As established by Fact 1, we need only focus on the time when $x$ and $y$ first appear in the domain of $\alpha$.
  1. 1.
    Consider $x,y\in S$. Note $\alpha$ is first defined as $\alpha_{S}\leftarrow\mathcal{D}_{S}$, so at this point $x$ and $y$ are both in its domain and $d_{\alpha}(x,y)=d_{\alpha_{S}}(x,y)\geq d(x,y)$, by the fact that all embeddings in $\mathcal{D}_{S}$ are expanding.
  2. 2.
    Now consider $x\in S,y\in X\setminus S$. We show that no matter the choice of $\set{K_{j}}$ (the partition of the outliers), the claim holds. In particular, consider some choice of $\set{K_{j}}$ with $y\in K_{i}$ for some $i$. Then let $\alpha_{i}$ be in the support of $\mathcal{D}_{K_{i}}$, and define $\alpha^{\prime}\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$, where $\alpha$ is the embedding at the start of this iteration in the algorithm. Also let $\gamma(u_{i})$ be the vertex common to the two embeddings that are being merged. Then we get from Definition 4
    $d_{\alpha^{\prime}}(x,y)$ $\geq\max\set{d_{\alpha}(x,\gamma(u_{i})),d_{\alpha_{i}}(\gamma(u_{i}),y)}$
    $\geq\max\set{d(x,\gamma(u_{i})),d(\gamma(u_{i}),y)}$
    $\geq d(x,y)/2$
    where the first line is because PerfectMerge is a perfect merge function, the second line is by the first part of this claim and the fact that $\alpha_{i}$ is expanding, and the third line is by applying the triangle inequality on $d(x,y)$.
  3. 3.
    Finally, consider $x,y\in X\setminus S$. Again, fix a choice of the partition $\set{K_{j}}$ from the algorithm. If $x,y\in K_{i}$ for the same $i$, let $\alpha_{i}$ be in the support of $\mathcal{D}_{K_{i}}$ and let $\alpha^{\prime}\leftarrow\texttt{PerfectMerge}(\alpha,\alpha_{i})$. Then $d_{\alpha^{\prime}}(x,y)=d_{\alpha_{i}}(x,y)\geq d(x,y)$, where we used the fact that we have a perfect merge function and $\alpha_{i}$ is expanding.
    Thus, we are left with the case that $x\in K_{i},y\in K_{j}$ for $i\neq j$. Wlog assume $i<j$. Then $x$ and $y$ first appear in the domain of $\alpha$ at iteration $j$ of the algorithm. Let $\alpha$ be the embedding before iteration $j$ and let $\alpha^{\prime}$ be the embedding that results from the merge in iteration $j$. Let $\alpha_{j}$ be drawn from the support of $\mathcal{D}_{K_{j}}$. Then, recalling that $x$ is in the domain of $\alpha$ prior to the merge, we get
    $d_{\alpha^{\prime}}(x,y)$ $\geq\max\set{d_{\alpha}(x,\gamma(u_{j})),d_{\alpha_{j}}(\gamma(u_{j}),y)}$
    $\geq\max\set{d(x,\gamma(u_{j}))/2,d(\gamma(u_{j}),y)}$
    $\geq d(x,y)/3,$
    where the first line is because we have a perfect merge function; the second line is by the second case of this claim, the fact that $\gamma(u_{j})\in S$, and the fact that all embeddings in $\mathcal{D}_{K_{j}}$ are expanding; the third is by the triangle inequality applied to $d(x,y)$. $\square$
0.A.3 Proof of Lemma 3
Finally, we prove Lemma 3
Proof of Lemma 3. Let $x\in Z_{1}$ and $y\in Z_{2}$ with $Z_{1}\cap Z_{2}=\set{u}$ and let $\alpha_{1}:Z_{1}\rightarrow T_{1}$, $\alpha_{2}:Z_{2}\rightarrow T_{2}$, where both $T_{1}$ and $T_{2}$ are $\beta$-HSTs with elements of $Z_{1}$ and $Z_{2}$ mapped to their leaves respectively. Let $\alpha\leftarrow\texttt{MergeHST}(\alpha_{1},\alpha_{2})$ be the resulting $\beta$-HST embedding from $Z_{1}\cup Z_{2}$ into $T$.
We first argue that $T$ is indeed a $\beta$-HST. Consider a node at height $i$ in $T$. If this node was originally in $T_{1}$, then any subtree that may have been added as a subtree of this node has height exactly i1i-1 (by the fact that they are copies of subtrees of the children of $\alpha_{2}(u)$’s ancestor at height $i$), so this node is still at height $i$ and is still labeled properly. Now consider the case that this is a copy of a node from $T_{2}$. If the original was at height $i$ in $T_{2}$, then this copy is also at height $i$, as its subtree is identical to that in $T_{2}$. Therefore, the node still retains the correct label.
We will show each of the perfect merge criteria separately. Let $LCA(x,y)$ be the least common ancestor of $\alpha(x)$ and $\alpha(y)$ in the final merged tree $T$.
  • First consider the case that $x,y\in Z_{1}$. Note that $T$ is identical to $T_{1}$ except with some extra children added to some internal nodes, and $\alpha$ maps $x$ and $y$ to exactly the same leaves as under $\alpha_{1}$, so their distance is unchanged.
  • Now consider $x,y\in Z_{2}$. Let $i^{*}$ be the height of the least common ancestor of $\alpha_{2}(x)$ and $\alpha_{2}(y)$ in $T_{2}$. Let $S_{i}(x)$ be the subtree of $T_{2}$ of height $i$ that contains $x$. Note that by definition of $i^{*}$, $S_{i^{*}-1}(x)\neq S_{i^{*}-1}(y)$ (which also holds for all $i<i^{*}$).
    First, consider the case that $S_{i^{*}-1}(x)\neq S_{i^{*}-1}(u)$ and $S_{i^{*}-1}(y)\neq S_{i^{*}-1}(u)$. Then the root of $S_{i^{*}-1}(x)$ and root of $S_{i^{*}-1}(y)$ are both made children of $\alpha(u)$’s ancestor at height $i^{*}$. Thus, their least common ancestor is still at level $i^{*}$ and their distance is unchanged.
    Next, consider the case that $S_{i^{*}-1}(x)=S_{i^{*}-1}(u)$ and $S_{i^{*}-1}(y)\neq S_{i^{*}-1}(u)$. (Note that at most one can be equal to $S_{i^{*}-1}(u)$ so we may assume this without loss of generality.) The first ancestor $\alpha(x)$ and $\alpha(y)$ could have in common is $\alpha(u)$’s ancestor at level $i^{*}$, as $\alpha(x)$’s first ancestor from $T_{1}$ is below this level (and thus all ancestors above this level are also from $T_{1}$), and $\alpha(y)$’s lowest ancestor in common with $u$ is at level $i^{*}$. Further, they do have this ancestor in common by construction, so the least common ancestor height under $\alpha$ is the same as under $\alpha_{2}$. Thus, we get $d_{\alpha}(x,y)=d_{\alpha_{2}}(x,y)$.
  • Finally consider the case that $x\in Z_{1}\setminus\set{u}$ and $y\in Z_{2}\setminus\set{u}$. Let $LCA(a,b)$ be the least common ancestor of $\alpha(a)$ and $\alpha(b)$ in $T$. Note that by definition of HSTs, for any set of nodes $\set{a,b,c,d}$, we have $d_{\alpha}(a,b)\geq d_{\alpha}(c,d)$ if and only if $LCA(a,b)$ is at or above the height of $LCA(c,d)$.
    We will show $d_{\alpha}(x,y)\geq d_{\alpha}(x,u)=d_{\alpha_{1}}(x,u)$ and $d_{\alpha}(x,y)\geq d_{\alpha}(y,u)=d_{\alpha_{2}}(y,u)$ separately.
    • $d_{\alpha}(x,y)\geq d_{\alpha}(x,u)$: Note that $LCA(x,y)$ is an original member of $T_{1}$, as nodes in $T_{1}$ only have ancestors from $T_{1}$. Further, by construction all ancestors of $\alpha(y)$ that are from $T_{1}$ are also ancestors of $\alpha(u)$. Thus, $LCA(x,y)$ must also be an ancestor of $\alpha(u)$ and it must be at or above the height of $LCA(x,u)$.
    • $d_{\alpha}(x,y)\geq d_{\alpha}(y,u)$: Note that $LCA(y,u)$ is $\alpha(y)$’s lowest height ancestor that was an original member of $T_{1}$, as all ancestors below that height are from a copied subtree from $T_{2}$. Further, since $\alpha(x)$ only has ancestors that were original members of $T_{1}$, $LCA(x,y)$ must be at or above the height of $LCA(y,u)$.
Appendix 0.B Omitted proofs from Section 4
In this appendix, we formally show that the rounding algorithm we present for HST LP succeeds in giving the promised approximation from Theorem 1.1. We also formalize the algorithm described in Section 4, and for completeness we write out the algorithms of Munagala et al. [31] used in this paper.
0.B.1 Bounding the optimal LP solution
To bound the size of the final outlier set obtained by our rounding algorithm, we first need to argue that the optimal solution to this LP has value at most $k$. This is expressed in Lemma 5.
Lemma 5
If HST LP is parameterized by $((X,d),k,c)$ and $(X,d)$ has a probabilistic $(k,c)$-outlier embedding into HSTs, then the optimal solution to HST LP has value at most $k$.
This lemma is true for the same reasons given in [31]. [^4] Note that in order to obtain Lemma 5, we need to have that the final embedding has no contraction, even on outliers, as we need to be able to guarantee that any node $j$ is represented at level $r$ by a node $i$ in $B_{j}(r)$. If we allow contraction for outliers, this may not be the case, and $j$ may be represented at each level by something that is actually quite far away from it. The primary change we have made is that $c$ is now a constant, and in finding a feasible solution to the LP, we need to assign values to the $\delta_{i}$ in addition to assigning the other variables as in [31]. Note that if we assign $\delta_{i}$ to be $1$ if $i$ is an outlier, then the right hand side of constraint 4 is an upper bound on the distortion for any pair of nodes under the nested composition of embeddings by Theorem 2.1. Note in particular that constraints 6 through 8 require that the underlying embedding be expanding on all pairs, even when one or both nodes may be an outlier. Thus, bounding the optimal solution to this LP is why it was essential to find nested compositions of embeddings for HSTs, rather than just Lipschitz extensions.
Proof of Lemma 5. Let $\mathcal{D}_{S}$ be a probabilistic embedding on some $S\subseteq X$ into HSTs with distortion $c$ such that $K:=|X\setminus S|$ has size at most $k$. By Theorem 2.1, there exists a probabilistic embedding $\mathcal{D}$ that is non-contracting, has distortion at most $4c$ on pairs of nodes in $S$, and has overall distortion at most $\zeta\cdot(c\log k+\log^{2}k)$ for some constant $\zeta$. Set the variables $z_{ijj^{\prime}}^{r},\gamma_{jj^{\prime}}^{r},$ $x_{ij}^{r}$, and $\delta_{i}$ as in Table 1. (That is, $z_{ijj^{\prime}}^{r}$ is the probability that $i$ is the “representative” of $j$ and $j^{\prime}$ at level $r$, $\gamma_{jj^{\prime}}^{r}$ is the probability that $j$ and $j^{\prime}$ are separated at level $r$, $x_{ij}^{r}$ is the probability that $i$ is the representative of $j$ at level $r$, and $\delta_{i}$ is $1$ if $i$ is an outlier and 0 otherwise.) By how we set the $\delta_{i}$ values, this solution clearly has value $k$, so we need only show it satisfies the constraints.
Each of the constraints in this LP is satisfied for essentially the same reason as in [31]. We will briefly state those reasons here, paying special attention to Constraint 4, as we did alter this constraint from [31] (though it is satisfied for essentially the same reasons as in [31]).
To see why constraint 4 is satisfied, let $\mu_{jj^{\prime}}^{r}$ be the probability that $j$ and $j^{\prime}$ are a distance exactly $r$ apart in the embedding sampled from $\mathcal{D}$. Note that if $j$ and $j^{\prime}$ are separated at level $r$, then their distance in that embedding is at least $r^{*}>r$. Thus, $\gamma_{jj^{\prime}}^{r}$ is the probability that the embedding drawn from the distribution places $j$ and $j^{\prime}$ at a distance strictly greater than $r$ apart, which means $\gamma_{jj^{\prime}}^{r}=\sum_{r^{*}\in M:r^{*}>r}\mu_{jj^{\prime}}^{r^{*}}$. Thus we get
$\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}$ $=\sum_{r\in M}r\cdot\sum_{r^{*}\in M:r^{*}>r}\mu_{jj^{\prime}}^{r^{*}}$
$=\sum_{r^{*}\in M}\mu_{jj^{\prime}}^{r^{*}}\sum_{r\in M:r^{*}>r}r$
$\leq\sum_{r^{*}\in M}r^{*}\mu_{jj^{\prime}}^{r^{*}}$
$=\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}\left[d_{\alpha}(j,j^{\prime})\right].$
The first line is by what we argued above; the second is by reordering the sums; the third is by the fact that $\sum_{r\in M:r<r^{*}}r\leq r^{*}$ (as all values in $M$ are powers of $2$); the fourth line is by definition of $\mu_{jj^{\prime}}^{r}$ as the probability that $j$ and $j^{\prime}$ are a distance exactly $r$ apart. Note that if at least one of $j,j^{\prime}$ is an outlier, then at least one of $\delta_{j}$ and $\delta_{j^{\prime}}$ is $1$, so the right hand side of constraint 4 is at least $\zeta\cdot c\cdot\log^{2}k\cdot d(j,j^{\prime})$, which is an upper bound on the expected distance between $j$ and $j^{\prime}$ under $\mathcal{D}$. If neither $j$ nor $j^{\prime}$ is an outlier, then the right hand side of constraint 4 is $4c\cdot d(j,j^{\prime})$, which is an upper bound on their expected distance under $\mathcal{D}$.
Now we can briefly describe why the other constraints are satisfied.
  1. 1.
    Constraint 5 is satisfied, as the probability that both $j$ and $j^{\prime}$ have $i$ as a representative at level $r$ is at most the probability that either of them individually has this representative.
  2. 2.
    Constraint 6 is satisfied because the embedding is non-contracting, so every node has a representative within a distance $r$ of it at level $r$.
  3. 3.
    Constraint 7 is because by non-contraction of the embedding, we have $x_{ij}^{r}=0$ if $j\notin B_{i}(r)$ so the left side is the probability that $j$ has some representative at level $r$, which happens with probability $1$.
  4. 4.
    Constraint 8 is satisfied by non-contraction of all embeddings, and constraints 9 and 10 are satisfied because all probabilities are in the range $[0,1]$.
0.B.2 Rounding the LP
We formally describe the rounding algorithm of Section 4 in Algorithm 3. This algorithm largely relies on the algorithms of [31] (presented here as Algorithms 4 and 5), but it takes the extra step of excluding from the domain of the output embedding all nodes in $X$ with optimal $\delta_{j}$ value above some threshold $\delta^{*}$. These excluded nodes form the outlier set for this embedding. Note that the outlier set is deterministic, as it is decided by a simple thresholding of the deterministic LP solution.
Algorithm 3 Algorithm for approximating outlier sets for HST embeddings
Input: Metric space $(X,d)$, target distortion $c$, approximation factor $\epsilon>0$
Output: An embedding $\alpha:S\rightarrow T$ for some HST $T$ and some $S\subseteq X$
1:for $k=1$ to $n$ do
2:  Let $LP_{k}$ be HST LP for $(X,d),c,k$ $\triangleright$ For weighted outlier sets, use $\min\sum_{i\in X}w_{i}\delta_{i}$ as the objective
3:  Find an optimal solution to $LP_{k}$ of value $v_{k}$
4:  Use the rounding algorithm of [31] (Algorithms 4 and 5) to sample an embedding $\alpha_{k}:X\rightarrow T$ for HST $T$
5:  $\delta^{*}\leftarrow\frac{\epsilon}{16\zeta\cdot\log^{2}k}$
6:  $K_{k}\leftarrow\set{j\in X:\delta_{j}\geq\delta^{*}\text{ in the optimal LP solution}}$
7:Let $k^{*}$ be the smallest value of $k$ with $v_{k}\leq k$ $\triangleright$ For weighted outlier sets, let $k^{*}$ minimize $v_{k}\log^{2}k$
8:Return $\alpha_{k^{*}}$ and $K_{k^{*}}$
Algorithm 4 (Algorithm 1 from [31]): Random selection of partitions for the nodes in $X$
1:for $r\in M$ in decreasing order do
2:  $S\leftarrow X$; $P_{i}^{r}\leftarrow\emptyset$ for all $i\in X$
3:  while $S\neq\emptyset$ do
4:   Choose a center $i\in X$ uniformly at random independent of past choices
5:   Choose $\ell_{i}^{r}\in[0,1]$ uniformly at random independently of past choices
6:   For each $j\in S\cap B_{i}(r)$, if $x_{ij}^{r}\geq\ell_{i}^{r}$, assign $j$ to $P_{i}^{r}$ and remove $j$ from $S$.
Algorithm 5 (Algorithm 2 from [31]): Embedding $(X,d)$ (with diameter $\Delta$) into an HST given randomly chosen partitions $P_{i}^{r}$ from Algorithm 4
1:Place a root node $w$ at the highest level with $\eta_{w}\leftarrow 2\cdot 2^{\lfloor\log\Delta\rfloor}$ and set $S_{w}\leftarrow X$
2:for $r\in M$ in decreasing order do
3:  for each node $w$ at the previous (parent) level with set $S_{w}$ do
4:   for each $i$ with $P_{i}^{r}\cap S_{w}\neq\emptyset$ do
5:     Place a child node $v$ with set $S_{v}\leftarrow P_{i}^{r}\cap S_{w}$ and $\eta_{v}\leftarrow 2r$
In order to show Algorithm 3 does in fact sample form a probabilistic $(O(\frac{k}{\epsilon}\log^{2}k),(32+\epsilon)c)$-outlier embedding, we will reference the following lemmas from [31].
Lemma 6(counterpart of Lemma 10 from [31])
The embedding produced by the rounding Algorithms 4 and 5 is non-contracting.
Lemma 7(counterpart of Lemma 11 from [31])
If $\delta_{j}$ is the optimal LP value of the variable $\delta_{j}$ and if $\alpha$ is sampled by Algorithms 4 and 5, we have $\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8\cdot(4+\zeta\cdot\log^{2}k\cdot(\delta_{j}+\delta_{j^{\prime}}))\cdot c\cdot d(j,j^{\prime})$
Note that Lemma 6 can be proven identically to [31], as the relevant algorithms and constraints involved in that proof are unchanged, but we briefly outline their reasoning here.
Proof of Lemma 6. As Algorithms 4 and 5 are unchanged from [31], we can use the same reasoning. In particular, for any pair of nodes $j,j^{\prime}$, if $j$ and $j^{\prime}$ are in the same part of the partition at the level corresponding to $r$, then there is some $i$ that is a distance at most $r$ from each of them. This means that the first level at which they share a parent is the level corresponding to $r$, and since we double all distances at the end of Algorithm 5 by scaling everything up by $2$, this implies non-contraction. ∎
Proof of Lemma 7. Lemma 11 of [31] stated that $\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8q$, where $q$ is the optimal LP solution value for their distortion variable. (Recall that for us, the distortion variable is replaced with a constant and we focus on our new outlier variables.) To prove this, they used only constraints 5 through 10 to show that $\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8\cdot\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}$ and then applied their version of constraint 4 to get the final result. As constraints 5 through 10 and the rounding algorithm remain unchanged in our algorithm, the statement $\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8\cdot\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}$ still holds. Thus, by our slightly modified version of constraint 4, we get $\mathop{\mathbb{E}}_{\alpha}[d_{\alpha}(j,j^{\prime})]\leq 8\cdot\sum_{r\in M}r\cdot\gamma_{jj^{\prime}}^{r}\leq 8\cdot(4+\zeta\cdot\log^{2}k\cdot(\delta_{j}+\delta_{j^{\prime}}))\cdot c\cdot d(j,j^{\prime})$, as stated in the lemma. ∎
Finally we can prove Theorem 1.1, which says that we can efficiently sample from an $(O(\frac{k}{\epsilon}\log^{2}k),(32+\epsilon)c)$ embedding.
Proof of Theorem 1.1. We know from Lemma 6 that the sampled embedding from Algorithm 3 is non-contracting. To consider its expected expansion, we need only consider pairs of nodes $j,j^{\prime}$ with $\delta_{j},\delta_{j^{\prime}}\leq\delta^{*}$. Then by Lemma 7, the expected distance between such a $j$ and $j^{\prime}$ is at most
$(32+8\zeta\cdot\log^{2}k\cdot 2\delta^{*})\cdot c\cdot d(j,j^{\prime})\leq(32+\epsilon)\cdot c\cdot d(j,j^{\prime}),$
by the choice of $\delta^{*}$ in Algorithm 3.
Additionally, note that the set of outliers produced by Algorithm 3 is deterministic and has size at most $k/\delta^{*}=O(\frac{k}{\epsilon}\log^{2}k)$, as Lemma 3 tells us that the optimal solution to HST LP has value at most $k$. ∎
0.B.2.1 A note on the runtime
Notably, Algorithm 4 (Algorithm 1 from [31]) runs in expected polynomial time, not deterministic polynomial time. Following the same reasoning as [31], we can select some additional error term $\epsilon^{\prime}$ and a sufficiently large constant $c^{\prime}$ and terminate the while loop after at most $c^{\prime}\cdot n\ln\frac{\Delta}{\epsilon^{\prime}}$ steps. If the algorithm would have otherwise exceeded that number of steps, we just output an HST in which all distances are $2^{\lceil\log\Delta\rceil}$. Using sufficient concentration bounds, the rounding scheme incurs a distortion at most $\epsilon^{\prime}$ higher than in the original form. If we want the distortion of the final embedding to be at most $(32+\epsilon)c$, we can select $\epsilon/2$ for $\epsilon^{\prime}$ in this section and for the approximation factor in choosing a threshold $\delta^{*}$. Note that because of this choice, the runtime of the algorithm depends on $1/\epsilon$ and $\Delta$ (specifically growing with $\ln(\frac{\Delta}{\epsilon})$), unlike the $\ell_{2}$ outlier embedding algorithm in [19].
0.B.3 Weighted outlier embeddings
In the weighted outlier set problem, we are given $\mathcal{X}=(X,d)$ and $c$, as well as a set of weights $w$ on the nodes, and our goal is to find a subset $K\subseteq X$ such that $X\setminus K\xhookrightarrow{c}_{p}\mathcal{H}$ and $w(K)$ is minimized. So far we have considered the unweighted version of this problem, in which $w(i)=1$ for all $i\in X$. In this subsection, we will discuss the slight changes needed to apply the results of Section 4 to the weighted outlier setting and thus obtain a bicriteria approximation for this problem.
The weights of the nodes affect the cost of leaving those nodes out, but they have no effect on how good the resulting embedding may be. This means that if we attempt to modify HST LP for the weighted setting, the weights should only influence the objective function. To solve this problem, we will use Algorithm 3 with the changes noted in red for weighted outlier embeddings. In the weighted setting, we are not necessarily concerned with simply minimizing the outlier set size, so intuitively we should select the LP solution corresponding to the $k$ that minimizes the objective value function for that LP. However, in order to account for the extra $\log^{2}k$ approximation factor incurred by our algorithm, we will instead look for a value of $k$ that minimizes the objective value of its LP solution, times $\log^{2}k$. The necessary changes are noted as comments in Algorithm 3. Below we give the proof of Corollary 1.
Proof of Corollary 1. Let $K$ be an outlier set of $\mathcal{X}$ that has size at most $\mathcal{K}(c)$, has weight at most $\text{OutlierCost}(c)$, and has $X\setminus K\xhookrightarrow{c}_{p}\mathcal{H}$. Define $k:=|K|$.
Let $k^{*}$ be the value chosen by the algorithm to minimize $v_{k^{\prime}}\log^{2}{k^{\prime}}$. Then
$v_{k^{*}}\cdot\log^{2}(k^{*})$ $\leq v_{k}\cdot\log^{2}k$
$\leq\text{OutlierCost}(c)\cdot\log^{2}\mathcal{K}(c),$
where the first line is by definition of $k^{*}$ and the second line is by the fact that $\delta_{i}=\begin{cases}1&i\in K\\ 0&else\end{cases}$ is part of a valid solution to HST LP with parameter $k$, and this solution has objective value at most $\text{OutlierCost}(c)$. (Note that this is part of a valid solution because only the objective function has changed, so all arguments with respect to feasibility of such a solution still go through.)
Now we consider the rounding step of the algorithm. We know that by selecting all $i$ with $\delta_{i}\geq\frac{\epsilon}{\zeta\cdot\log^{2}k^{*}}$ in the LP solution corresponding to $k^{*}$, we obtain a solution with at most $(32+\epsilon)c$ distortion on the non-outliers (again following identical reasoning to that in the unweighted case). Thus, we are just left with bounding the weight of the outlier set. We have that the total cost is at most $v_{k^{*}}\cdot\frac{\zeta\cdot\log^{2}(k^{*})}{\epsilon}$ as a result of our rounding, as $v_{k^{*}}$ is the value of the objective function under this solution. This implies an outlier set of weight at most $O(\frac{1}{\epsilon}\cdot\text{OutlierCost}(c)\cdot\log^{2}\mathcal{K}(c))$. ∎
Appendix 0.C Probabilistic Lipschitz extensions
While the primary focus of this paper is on strong nested embeddings and their applications to finding outlier HST embeddings, in this section we will also briefly mention how an “oblivious” Lipschitz extension algorithm for deterministic embeddings can also serve as a Lipschitz extension algorithm for probabilistic embeddings. In particular, we consider the case that we have some “oblivious” algorithm that takes in a deterministic embedding $\alpha:S\xhookrightarrow{c_{S}}Y$ for $S\subseteq X$ and samples from a probabilistic Lipschitz extension of that embedding. We will show here that if we have a probabilistic embedding $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$ instead, we can sample any embedding from $\mathcal{D}_{S}$ and run the existing Lipschitz extension algorithm on it in order to obtain a sample from a Lipschitz extension of $\mathcal{D}_{S}$. Notably, the original Lipschitz extension (and corresponding algorithm) must be oblivious, in that for each $x\notin S$, the extension always maps $x$ to the same spot as some $u\in S$, and the choice of such $u$ is completely independent of the input embedding $\alpha_{S}$.
We first define probabilistic and oblivious Lipschitz extensions.
Definition 5
A probabilistic Lipschitz extension of a deterministic embedding with extension factor $L$ of a deterministic embedding $\alpha_{S}:S\xhookrightarrow{c_{S}}Y$ for some space $Y$ and $S\subseteq X$ is a probabilistic embedding $\mathcal{D}:X\xhookrightarrow{}_{p}Y$ such that for all $\alpha$ in the support of $\mathcal{D}$ and all $u\in S$, $\alpha(u)=\alpha_{S}(u)$; and for all $x,y\in X$ $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(x,y)]\leq L\cdot c_{S}\cdot d(x,y)$.
Further, we say that a function $f$ mapping input embeddings $\alpha_{S}$ to Lipschitz extensions $\mathcal{D}$ is anoblivious Lipschitz extension function if for all $x\in X\setminus S$, $\alpha(x)\in\set{\alpha(u):u\in S}$ with probability $1$, and for all $u\in S$, $\Pr_{\alpha\sim f(\alpha_{S})}[\alpha(x)=\alpha(u)]$ is the same for all inputs $\alpha_{S}$.
Notably, Lipschitz extensions $\alpha$ on deterministic embeddings $\alpha_{S}$ require that $\alpha(u)=\alpha_{S}(u)$ for all $u\in S$ with probability $1$. In the randomized setting, $u$ is not mapped to a fixed place in a metric space, so we need to clarify what we mean by a Lipschitz extension. In particular, we will require not only that the distribution of $u$ under the new embedding be the same as under the old embedding, but also that the joint distribution of the embedding on all points in $S$ is unchanged by the new embedding. Because of this, we get that the output embedding has the same expected distortion on any pair of nodes as under the original embedding, but in fact this is a much stricter requirement.
We define Lipschitz extensions of probabilistic embeddings below.
Definition 6
Let $(X,d_{X})$ be a metric space, let $\mathcal{Y}$ be a set of metric spaces, and $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$ for $S\subseteq X$. Then $\mathcal{D}:X\xhookrightarrow{}_{p}\mathcal{Y}$ is a Lipschitz extension of $\mathcal{D}_{S}$ with extension factor $L$ if:
  1. 1.
    The distribution of $(\alpha(x_{i}))_{x_{i}\in S}$ is unchanged under $\mathcal{D}$ compared to $\mathcal{D}_{S}$;
  2. 2.
    For all $x,y\in X$,
    $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(x,y)]$ $\leq L\cdot c_{S}\cdot d_{X}(x,y).$
    In other words, the expected expansion under $\mathcal{D}$ should be bounded by $L\cdot c_{S}$.
We now argue that if there are oblivious probabilistic Lipschitz extension functions of deterministic embeddings, they imply the existence of Lipschitz extensions of probabilistic embeddings.
Theorem 0.C.1
Let $f$ be an oblivious probabilistic Lipschitz extension function that maps deterministic embeddings $\alpha_{S}:S\xhookrightarrow{c_{S}}Y$ for any $Y\in\mathcal{Y}\cup\set{S}$ to embeddings on a set $X$ with extension factor $L$.
Then consider $\mathcal{D}_{S}:S\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$ and let $g$ be an algorithm that samples $\alpha_{S}\sim\mathcal{D}_{S}$ and outputs $f(\alpha_{S})$. Let $\mathcal{D}^{*}$ be the distribution produced by running $g(\mathcal{D}_{S})$ and sampling an embedding from the resulting distribution. Then $\mathcal{D}^{*}$ is a Lipschitz extension of $\mathcal{D}_{S}$ with extension factor $L$.
Proof
Note that we can sample from $\mathcal{D}^{*}$ by sampling $\alpha_{S}$ from $\mathcal{D}_{S}$ and then sampling from $f(\alpha_{S})$. Further, for any $\alpha$ in the support of $f(\alpha_{S})$, $(\alpha(u))_{u\in S}$ is fixed to be exactly $(\alpha_{S}(u))_{u\in S}$. Thus, $\mathcal{D}^{*}$ has the same distribution on $S$ as does $\mathcal{D}_{S}$.
Now we bound expansion for an arbitrary pair of nodes $x,y\in X\setminus S$. Define $K_{u}$ to be the random set equal to $\set{x\in X\setminus S:\alpha(x)=\alpha_{S}(u)}$ when $\alpha$ is drawn from $f(\alpha_{S})$. Note that for $u\neq v\in S$, $\alpha_{S}(u)\neq\alpha_{S}(v)$ because $\alpha_{S}$ is expanding. This implies that the random sets $K_{u}$ form a random partition of $X\setminus S$.
Further, we get that the probability of any particular partition $\set{K_{u}}$ occurring is the same for all distributions in the image of $f$, as $f$ produces an oblivious Lipschitz extension and thus the input to $f$ cannot affect these probabilities.
Let $u_{x}$ be the random variable corresponding to the $u\in S$ such that $\alpha(x)=\alpha_{S}(u)$. Then we get
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)]$ $=\mathop{\mathbb{E}}_{\set{K_{u}}}[\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}^{*}}[d_{\alpha}(x,y)|\set{K_{u}}$]]
$=\mathop{\mathbb{E}}_{\set{K_{u}}}[\mathop{\mathbb{E}}_{\alpha_{S}\sim\mathcal{D}_{S}}[d_{\alpha_{S}}(u_{x},u_{y})|\set{K_{u}}$]]
$\leq\mathop{\mathbb{E}}_{\set{K_{u}}}[c_{S}\cdot d(u_{x},u_{y})]$
$\leq c_{S}\cdot L\cdot d(x,y).$
The first line comes from the law of total probability. The second line comes from the definition of $u_{x}$ and $u_{y}$. The third line comes from the fact that if $\set{K_{u}}$ is fixed, then $u_{x}$ and $u_{y}$ are fixed, and the fact that the expected distortion between any pair of nodes in $S$ is $c_{S}$ under $\alpha_{S}$. Note that we were able to take the expectation over $\set{K_{u}}$ without setting the choice of $\alpha_{S}\sim\mathcal{D}_{S}$ in the first line due to the obliviousness of the Lipschitz extension.
To see correctness of the fourth line, let $\iota$ be the identity map from $S$ to $S$. Note that $\mathop{\mathbb{E}}_{\set{K_{u}}}[d(u_{x},u_{y})]=\mathop{\mathbb{E}}_{\alpha\sim f(\iota)}[d_{\alpha}(x,y)]$, and $f(\iota)$ is a probabilistic Lipschitz extension of the identity function with extension factor $L$. (Further, $\iota$ is in the domain of $f$ because it maps $S$ to $S\in\mathcal{Y}\cup\set{S}$.) ∎
We note that it is essential that the probabilistic Lipschitz extension of the deterministic embedding used in the statement of the above lemma be oblivious. If this is not the case, then sampling from $\mathcal{D}_{S}$ and then applying the extension to the resulting embedding could create a distribution with very high expansion on pairs that are not in $S$. In particular, the worst expansion between any pair of nodes in the drawn embedding can be quite high with high probability, even if the average expansion for any given pair is small. The definition of a Lipschitz extension of a deterministic embedding only requires that the final expansion in the new embedding be bounded with respect to the worst case expansion for any pair in the original embedding. Thus, if we try to apply the arguments in Lemma 0.C.1 to a non-oblivious extension, we may not obtain the same result.
Further, we note that Algorithm 4.1 in [19] in fact takes in any embedding $\alpha_{S}:S\xhookrightarrow{c_{S}}Y$ for any metric space $Y$ and samples from an oblivious probabilistic Lipschitz extension of $\alpha_{S}$ with extension factor $125H_{k}$, where $k:=|X\setminus S|$. Thus, Theorem 0.C.1 implies that for any $\mathcal{Y}$ and any $\mathcal{D}_{S}\xhookrightarrow{c_{S}}_{p}\mathcal{Y}$, there exists a Lipschitz extension $\mathcal{D}$ of $\mathcal{D}_{S}$ with extension factor $125H_{k}$.
Appendix 0.D Ultrametrics
We can also consider how our results affect embeddings into another popular class of metrics called ultrametrics, which are defined as follows.
Definition 7
A metric $(X,d)$ is an ultrametric if for all $x,y,z\in X$, $d(x,y)\leq\max\set{d(x,z),d(z,y)}$.
Note that HSTs form a subset of the set of ultrametrics. Further, because HSTs are a subset of all ultrametrics and any ultrametric embeds into an HST with distortion at most $2$ [12], we also obtain the following corollary to Theorem 1.1.
Corollary 2
Given a metric $(X,d)$ and target distortion $c$, let $k$ be the optimal outlier set size for probabilistically embedding $(X,d)$ into ultrametrics. Then for any constant $\epsilon>0$, there exists an algorithm that efficiently samples from a probabilistic embedding of some $S\subseteq X$ into ultrametrics with $|X\setminus S|\leq O(\frac{1}{\epsilon}k\log^{2}k)$ and distortion at most $(64+\epsilon)c$.
Appendix 0.E More applications
In Section 5 we discussed the Minimum Cost Communication Tree (MCCT) problem, and we stated that there was an algorithm for approximating the outlier version of the problem when the input is a tree. We begin our applications section by presenting that algorithm and discussing how we can combine it with our outlier embedding algorithm. [^5] An intuitive approach here might be to greedily remove the node $i$ maximizing $\sum_{j}d(i,j)$, but this does not get a finite approximation ratio, even for trees.
0.E.1 MCCT with outliers
Let $(T,d)$ be a tree metric, and let $k$ be an integer. Then we define variables $x_{ij}$ for each pair of nodes $ij$ and variables $\delta_{i}$ for each variable $i$. We can use them to make the following linear program.
$\min\sum_{i,j\in T}x_{ij}\cdot r_{ij}\cdot d(i,j)\quad s.t.$ (MCCT Outlier LP)
$\forall i,j\in T:\quad$ $\delta_{i}+\delta_{j}+x_{ij}\geq 1$
$\sum_{i\in T}\delta_{i}\leq k$
$\forall i,j\in T:\quad$ $\delta_{i},x_{ij}\in[0,1]$
Claim 2
Let $(T,d)$ be a tree metric such that there exists a non-contracting embedding $\alpha:T\rightarrow\mathcal{H}$ with a set $K\subseteq T$, $|K|\leq k$, and $\sum_{i,j\in T\setminus K}r_{ij}\cdot d_{\alpha}(i,j)\leq D$.
Then the following procedure finds an embedding $\alpha^{\prime}$ and subset $K^{\prime}\subseteq T$ with $|K^{\prime}|\leq 3k$ and $\sum_{i,j\in T\setminus K^{\prime}}d_{\alpha^{\prime}}(i,j)\leq 24D$.
  1. 1.
    Solve MCCT Outlier LP and let $x_{ij},\delta_{j}$ be the resulting values
  2. 2.
    Let $K^{\prime}=\set{i:\delta_{i}\geq 1/3}$
  3. 3.
    Embed $T\setminus K^{\prime}$ into a tree $T^{\prime}$ with distortion at most $8$, using the algorithm of Gupta [25]
Proof
First, MCCT Outlier LP has a solution of value at most $D$. In particular, let $K$ be as in the claim statement and consider setting $\delta_{i}=\begin{cases}1&i\in K\\ 0&else\end{cases}$ and $x_{ij}=\begin{cases}1&\delta_{i}=\delta_{j}=0\\ 0&else\end{cases}$. As there are only $k$ outliers, the second constraint is clearly satisfied, and we have defined $x_{ij}$ such that the other constraints are also satisfied. Further, $x_{ij}$ is 0 if either $i$ or $j$ is an outlier in $K$, so the objective function value is just $\sum_{i,j\in T\setminus K}r_{ij}d(i,j)$, which by the claim statement is bounded by $D$.
Let $K^{\prime}$ be as in the claim. As the $\delta_{i}$ sum to at most $k$, we know $|K^{\prime}|\leq 3k$. Consider any $x_{ij}$ with $i,j\notin K^{\prime}$ We have $x_{ij}\geq 1-\delta_{i}-\delta_{j}\geq 1/3$ by how we chose $K^{\prime}$. The cost associated with a given pair $ij$ of non-outliers in our solution is $r_{ij}d(i,j)$, but that same pair contributed at least $r_{ij}d(i,j)/3$ to the optimal objective function value, so our final cost is at most $3$ times the objective, which is at most $3D$. At this point, we have an isometric (distortion $1$) embedding of the metric induced by $T\setminus K^{\prime}$ into the original metric $T$ itself. Thus, we can view $K^{\prime}$ as a set of Steiner points in our original metric.
Finally we consider the loss incurred applying the algorithm of Gupta [25]. This creates a tree embedding that removes Steiner points and incurs distortion at most $8$, so the cost of this embedding on $T\setminus K^{\prime}$ is at most $24D$, as desired. This step is what prevents us from applying this algorithm directly to general metrics. ∎
Note additionally that if $(T,d)$ is an HST metric, then we can “remove” the bad points we’ve found by just deleting the leaves corresponding to those nodes, and as a result we do not incur the extra factor of $8$ on the distortion.
Converting to an approximation on general metrics
In Section 5, we argued that we can use the outlier embeddings algorithm from this paper to obtain a solution with at most $O((1+\epsilon)c_{k}(\mathcal{X})(OPT_{k}))$ cost and $O(\frac{k}{\epsilon}\log^{2}k)$ outliers, where $c_{k}(\mathcal{X})$ is the minimum distortion for embedding $\mathcal{X}$ into HSTs with at most $k$ outliers and $OPT_{k}$ is the cost of the optimal solution that is allowed to select $k$ outliers. We will discuss this in additional detail here.
In particular, assume that an optimal algorithm for the $k$-outlier MCCT problem removes a subset $K^{*}\subseteq X$. Let $S=X\setminus K^{*}$. The cost of this algorithm, which is $OPT_{k}$, is at least $\sum_{i,j\in S}r_{ij}d(i,j)$. Let $OPT_{\alpha,k}$ be the cost of the optimal $k$-outlier solution on the metric $d_{\alpha}$ (i.e. the optimal solution after the embedding $\alpha$ is applied). We get that if $\mathcal{D}$ is a probabilistic embedding with expected distortion at most $c$, then $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[OPT_{\alpha,k}]\leq E_{\alpha}[\sum_{i,j\in S}r_{ij}d_{\alpha}(i,j)]\leq c\sum_{i,j\in S}r_{ij}d(i,j)\leq 8c\cdot OPT_{k}$, where we used the fact that the optimal cost for the $k$-outlier problem on $d_{\alpha}$ is bounded by the cost of embedding all points but $K^{*}$ into a tree, and we applied the Steiner point removals results of Gupta [25]. Consider probabilistically embedding a given metric into trees using $\mathcal{D}$ and then applying the outlier approximation algorithm from Claim 2. For an input metric $d_{\alpha}$, this algorithm produces a solution with at most $3k$ outliers and cost at most $3OPT_{\alpha,k}$. Thus, the expected cost of this solution is at most $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[3OPT_{\alpha,k}]\leq 24c\cdot OPT_{k}$, where $c$ is the distortion of embedding $\mathcal{D}$. This is enough to obtain the results we claimed in Section 5.
The outlier embedding results from this paper let us improve the factor $c$ that we can obtain for such an approximation if we are willing to increase the number of outliers. Note also that because $\epsilon$ is a tunable parameter in our algorithm, it does not necessarily need to be a constant, so if we are more interested in minimizing the number of outliers, we can also obtain a solution of cost at most $O(\log^{2}k\cdot c_{k}(\mathcal{X})\cdot OPT_{k})$ with only $O(k)$ outliers.
0.E.2 Some metrics have small outlier sets
To motivate this work, we may ask: do there exist metrics for which removing a small number of points results in a significantly better probabilistic embedding into HSTs? In this subsection, we discuss types of metrics for which this is the case. In particular, consider the distance metric on an unweighted graph obtained by “linking” two unweighted graphs through an edge, the first graph being an expander on $n$ nodes and the second being the complete graph on $n-\log n$ nodes. Bartal [10] shows that, for any $x$, certain expander graphs on $x$ nodes incur a distortion $\Omega(\log x)$ when probabilistically embedding into HSTs. Thus, the graph we have described requires $\Omega(\log\log n)$ distortion for HST embeddings (inheriting the distortion bound of the expander subgraph), whereas the same graph with the expander removed can be embedded with $O(1)$ distortion.
This graph is highly specific and the “good” and “bad” parts of the graph are easy to spot. We will use this graph as an example in the coming subsections because its simplicity makes it easy to analyze, but we also note that the class of metrics whose optimal distortion can be improved by removing a small number of vertices is much more diverse than this simple example.
To demonstrate a large family of metrics with good outlier embeddings, we will turn to the notion of metric composition as defined by Bartal et al. [12].
Definition 8(Metric composition [12])
Let $(M,d_{M})$ be a metric space with minimum distance exactly $1$, and let $\mathcal{N}=\set{(N_{x},d_{N_{x}})}_{x\in M}$ be a set of metric spaces. Let $\Delta:=\max_{x\in M}\set{diameter(N_{x})}$. Then the $\beta$-composition of $M$ and $\set{N_{x}}$ denoted $M^{*}=M_{\beta}[\mathcal{N}]$ is a metric on the vertex set $\set{(x,u):x\in M,u\in N_{x}}$ such that distances in $M^{*}$ are
$d_{M^{*}}((x,u),(y,v)):=\begin{cases}d_{N_{x}}(u,v)&x=y\\ \beta\cdot\Delta\cdot d_{M}(x,y)&else\end{cases}.$
We provide a visualization of this composition in Figure 2. Bartal et al. [12] show that for any $M$ and $\mathcal{N}$, $M_{\beta}[\mathcal{N}]$ defines a proper metric if $\beta\geq 1/2$. While it’s easy to factor a composition back into its original components if $\beta$ is quite large, it’s less clear how to do so when $\beta=1/2$.
$N_{x}$$N_{y}$$N_{z}$$\beta\Delta\cdot d_{M}(x,y)$$\beta\Delta\cdot d_{M}(y,z)$$\beta\Delta\cdot d_{M}(x,z)$$\Delta:=\max_{w\in\{x,y,z\}}(diameter(N_{w}))$
Figure 2: A visual example of metric composition where $M$ is some metric on the set $\set{x,y,z}$.
The highest distortion required to embed a metric in $\set{M}\cup\mathcal{N}$ into HSTs is a lower bound on the minimum distortion to embed $M_{\beta}[\mathcal{N}]$ into HSTs, and careful composition reveals that in fact this quantity is an upper bound as well (up to constant factors).
0.E.3 Buy-at-bulk
In the (uniform) buy-at-bulk problem, [^6] We assume the metric version of the problem rather than the graph version, as Awerbuch et al. [6] note that they are equivalent. we are given a metric space $\mathcal{X}$ as well as a set of demand pairs $(s_{i},t_{i})$, each with a demand $d_{i}$ and a sub-additive cost function $g$. (We will assume wlog that $g(1)=1$, as this is always possible by rescaling.) The goal is to select a single path $P_{i}$ for each demand pair in a way that minimizes $\sum_{e}\ell_{e}g(\sum_{i:e\in P_{i}}d_{i})$, where $\ell_{e}$ is the length of edge $e$ and all the sum is over all edges. This has natural applications in many areas of operation research.
A key result of Awerbuch et al. [6] states that if $\mathcal{D}$ is a probabilistic tree embedding of metric $(X,d)$ with distortion $c$, then the optimal cost of the buy-at-bulk problem on $(X,d)$ is at most a factor of $c$ less than the expected value of the optimal solution to the same problem on the embedded metric. The buy-at-bulk problem is trivial on trees, where there is only one path between each pair of nodes. Thus, using the results of Munagala et al. [31] we can obtain an expected approximation ratio $O(c)$, where $c$ is the optimal distortion for probabilistically embedding the input into HSTs.
In this subsection, we will work to improve this approximation ratio for certain metrics by solving the problem separately on a set of requests only involving a set of points with a low distortion HST embedding and then on the remaining set of requests, which we seek to ensure has relatively small total cost. We formalize this idea in Algorithm 6.777Algorithm 6 actually gives a more general framework for solving the buy-at-bulk or dial-a-ride problem, and as a result it asks for an oracle that solves the underlying problem on HSTs. For the buy-at-bulk problem on trees, the only feasible solution is optimal, so $\mathcal{A}$ is trivial. To outline this approach, we must first define some terms that will help us set up a “weighted outlier” problem on the input metric space.
In particular, consider a buy-at-bulk instance $(\mathcal{X},\set{(s_{i},t_{i},d_{i})}_{i=1}^{m})$. Let $OPT_{i}$ be the cost of the optimal solution to the buy-at-bulk instance $(\mathcal{X},(s_{i},t_{i},d_{i}))$ (i.e. the same instance but with only demand pair $i$). This is easy to solve optimally, as we can simply pick a shortest $s_{i}$-$t_{i}$ path and purchase $d_{i}$ capacity on that path. We will consider $OPT_{i}$ to be the “cost” of demand pair $i$. In order to decide the “cost” of removing a node $x$ from $\mathcal{X}$, we will sum the costs of all demand pairs that involve $x$. That is, let $R_{x}:=\set{i:x\in\set{s_{i},t_{i}}}$, and we will set the weight of $x$ to be $w(x):=\sum_{i\in R_{x}}OPT_{i}$. Further, for any $K\subseteq X$, let $w(K)=\sum_{x\in K}w(x)$. This weighting of the points of $\mathcal{X}$ sets up a weighted outlier problem for $\mathcal{X}$.
For a given metric $\mathcal{X}$ with weights $w$ and any choice of $c\geq 1$, we will define $\text{OutlierCost}(c):=\min_{K\subseteq X;(X\setminus K)\xhookrightarrow{c}_{p}\mathcal{H}}w(K)$ to be the cost of the cheapest set we can leave out while still embedding the remaining metric into HSTs with at most $c$ distortion. We will also be interested in the size of such a cheap set, so we define $\mathcal{K}(c):=|\arg\min_{K\subseteq X;(X\setminus K)\xhookrightarrow{c}_{p}\mathcal{H}}w(K)|$. In Claim 3, we give a bound on the cost of Algorithm 6. Note that for some constant $\eta$, if $c=\eta\log n$, [^8] In this section, we let $\eta$ be a constant such that for all $n$-point metric spaces $X$, $X\xhookrightarrow{\eta\log n}_{p}\mathcal{H}$. (Such a constant is guaranteed to exist by [23].) we have $\text{OutlierCost}(c)=0$, so Algorithm 6 never obtains worse than an $O(\log n)$ approximation.
Algorithm 6 Approximation algorithm for dial-a-ride or buy-at-bulk
Input: metric $\mathcal{X}$, request set $\set{(s_{i},t_{i})}_{i=1}^{m}$, approximation algorithm $\mathcal{A}$ for buy-at-bulk or dial-a-ride on HSTs, integer $\ell$ if dial-a-ride and demands $\set{d_{i}}$ if buy-at-bulk
Output: A buy-at-bulk or dial-a-ride solution
1:for $x=1$ to $n$ do
2:  $R_{x}\leftarrow\set{i:x\in\set{s_{i},t_{i}}}$
3:  Let $OPT_{i}$ be the cost of an optimal solution that only fulfills only the $i$th request
4:  $w_{x}\leftarrow\sum_{i\in R_{x}}OPT_{i}$
5:for $i=0$ to $\log\log n$ do
6:  Let $c_{i}\leftarrow\eta\log n/2^{i}$ $\triangleright$ here $\eta$ is the appropriate constant from [23]
7:  Run the weighted version of Algorithm 3 with input distortion $c_{i}$ and $\epsilon=1$
8:  Receive an outlier set $K_{i}$ of cost $w^{i}$ and an embedding $\alpha_{i}$ of distortion $2c_{i}$
9:  Embed $X\setminus K_{i}$ into HSTs
10:  Use $\mathcal{A}$ to solve the relevant problem on requests $R_{1}:=\set{j:\set{s_{j},t_{j}}\cap K_{i}=\emptyset}$ and nodes in $X\setminus K_{i}$
11:  Let $sol_{i}^{1}$ be the solution obtained
12:  Let $R_{2}:=\set{j:\set{s_{j},t_{j}}\cap K_{i}\neq\emptyset}=R\setminus R_{1}$
13:  Find a minimum distortion embedding of $X_{i}=K_{i}\cup\set{x:\exists y\in K_{i}\text{ s.t. }\set{x,y}\in R}$ into HSTs
14:  Use $\mathcal{A}$ to find a solution to the problem with node set $X_{i}$ and request set $R_{2}$
15:  Let $sol_{i}^{2}$ be the solution obtained
16:  Find the “naive” solution to the problem with node set $X_{i}$ and request set $R_{2}$
17:  Let $sol_{i}^{3}$ be the solution obtained
18:  Let $sol_{i}$ be the solution obtained by combining solution $sol_{i}^{1}$ with either $sol_{i}^{2}$ or $sol_{i}^{3}$, whichever is cheaper
19:Return the solution of lowest cost
Claim 3
Let $(\mathcal{X},R=\set{(s_{i},t_{i},d_{i})}_{i=1}^{m})$ be an instance of the buy-at-bulk problem with an overall solution of cost at most $OPT$.
Then Algorithm 6 on this instance (and with the trivial oracle $\mathcal{A}$ for buy-at-bulk) returns a solution of cost at most
[…]
In fact, for any $c$ directly considered by the algorithm in line 6, the algorithm’s cost is bounded by $O({c\cdot OPT+\text{OutlierCost}(c)\cdot\log^{2}(\mathcal{K}(c))}$.
First we briefly justify why proving the second part of the statement is sufficient to prove the overall statement, which minimizes over all $c$ not just the $c_{i}$ considered by the algorithm. First, for all $c\geq\eta\log n=c_{0}$, $\text{OutlierCost}(c)=0$, so we need only justify this for $c<\eta\log n$. Let $c$ be the value that minimizes $c\cdot OPT+\text{OutlierCost}(c)\log^{2}(n)$, and note that any $c_{i}>c$ has optimal outlier set value at most equal to that of $c$. As the algorithm will consider some value of $c_{i}$ that is bigger than $c$ and at most $\eta$ times its value (if $c\geq\eta/2$, it will in fact consider a choice of $c_{i}$ that is at most twice this value), we will end up with the desired approximation.
Proof of Claim 3. We show that if index $i$ is selected as the optimal solution by the algorithm, then the cost of the solution is at most $O(c_{i}\cdot OPT+\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i})))$. Let $K_{i}$ be the outlier set from line 8 of Algorithm 6.
Let $S_{1}$ be the set of pairs with neither endpoint in $K_{i}$ and let $S_{2}$ be the remaining pairs. We know that the cost of the solution we obtain for pairs in $S_{1}$ is at most $c_{i}\cdot OPT$, as the tree embedding incurs distortion $c_{i}$ and we get an optimal solution for the resulting tree.
Now consider the cost associated with our purchase of edges for pairs in $S_{2}$. The total cost of purchasing for each demand pair independently is at most $\sum_{i\in S_{2}}OPT_{i}$. (As the cost function is sub-additive, buying more capacity on a single edge never costs more than splitting that capacity into two parts and buying them separately.) We know that $\sum_{i\in S_{2}}OPT_{i}=O(\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i})))$ by Algorithm 3’s guarantees on the outlier set cost. ∎
An example where the approximation ratio improves
Let $n$ be such that $\log n$ is an even integer dividing $n$. Let $M$ be an HST metric of size $n/\log n$ and diameter $\ell\geq\log^{4}n$, and let $\mathcal{N}=\set{N_{x}}_{x\in M}$ be a set of HST metrics, each of size $\log n$. Let $\mathcal{N}^{\prime}=\set{N_{x}^{\prime}}_{x\in M}$ be the same as $\mathcal{N}$, but for one $x^{*}\in M$, replace $N_{x^{*}}$ with the metric of an unweighted expander graph that does not embed well into HSTs. Let $M^{*}:=M_{1/2}[\mathcal{N}^{\prime}]$ be a metric of size $n$.
Consider an instance of the buy-at-bulk problem on $M^{*}$ defined as follows
  • For each $x\in M$, divide the nodes of $N_{x}^{\prime}$ into pairs $(u,v)$, and make $((x,u),(x,v))$ a demand pair in $M^{*}$ with demand $1$.
  • Let $x_{1},x_{2}\neq x^{*}$ be a pair of nodes in $M$ with distance $\ell$. (We assume $x^{*}$ is chosen such that such a pair exists.) Select some $u\in N_{x_{1}}$ and $v\in N_{x_{2}}$ and add $((x_{1},u),(x_{2},v))$ as a demand pair with demand $1$.
Note that $M^{*}$ requires distortion $\Omega(\log\log n)$ to probabilistically embed into HSTs (due to the presence of an expander graph of size $\log n$ as a submetric), so if we run the algorithm of Munagala et al. [31] and buy paths on the resulting tree, we could potentially pay as much as $\Omega(\log\log n\cdot OPT)$, depending on which distances were distorted.
Now consider the result of running Algorithm 6 on this instance. Note that if we remove from $M^{*}$ the nodes corresponding to $N_{x^{*}}^{\prime}$, we get a metric that embeds into HSTs with constant distortion. We also know that the outlier cost of the nodes corresponding to $N_{x^{*}}^{\prime}$ is at most $\log^{2}n$, as there are $O(\log n)$ demand pairs associated with these nodes; each such pair has demand $1$; nodes in $N_{x^{*}}^{\prime}$ only participate in demand pairs with other nodes in $N_{x^{*}}^{\prime}$; and the diameter of $N_{x^{*}}^{\prime}$ is at most $\log n$ (due to being the path metric of an unweighted graph on $\log n$ nodes). Thus, Algorithm 6 gives us a solution of cost at most $O(OPT+\log^{4}n)=O(OPT)$, where we used the fact that $OPT\geq\ell\geq\log^{4}n$, as there is some demand pair with distance at least $\ell$.
0.E.4 Dial-a-ride
In the dial-a-ride problem, we have a van of capacity $\ell$ and we are given a metric space $(X,d)$ as well as a set of transportation requests $\set{(s_{i},t_{i})}_{i=1}^{m}$ from $m$ people such that person $i$ wants the van to transport them from location $s_{i}$ to location $t_{i}$. The goal of the problem is to find a schedule of pickups and drop-offs for the van such that: [^9] Technically we should also require that the tour start and end at some specified “depot” location, but solving this relaxed version of the problem and then adding on beginning and ending drops from/to the depot only worsens the overall approximation ratio by a constant factor (as each of the new trips we add has cost at most $OPT$), so we will discuss this simplified version instead.
  1. 1.
    The number of passengers in the van is the number that have been picked up minus the number that have been dropped off, and it is always between 0 and $\ell$,
  2. 2.
    For any customer $i$, that customer is picked up at their starting location $s_{i}$ and dropped off later at their destination $t_{i}$, and
  3. 3.
    The cost of the solution is the sum of distances traveled by the van between each consecutive stop on its schedule.
This problem is NP-hard, as it generalizes the Traveling Salesman Problem [18]. Note that if we serve only a single request $i$, the problem is easy and incurs cost $d(s_{i},t_{i})$ in the optimal solution. Thus, computing $OPT_{i}$ for each request $i$ is efficient, and we can run Algorithm 6 for this problem.
However, we will need to use a more complicated algorithm $\mathcal{A}$ to solve the dial-a-ride problem on HSTs than we did for buy-at-bulk. In particular, we will consider an algorithm $\mathcal{A}$ that gets approximation ratio $r_{n,m,\ell}$ for dial-a-ride instances on HSTs with $n$ nodes, $m$ requests, and $\ell$ van capacity. In Claim 4, we consider the approximation ratio achieved by Algorithm 6 with such an input $\mathcal{A}$. Again note that for $c=\eta\cdot\log n$, $\text{OutlierCost}(c)=0$, so the worst-case approximation ratio is $O(\log n)$.
Claim 4
Let $\mathcal{A}$ be an approximation algorithm for the dial-a-ride problem on HSTs that achieves approximation ratio $r_{n,m,\ell}$ for instances of $n$ nodes, $m$ requests, and $\ell$ van capacity.
Let $(\mathcal{X},\set{(s_{i},t_{i})}_{i=1}^{m},\ell)$ be a dial-a-ride input such that $OPT$ is the cost of an optimal solution for this instance, $\Delta$ is the diameter of $\mathcal{X}$, and $\text{OutlierCost}(c)$ and $\mathcal{K}(c)$ are as in Section 0.E.3. Then Algorithm 6 achieves a feasible solution with expected cost at most
[…]
In fact, for any $c$ directly considered by the algorithm in line 6, the algorithm’s cost is bounded by
$O({c\cdot OPT+\text{OutlierCost}(c)\cdot\log^{2}(\mathcal{K}(c))}$.
Note that this algorithm’s performance is generally better on instances whose optimal cost is much bigger than $\Delta$ (such as instances in which the total number of requests is large compared to $\ell\Delta$).
Charikar et al. [16] give an algorithm $\mathcal{A}$ for dial-a-ride on HSTs with $r_{n,m,\ell}=O(\sqrt{\ell})$. Thus, we get an algorithm that produces a solution such that for any $c\geq 1$, the solution’s cost is at most $O(\sqrt{\ell}\cdot c\cdot OPT+\Delta\cdot\text{OutlierCost}(c)\log^{2}n)$. If each location is involved in at most $t$ requests (i.e. the demand is “spread out”), then the solution also has cost at most $O(\sqrt{\ell}\cdot(c+\log(t\cdot\mathcal{K}(c)))\cdot OPT)$.
As in Claim 3, to prove Claim 4, we only need to prove the given bound for each value of $i$ tested by the algorithm.
Proof of Claim 4. We will show a slightly stronger version of the claim by showing that for each $i$, the cost associated with the solution for $c_{i}$ is bounded by $O(r_{n,m,\ell}\cdot c_{i}\cdot OPT+\text{OutlierCost}(c_{i})\log^{2}\mathcal{K}(c_{i})\cdot\Delta)$ and also by $O(r_{n,m,\ell}\cdot c_{i}+r_{2|R_{2}^{i}|,m,\ell}\log(|R_{2}^{i}|)\cdot OPT)$, where $R_{2}^{i}$ is the set of requests that intersect $K_{i}$ in iteration $i$ of the algorithm. (Only the first of these two bounds will be necessary to show the claim.) For the rest of this proof we will fix a choice of $i$ and consider the solution the algorithm obtains for that choice of $i$ and we refer to $R_{2}^{i}$ as $R_{2}$.
Let $K_{i}$ be the outlier set from line 8 of Algorithm 6, let $R_{1}$ be the set of requests $(s_{j},t_{j})$ that contain a point in $K_{i}$, and recall that $R_{2}$ is the set of remaining requests. Note that each solution the algorithm finds is feasible, as we handle all requests in $R_{1}$ according to some feasible solution given by $\mathcal{A}$, and then we handle all requests in $R_{2}$ according to another feasible solution from $\mathcal{A}$ or by a naive algorithm for this set of requests.
We will divide the cost of the algorithm’s solution into three sections: the cost of servicing requests in $R_{1}$, the cost of servicing requests in $R_{2}$, and the cost of moving the van from the last drop-off location in the $R_{1}$ solution to the first pickup location in the $R_{2}$ solution.
  1. 1.
    First consider the cost of the solution we obtain for servicing requests in $R_{1}$. Our expected cost is at most $r_{n,m,\ell}$ times the optimum solution on the tree this solution is embedded in, and the expected distortion of that embedding is bounded by $2c_{i}$ (as we potentially lose a factor of $2$ by setting $\epsilon=1$). Thus, the total cost of this part of the solution is bounded by $2r_{n,m,\ell}\cdot c_{i}\cdot OPT$ (noting that the optimal solution for $R_{1}$ is no more than the optimal solution for $R$).
  2. 2.
    Now consider the cost of the second part of the solution. From the analysis of the weighted version of Algorithm 3, we know that the total cost of the outlier set is at most $O(\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i}))$ This means
    $\sum_{i\in R_{2}}d(s_{i},t_{i})$ $\leq\sum_{x\in K_{i^{*}}}\sum_{i\in R_{x}}d(s_{i},t_{i})$
    $\leq\sum_{x\in K_{i^{*}}}w_{x}$
    $\leq O(\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i})).$
    The first line comes immediately from the definition of $R_{x}$ (the set of requests involving $x$) and $R_{2}$; the second comes from the definition of $w_{x}$; and the third comes from our bound on the outlier set value of Algorithm 3.
    Consider a naive algorithm that simply picks an arbitrary order on the requests and services the pick-ups and drop-offs one at a time in that order. The cost of the $j$th pickup and drop-off is $d(s_{j},t_{j})$, and the cost of moving from $t_{j}$ to $s_{j+1}$ is $d(t_{j},s_{j+1})\leq\Delta$. In total we service at most $\text{OutlierCost}(c_{i})$ requests in $R_{2}$, as we assume each request has a cost at least $1$. Thus, the total cost incurred on this part of the solution is bounded by $O(\text{OutlierCost}(c_{i})\log^{2}\mathcal{K}(c_{i})+\text{OutlierCost}(c_{i})\cdot\Delta)$.
    Additionally, we know that the optimal cost on this set of requests is bounded by $OPT$, and the number of points we need to embed to get a solution is at most $2|R_{2}|$. Thus, we can use a standard Fakcharoenphol et al. [23] embedding on the set of points involved in this request and we get a solution of value at most $O(r_{2|R_{2}|,m,\ell}\log|R_{2}|OPT)$.
  3. 3.
    Finally, consider the cost of the algorithm moving from the last drop-off spot $t_{j}$ in $R_{1}$ to the first pickup spot $s_{j^{\prime}}$ in $R_{2}$. This total cost is $d(t_{j},s_{j^{\prime}})$, but this is upper bounded by $OPT$, as at some point any optimal algorithm must visit both locations.